### v3

# Simple Estimation of Semiparametric Models with Measurement Errors

Abstract
We develop a practical way of addressing the Errors-In-Variables (EIV) problem in the Generalized Method of Moments (GMM) framework. We focus on the settings in which the variability of the EIV is a fraction of that of the mismeasured variables, which is typical for empirical applications. For any initial set of moment conditions our approach provides a “corrected” set of moment conditions that are robust to the EIV. We show that the GMM estimator based on these moments is $\sqrt{n}$-consistent, with the standard tests and confidence intervals providing valid inference. This is true even when the EIV are so large that naive estimators (that ignore the EIV problem) are heavily biased with their confidence intervals having 0% coverage. Our approach involves no nonparametric estimation, which is especially important for applications with many covariates, and settings with multivariate or non-classical EIV. In particular, the approach makes it easy to use instrumental variables to address EIV in nonlinear models.
Keywords: errors-in-variables, nonstandard asymptotic approximation, nonclassical measurement errors, nonparametric identification, instrumental variables
1 Introduction
Measurement errors are a common problem for empirical studies. Addressing the Errors-In-Variables (EIV) bias in nonlinear models requires elaborate strategies. [^1] See Hausman et al. (1991, 1995); Newey (2001); Schennach (2007); Li (2002); Schennach (2004); Chen et al. (2005); Hu and Schennach (2008); Schennach (2014); Wilhelm (2019), among others. Despite the fundamental theoretical progress in identification and estimation of nonlinear models with EIV, the problem of EIV is still rarely addressed in empirical work outside of linear specifications.
The goal of this paper is to develop a simple and practical approach to estimation of nonlinear semiparametric models that can be expressed in the form of general moment conditions
$\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta)]=0\text{ iff }\theta=\theta_{0}\text{,}$ (1)
where $g\left(\cdot\right)$ is a vector of moment functions and $\theta_{0}$ is the parameter vector of interest. The researcher has a random sample of $\left\{X_{i},S_{i}\right\}_{i=1}^{n}$, where scalar or vector $X_{i}$ is a mismeasured version of unobserved $X_{i}^{\ast}$ with measurement error $\varepsilon_{i}$:
$X_{i}=X_{i}^{\ast}+\varepsilon_{i}.$
The measurement error can be classical or non-classical. We will refer to $g\left(\cdot\right)$ as the original moment function, since it would have been valid had the researcher observed $X_{i}^{\ast}$. A naive GMM estimator (that ignores the EIV and uses $X_{i}$ in place of $X_{i}^{\ast}$) based on $g\left(\cdot\right)$ is biased because $\mathbb{E}[g(X_{i},S_{i},\theta_{0})]\neq 0$, in contrast to equation (1).
Example (Nonlinear Regression, NLR).
Let $Y_{i}$ denote a scalar outcome, and let $X_{i}^{\ast}$ and $W_{i}$ be the covariates. Suppose
$E\left[Y_{i}|X_{i}^{\ast},W_{i}\right]=\rho\left(X_{i}^{\ast},W_{i},\theta_{0}\right)$ (2)
for some function $\rho$ known up to the parameter $\theta$. For example, in the Logit model, $Y_{i}$ is binary, $\rho\left(x,w,\theta\right)\equiv 1\left/\left(1+\exp\left(-\left(\theta_{x}^{\prime}x+\theta_{w}^{\prime}w\right)\right)\right)\right.$, and $\theta\equiv\left(\theta_{x}^{\prime},\theta_{w}^{\prime}\right)^{\prime}$.
Suppose the researcher has an instrumental variable $Z_{i}$. Then, they can use
$g\left(y,x,w,z;\theta\right)\equiv\left(y-\rho\left(x,w,\theta\right)\right)\varphi\left(x,w,z\right)$
as the original moment function, where $\varphi\left(x,w,z\right)$ is a vector that, for example, can include $x$, $z$, $w$, their powers and/or interactions. [^2] Note that the moment condition (1) is stated in terms of the true (correctly measured) $X_{i}^{\ast}$. Determining what functions $g\left(\cdot\right)$ (or $h\left(\cdot\right)$ in the NLR model) satisfy this moment condition does not involve any consideration of the measurement errors and hence is straightforward. Here $S_{i}=\left(Y_{i},W_{i},Z_{i}\right)$. $\blacksquare$
Even in this well-studied example of nonlinear regression, estimation in the presence of the EIV is a difficult problem. Importantly, nonlinear instrumental variable regression estimator cannot be used, since it is inconsistent in the presence of EIV (Amemiya, 1985). The existing approaches typically require nonparametric estimation that can be impractical in many empirical applications. In contrast, in this paper, we develop an alternative class of estimators, that are essentially GMM estimators that modify the original moment functions $g\left(\cdot\right)$ in a way that makes the moment conditions robust to the EIV. In particular, our approach makes it easy to use instrumental variables to address EIV in nonlinear models.
To provide a practical estimation approach for the general class of models (1), we focus on the empirical settings in which the researcher believes the variability of the measurement error to be at most a fraction of the variability of the mismeasured variable, i.e., the noise-to-signal ratio $\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}}$ to be moderate, e.g., $\tau\lesssim 0.5$ (we later discuss how the appropriateness of this assumption can be checked). The absolute magnitude of the measurement error $\sigma_{\varepsilon}$ does not need to be small. Focusing on these settings allows us to isolate the most important aspects of the problem and, as result, to develop a simple estimator, which does not require any nonparametric estimation or simulation. Such simple estimation becomes possible because in these settings we can obtain a simple approximation of the EIV bias of the moment conditions as a function of $\theta$.
We propose to bias correct the original moments $g\left(\cdot\right)$, which in turn removes the bias of the corresponding estimator of $\theta_{0}$. This bias correction depends on some moments of the distribution of the measurement errors that are unknown. Another difficulty is that the estimators of some components of the bias correction themselves may need to be bias corrected. To address these issues, we develop the corrected moment conditions, which depend on $\theta$ and additional parameters $\gamma$ that govern the bias correction. The true parameter value $\gamma_{0}$ is associated with (possibly conditional) low-order moments of $\varepsilon_{i}$. Despite some theoretical subtleties with the construction of the corrected moment conditions, their practical implementation is straightforward and they can be automatically computed for any original moment function $g\left(\cdot\right)$.
We introduce the Measurement Error Robust Moments (MERM) estimator, which is a GMM estimator that uses the corrected moment conditions to jointly estimate parameters $\theta_{0}$ and $\gamma_{0}$. The estimator can be computed using any standard software for GMM estimation. Joint estimation of parameters $\theta_{0}$ and $\gamma_{0}$ using the corrected moment conditions effectively robustifies moment conditions $g\left(\cdot\right)$ against the impact of the measurement errors.
To make these ideas precise and to study the properties of the proposed estimators, we develop an asymptotic theory using a nonstandard asymptotic approximation that models $\tau$ as slowly shrinking with the sample size. Standard asymptotics considers $\tau$ to be constant, which implies that as $n\rightarrow\infty$ the bias of a naive estimator dwarfs its sampling variability: the bias is constant while the standard errors shrink proportionally to $1/\sqrt{n}$. As a result, under the standard asymptotics, the problem of removing the EIV bias becomes central in the analysis, with relatively little attention paid to the sampling variability of estimators. However, this focus does not seem to be appropriate in many empirical applications, in which the researcher does not expect the potential EIV bias to be several orders of magnitude larger than the standard errors. [^3] Such empirical settings appear to be widespread. Although the concerns about measurement errors are often raised, the majority of applied work does not explicitly correct the EIV bias in nonlinear models, and instead implicitly or explicitly argues or conjectures that the EIV bias is likely not to be too large. See also the review of Bound et al. (2001). By considering $\tau$ as drifting towards zero with the sample size, our approach provides a better guidance on construction of EIV robust estimators with good finite sample properties when $\tau$ is small or moderate. [^4] Nonstandard asymptotic approximations with drifting parameters are often used to obtain better approximations of the finite sample behavior of estimators and tests. For example, in the instrumental variable regression settings, to consider the settings with relatively small first stage coefficients, Staiger and Stock (1997) model them as shrinking with $n$. It is important to keep in mind that such nonstandard asymptotic approximations are merely mathematical tools. One should not take them literally and think of parameters somehow changing if more data is collected.
Using this approximation, we show that the proposed estimation approach indeed addresses the EIV problem. The MERM estimator is shown to be $\sqrt{n}$-consistent and asymptotically normal and unbiased. The standard confidence intervals and tests for GMM estimators are also valid for the MERM estimator. Additionally, the standard GMM arsenal of assessment tools can be applied to the MERM estimator, allowing one to test model identification, conduct valid inference, and perform model specification diagnostics.
The usefulness of a large sample theory is measured by its ability to approximate the finite sample properties of the estimators and inference procedures. Thus, we study the MERM estimators in a variety of simulation experiments. The results confirm that the nonstandard asymptotic theory indeed provides a good approximation of the finite sample properties of the estimators even in the settings with relatively large EIV. In some of the simulation experiments, the EIV are so large that for the naive estimators’ standard $95\%$ confidence intervals have actual coverages of $0\%$ in finite samples, due to the magnitude of the EIV bias. At the same time, even in these settings the MERM estimators perform well, removing the EIV bias and providing confidence intervals with the correct coverage. In particular, the simulation results show that despite the simplicity of implementation, the MERM estimators can compete with and outperform semi-nonparametric estimators.
The MERM estimator is structurally different from the existing approaches that require nonparametric estimation of some nuisance parameters, for example, of the density $f_{X_{i}^{\ast}|W}$. Avoiding nonparametric estimation has at least two advantages. First, since the majority of empirical applications include at least a handful of additional covariates $W_{i}$, nonparametric estimation is often infeasible due to the curse of dimensionality. Because the MERM estimator does not involve any nonparametric estimation, it can be used in applications with a relatively large number of additional covariates $W_{i}$, and remains feasible even in the more complicated settings, including multi-equation and structural models, and applications with multiple mismeasured variables $X_{i}$. Second, estimation of infinite-dimensional nuisance parameters is typically more demanding towards the sources of identification available in the data, for example, requiring an instrumental variable with a large support (continuously distributed). In contrast, having a discrete instrument is sufficient for the MERM approach because the nuisance parameter $\gamma_{0}$ is finite-dimensional.
The simplicity and practicality of the MERM approach do come at a cost: there is a limit on the magnitude of the measurement errors it can handle. For example, one generally should not expect the MERM approach to work well when $\tau>1$, i.e., when the noise dominates the signal; in this case the researcher should seek an alternative estimation method.
We view the MERM approach as providing a bridge between the settings in which the measurement errors are guaranteed to be absent or negligible, and the settings where the measurement errors are so large that one has to use the relatively more complicated estimators from the earlier literature (if they exist at all for the model of interest).
Related Literature Chen et al. (2011), Schennach (2016), and Schennach (2020) provide excellent overviews of the measurement error literature.
The existing semiparametric approaches to estimation and inference in models with EIV involve nonparametric estimation of infinite-dimensional nuisance parameters (e.g., Chesher2000WP,Li2002JoE,Schennach2004Ecta,Schennach2007Ecta,HuSchennach2008Ecta,SchennachHu2013JASA,Song2015JoE), simulation (e.g., Schennach2014Ecta-ELVIS), or both (e.g., Newey2001REStat,WangHsiao2011JoE)))). The exceptions include models with linear and polynomial regression functions (see HINP1991JoE,HausmanNeweyPowell1995JoE)))), and Gaussian control variable models such as Probit and Tobit with endogeneity (see SmithBlundell1986Ecta,RiversVuong1988JoE).
To the best of our knowledge, this paper is the first to provide an approach for $\sqrt{n}$-consistent and asymptotically normal and unbiased estimation of general GMM models with EIV that does not require any nonparametric estimation (or simulation).
We are able to provide such an estimator because we focus on the models with moderate measurement errors. Modeling the variance of the measurement error as shrinking to zero with the sample size is a popular approach in Statistics. The method has been proposed by Wolter and Fuller (1982), who used it to construct an approximate MLE estimator of a nonlinear regression model with Gaussian errors. Following their approach, the Statistics literature has mainly focused on the settings where the moments of the EIV needed to bias correct the estimators are either known or can be directly estimated from the available data such as repeated measurements (e.g., CarrollStefanski1990JASA,CarrollEtAl2006Book-ME). In Economics, such data are relatively rare. The use of approximations with shrinking variance of measurement errors in Econometrics literature has been pioneered by Kadane (1971), Amemiya (1985), and Chesher (1991). Such approximations have been used to check the sensitivity of naive estimators to the EIV by considering how the estimates change as the unknown moments of the measurement errors vary within some set of plausible values, e.g., see Chesher and Schluter (2002), Chesher et al. (2002), Battistin and Chesher (2014), Chesher (2017), and Hong and Tamer (2003).
This paper differs from the earlier literature in several ways. First, it presents a way to estimate the unknown nuisance parameters (moments of the measurement errors) jointly with the parameters of interest. As a result, the approach can, for example, use instrumental variables as a source of identification. Second, the method applies to a very general class of semiparametric models specified by moment conditions. Third, the MERM approach allows the measurement errors to have larger magnitudes than most of the papers in the earlier literature; this is achieved by the MERM approach recursively bias correcting the bias correction terms. Fourth, our approach allows relaxing the assumption of classical measurement errors.
The most widespread approach to identification of the EIV models in economic applications is to use instrumental variables, e.g., see Hausman et al. (1991); Newey (2001); Schennach (2007); Wang and Hsiao (2011). In a recent paper, Hahn et al. (2021) reconsider the regression model in Amemiya (1990) using a bias correction similar to ours. When proper excluded variables are not available, researchers have considered using higher moments of $X_{i}$ as instruments, e.g., see Reiersøl (1950); Lewbel (1997); Erickson and Whited (2002); Schennach and Hu (2013); Ben-Moshe et al. (2017). When available, repeated measurements can also be used to identify the model, e.g., see Hausman et al. (1991); Li and Vuong (1998); Li (2002); Schennach (2004). The MERM estimator accommodates these identification approaches within a unified estimation framework.
The power of the general MERM approach can be illustrated in the NLR model. For example, when a candidate instrumental variable is available, the conditions it needs to satisfy are much weaker than what is required by many existing approaches. Availability of a discrete instrument is sufficient for identification; and the instrument is allowed to have heterogeneous impact on covariates $X_{i}^{\ast}$. One can also take a nonclassical, nonlinear (e.g., discretized or censored), or biased measurement of $X_{i}^{\ast}$ as an instrument in the MERM approach. We discuss identification in Section 2.4. In addition, in a related paper Evdokimov and Zeleneev (2022) study nonparametric regression with EIV using the $\tau\rightarrow 0$ approximation, and demonstrate that the MERM approach can also be motivated from a nonparametric perspective.
Kitamura et al. (2013); Andrews et al. (2017); Armstrong and Kolesár (2021); Bonhomme and Weidner (2022), among others, develop tools for estimation and inference in GMM, which are robust to general perturbation or misspecification of the true data generating process. They focus on the settings in which these perturbations are sufficiently small, so that naive estimators remain $\sqrt{n}$-consistent, and their biases are of the same order of magnitude as their standard errors. In contrast, we focus on more specific forms of data contamination due to the EIV. This allows the MERM approach to remain valid even in the settings with larger measurement errors, in which naive estimators may have slower than $\sqrt{n}$ rates of convergence.
The MERM approach also provides a useful foundation for dealing with EIV in more complicated settings. Evdokimov and Zeleneev (2018) utilize the MERM framework to address an issue of nonstandard inference, which turns out to arise generally when EIV models are identified using instrumental variables. Evdokimov and Zeleneev (2019) extend the analysis of this paper to long panel and network settings.
Organization of the paper Section 2 introduces the Moderate Measurement Error framework and the proposed MERM estimator. Section 3 presents several Monte Carlo experiments that illustrate finite sample properties of the MERM estimators. Section 4 considers several extensions of the framework.
2 Moderate Measurement Errors Framework
To present the main ideas we first consider the case of univariate $X_{i}^{\ast}$ and classical measurement error $\varepsilon_{i}$. Later we consider multivariate $X_{i}^{\ast}$ and non-classical measurement errors $\varepsilon_{i}$.
To develop a practical estimation approach for general moment condition models we focus on the settings in which $\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}}$ is small or moderate. We consider an asymptotic approximation with $\tau_{n}\equiv\tau\rightarrow 0$ as $n\rightarrow\infty$. Note that economically meaningful parameters are usually invariant to rescaling of $X_{i}^{\ast}$. Likewise, the extent of the EIV problem does not change with such rescaling.
[…]
2.1 Special Case: Quadratic Expansion
For clarity, we first consider a simple special case of the general approach. Let us denote $g_{x}^{(k)}\left(x,s,\theta\right)\equiv\partial^{k}g\left(x,s,\theta\right)/\partial x^{k}$. Since $\mathbb{E}[\left|\varepsilon_{i}\right|^{k}]\propto\tau_{n}^{k}\rightarrow 0$ as $n\rightarrow\infty$, under some regularity conditions, we can write the quadratic Taylor expansion of function $g(X_{i},S_{i},\theta)=g(X_{i}^{\ast}+\varepsilon_{i},S_{i},\theta)$ around $\varepsilon_{i}=0$ as
$\mathbb{E}[g(X_{i},S_{i},\theta)]$ $=\mathbb{E}\left[g(X_{i}^{\ast},S_{i},\theta)+g_{x}^{(1)}(X_{i}^{\ast},S_{i},\theta)\varepsilon_{i}+\dfrac{1}{2}g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta)\varepsilon_{i}^{2}\right]+O(\mathbb{E}[\left|\varepsilon_{i}\right|^{3}])$
$=\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta)]+\dfrac{\mathbb{E}[\varepsilon_{i}^{2}]}{2}\mathbb{E}[g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta)]+O(\tau_{n}^{3}),$ (3)
where the second equality holds because because $\varepsilon_{i}$ and $\left(X_{i}^{\ast},S_{i}\right)$ are independent, and $\mathbb{E}\left[\varepsilon_{i}\right]=0$.
This expansion implies that $\mathbb{E}[g(X_{i},S_{i},\theta_{0})]=O\left(\sigma_{\varepsilon}^{2}\right)=O(\tau_{n}^{2})$. As a result, a naive estimator that ignores the EIV and uses $X_{i}$ in place of $X_{i}^{\ast}$ has EIV bias of order $\tau_{n}^{2}$. [^6] For example, consider a linear regression with a scalar mismeasured regressor. The bias of the naive OLS estimator of the slope parameter $\theta_{01}$ is $-\theta_{01}\frac{\tau_{n}^{2}}{1+\tau_{n}^{2}}=-\theta_{01}\tau_{n}^{2}+O\left(\tau_{n}^{4}\right)$. Bias of the naive estimator should be compared with its standard error, which is of order $n^{-1/2}$. Bias of the naive estimator is not negligible, unless the measurement error is rather small (theoretically, unless $\tau_{n}^{2}=o\left(n^{-1/2}\right)$). In particular, tests and confidence intervals based on the naive estimator are invalid and can provide highly misleading results. Moreover, if $\tau_{n}^{2}$ shrinks at a rate slower than $O\left(n^{-1/2}\right)$, the rate of convergence of the naive estimator is slower than $\sqrt{n}$.
Suppose $\tau_{n}=o\left(n^{-1/6}\right)$. Then, $O(\tau_{n}^{3})=o\left(n^{-1/2}\right)$ and we can rearrange equation (3) as
$\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta)]=\mathbb{E}[g(X_{i},S_{i},\theta)]-\frac{\mathbb{E}[\varepsilon_{i}^{2}]}{2}\mathbb{E}\left[g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta)\right]+o(n^{-1/2}).$ (4)
The left-hand side of this equation is exactly the moment condition (1) that we would like to use for estimation of $\theta_{0}$. The first term on the right-hand side involves only observed variables, and can be estimated by the sample average $\overline{g}(\theta)\equiv n^{-1}\sum_{i=1}^{n}g(X_{i},S_{i},\theta).$ The second term on the right-hand side can be thought of as a bias correction that removes the EIV-bias from the expected moment function $\mathbb{E}[g(X_{i},S_{i},\theta)]$.
The idea of the MERM estimator we propose is to make use of expansions such as (4) to bias correct the moment condition $\mathbb{E}[g(X_{i},S_{i},\theta)]$, which in turn removes the bias of the estimator of the parameters of interest $\theta_{0}$. To perform the bias correction we need to estimate two quantities: $\mathbb{E}[\varepsilon_{i}^{2}]$ and $\mathbb{E}[g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta)]$.
First, we show that in equation (4) we can substitute $\mathbb{E}[g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta)]$ with $\mathbb{E}[g_{x}^{(2)}(X_{i},S_{i},\theta)]$, which in turn can be estimated by $\overline{g}_{x}^{(2)}(\theta)\equiv n^{-1}\sum_{i=1}^{n}g_{x}^{(2)}(X_{i},S_{i},\theta)$. By the Taylor expansion around $\varepsilon_{i}=0$ similar to equation (3), we can show that $\mathbb{E}[g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta)]=\mathbb{E}[g_{x}^{(2)}(X_{i},S_{i},\theta)]+O(\tau_{n}^{2})$ and hence
$\frac{1}{2}\mathbb{E}[\varepsilon_{i}^{2}]\left(\mathbb{E}\left[g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta)\right]-\mathbb{E}\left[g_{x}^{(2)}(X_{i},S_{i},\theta)\right]\right)=\mathbb{E}[\varepsilon_{i}^{2}]O\left(\tau_{n}^{2}\right)=O\left(\tau_{n}^{4}\right).$ (5)
Here $O\left(\tau_{n}^{4}\right)=o\left(n^{-1/2}\right)$ because we assume that $\tau_{n}=o\left(n^{-1/6}\right)$. The idea behind this substitution is that the bias of order $O(\tau_{n}^{2})$ in $\mathbb{E}[g_{x}^{(2)}(X_{i},S_{i},\theta)]$ can be ignored because it is multiplied by $E\left[\varepsilon_{i}^{2}\right]=O\left(\tau_{n}^{2}\right)$. [^7] Such substitutions of $X^{\ast}$ with $X$ have been used in other contexts, e.g., Chesher and Schluter (2002). With the substitution, we can rearrange equation (4) and write it as
$\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta)]=\mathbb{E}\left[g(X_{i},S_{i},\theta)-\frac{\mathbb{E}[\varepsilon_{i}^{2}]}{2}g_{x}^{(2)}(X_{i},S_{i},\theta)\right]+o(n^{-1/2}).$ (6)
Second, we propose estimating the unknown $\mathbb{E}[\varepsilon_{i}^{2}]$ together with the parameter of interest $\theta$. Specifically, let $\gamma_{02}\equiv\mathbb{E}[\varepsilon_{i}^{2}]/2$ denote the true value of parameter $\gamma_{2}$, and consider the following corrected moment function:
$\psi(X_{i},S_{i},\theta,\gamma)\equiv g(X_{i},S_{i},\theta)-\gamma_{2}g_{x}^{(2)}(X_{i},S_{i},\theta).$ (7)
Function $\psi$ is a moment function parameterized by $\theta$ and $\gamma$, and
$\mathbb{E}[\psi(X_{i},S_{i},\theta_{0},\gamma_{02})]=\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta_{0})]+o\left(n^{-1/2}\right)=o\left(n^{-1/2}\right),$ (8)
where the first equality follows from equation (6) and the definition of $\gamma_{02}$, and the second equality follows from equation (1). Hence, the corrected moment conditions $\psi$ can be used to jointly estimate the true parameters $\theta_{0}$ and $\gamma_{02}$ by a GMM estimator. [^8] In the moment condition settings, having $o\left(n^{-1/2}\right)$ is equivalent to having 00 on the right-hand side of equation (8).
Remark 1.
If $\mathbb{E}[\varepsilon_{i}^{3}]=0$ (e.g., if the distribution of $\varepsilon_{i}$ is symmetric), the remainder in equation (3) is of a smaller order $O(\tau_{n}^{4})$. Hence, the corrected moments (8) remain valid for larger values of $\tau_{n}$, requiring only the weaker condition $\tau_{n}=o(n^{-1/8})$. The bias of the naive estimators in this case can be as large as $o(n^{-1/4})$.
2.2 General Case: Expansion of order $K$
The quadratic expansion of equation (3) can be extended to general order $K\geq 2$. Considering larger $K$ theoretically allows $\tau_{n}$ converging to zero at a slower rate. In finite samples this corresponds to the asymptotics providing good approximations for larger values of $\tau_{n}$, i.e., large measurement errors. Expanding $g(X_{i}^{\ast}+\varepsilon_{i},S_{i},\theta)$ around $\varepsilon_{i}=0$ we have,
$\mathbb{E}[g(X_{i},S_{i},\theta)]=\mathbb{E}\left[g(X_{i}^{\ast},S_{i},\theta)+\sum_{k=1}^{K}\frac{\varepsilon_{i}^{k}}{k!}g_{x}^{(k)}(X_{i}^{\ast},S_{i},\theta)\right]+O\left(\mathbb{E}\left[\left|\varepsilon_{i}\right|^{K+1}\right]\right).$ (9)
The above special case of quadratic expansion corresponds to $K=2$.
The approximation we consider is formalized by the following assumption.
Assumption MME.
(Moderate Measurement Errors) (i) $\tau_{n}=o(n^{-1/\left(2K+2\right)})$ for some integer $K\geq 2$; and (ii) $\mathbb{E}[\left|\varepsilon_{i}\right|^{L}]\leq C\sigma_{\varepsilon}^{L}$ for some $L\geq K+1$ and $C>0$.
Assumption 2.1 (i) limits the magnitude of the measurement errors and implies that $\tau_{n}^{K+1}=o\left(n^{-1/2}\right)$. Assumption 2.1 (ii) implies that $\mathbb{E}[\left|\varepsilon_{i}\right|^{k}]=O\left(\sigma_{\varepsilon}^{k}\right)$, and requires the tails of $\varepsilon_{i}/\sigma_{\varepsilon}$ to be sufficiently thin. Together, parts (i) and (ii) imply that $\mathbb{E}[\left|\varepsilon_{i}\right|^{K+1}]=O\left(\tau_{n}^{K+1}\right)=o\left(n^{-1/2}\right)$, and hence ensure that the remainder in equation (9)  is negligible. Using $\mathbb{E}\left[\varepsilon_{i}|X_{i}^{\ast},S_{i}\right]=0$ to further simplify this expansion and rearranging the terms we obtain
$\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta)]=\mathbb{E}[g(X_{i},S_{i},\theta)]-\sum_{k=2}^{K}\frac{\mathbb{E}[\varepsilon_{i}^{k}]}{k!}\mathbb{E}\left[g_{x}^{(k)}(X_{i}^{\ast},S_{i},\theta)\right]+o(n^{-1/2}).$ (10)
This equation is the general expansion analog of equation (4). The summation on the right hand side is the bias correction term, which we use to construct the MERM estimator.
It turns out that for $K\geq 4$, estimation of $\mathbb{E}[g_{x}^{(k)}(X_{i}^{\ast},S_{i},\theta)]$ is more intricate than in the case of $K=2$, and the substitution we made in equation (6) no longer works. Larger values of $K$ allow for larger values of $\tau_{n}$ and hence larger EIV biases of naive estimators $n^{-1}\sum_{i=1}^{n}g_{x}^{(k)}(X_{i},S_{i},\theta)$. The expansion of order $K$ includes terms up to the order $\tau_{n}^{K}$, with the asymptotically negligible remainder of order $O\left(\tau_{n}^{K+1}\right)$. For $K\geq 4$, terms of order $\tau_{n}^{4}$ are not negligible. This implies that we cannot ignore the EIV bias that would arise from substituting $\mathbb{E}[g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta)]$ with $\mathbb{E}[g_{x}^{(2)}(X_{i},S_{i},\theta)]$ in equation (10), because this bias is of order $O\left(\tau_{n}^{4}\right)$ according to equation (5). To address this problem, we instead replace $\mathbb{E}[g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta)]$ with the bias corrected expression $\mathbb{E}[g_{x}^{(2)}(X_{i},S_{i},\theta)]-\left(\mathbb{E}[\varepsilon_{i}^{2}]/2\right)\mathbb{E}[g_{x}^{(4)}(X_{i},S_{i},\theta)]$. Thus, for $K\geq 4$, one needs to bias correct the estimator of the bias correction term. Moreover, for larger $K$ one needs to bias correct the bias correction of the bias correction term and so on.
Fortunately, we show that these bias corrections can be constructed as linear combinations of the expectations of the higher order derivatives of $g_{x}^{\left(k\right)}(X_{i},S_{i},\theta)$. Let us define the following corrected moment function:
$\psi(X_{i},S_{i},\theta,\gamma)\equiv g(X_{i},S_{i},\theta)-\sum_{k=2}^{K}\gamma_{k}g_{x}^{(k)}(X_{i},S_{i},\theta),$ (11)
where $\gamma=(\gamma_{2},\dots,\gamma_{K})^{\prime}$ is a $K-1$ dimensional vector of parameters. Let $\gamma_{0}\equiv(\gamma_{02},\dots,\gamma_{0K})^{\prime}$ denote the vector of true parameters $\gamma_{0k}$, defined as
$\gamma_{02}\equiv\frac{\mathbb{E}\left[\varepsilon_{i}^{2}\right]}{2}\text{,}\;\;\gamma_{03}\equiv\frac{\mathbb{E}\left[\varepsilon_{i}^{3}\right]}{6},\;\;\text{and}\;\;\gamma_{0k}\equiv\frac{\mathbb{E}\left[\varepsilon_{i}^{k}\right]}{k!}-\sum_{\ell=2}^{k-2}\frac{\mathbb{E}\left[\varepsilon_{i}^{k-\ell}\right]}{(k-\ell)!}\gamma_{0\ell}\;\;\text{for}\;\;k\geq 4.$ (12)
We formalize this discussion below.
Assumption CME.
$\varepsilon_{i}$ is independent from $(X_{i}^{\ast},S_{i})$ and $\mathbb{E}[\varepsilon_{i}]=0$.
Assumption 2.2 is the classical measurement error assumption. We relax this assumption later in Section 4.2. The following lemma establishes validity of the corrected moment conditions under Assumptions 2.1, 2.2, and some mild regularity conditions provided in Appendix A.
Lemma 1.
Under Assumptions 2.1, 2.2 and A.1 in Appendix A,
$\mathbb{E}[\psi(X_{i},S_{i},\theta_{0},\gamma_{0})]=\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta_{0})]+o\left(n^{-1/2}\right)=o\left(n^{-1/2}\right).$
Lemma 1 implies that the corrected moment conditions $\psi$ are valid and can potentially be used to jointly estimate parameters $\theta_{0}$ and $\gamma_{0}$. The total number of parameters to be estimated is now $\dim\left(\theta\right)+K-1$. Thus, joint estimation of $\theta_{0}$ and $\gamma_{0}$ requires that $\dim\left(\psi\right)=\dim\left(g\right)\geq\dim\left(\theta\right)+K-1$, i.e., that the original moment conditions $g$ include sufficiently many overidentifying restrictions. For example, the overidentifying restrictions can be constructed by using an instrumental variable; we discuss this in more detail below.
2.3 Measurement Error Robust Moments (MERM) estimator
The MERM estimator jointly estimates the parameters $\theta_{0}$ and $\gamma_{0}$ using moment conditions $\psi$. It is convenient to define the joint vector of parameters
$\beta\equiv(\theta^{\prime},\gamma^{\prime})^{\prime},\ \beta_{0}\equiv(\theta_{0}^{\prime},\gamma_{0}^{\prime})^{\prime},\text{ }\hat{\beta}\equiv(\hat{\theta}^{\prime},\hat{\gamma}^{\prime})^{\prime},$
and the parameter space $\mathcal{B}\equiv\Theta\times\Gamma$, where $\Theta$ and $\Gamma$ are the parameter spaces for $\theta$ and $\gamma$. Then, MERM estimator is the GMM estimator (Hansen1982Ecta):
$\hat{\beta}\equiv\operatorname*{\mathrm{arg}\!\min\limits}_{\beta\in\mathcal{B}}\hat{Q}(\beta),\qquad\hat{Q}(\beta)\equiv\overline{\psi}(\beta)^{\prime}\hat{\Xi}\overline{\psi}(\beta),$ (13)
where $\overline{\psi}(\beta)\equiv n^{-1}\sum_{i=1}^{n}\psi_{i}(\beta)$, $\psi_{i}(\beta)\equiv\psi\left(X_{i},S_{i},\beta\right)$, $\hat{\Xi}$ is a weighting matrix, and $\hat{Q}(\beta)$ is the standard GMM objective function.
While Lemma 1 establishes validity of the corrected moment restrictions $\psi$, the MERM estimator also relies on $\beta_{0}$ being identified from $\psi$. This requirement is formalized by the following assumption.
Assumption ID.
  1. (i)
    the Jacobian $\Psi^{\ast}$ has full column rank, where
    $\Psi^{\ast}$ $\equiv\mathbb{E}\left[\nabla_{\theta}\psi(X_{i}^{\ast},S_{i},\theta_{0},0),\nabla_{\gamma}\psi(X_{i}^{\ast},S_{i},\theta_{0},0)\right]$
    $=\mathbb{E}\left[\nabla_{\theta}g(X_{i}^{\ast},S_{i},\theta_{0}),-g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta_{0}),\dots,-g_{x}^{(K)}(X_{i}^{\ast},S_{i},\theta_{0})\right];$
  2. (ii)
    $\mathbb{E}\left[\psi(X_{i}^{\ast},S_{i},\theta,\gamma)\right]=0$ iff $\theta=\theta_{0}$ and $\gamma=0$.
Assumptions 2.3 (i) and (ii) are the standard GMM local and global identification conditions applied to the moment function $\psi(X_{i}^{\ast},S_{i},\theta,\gamma)$. These are high-level conditions, which we will return to later in the paper. The moment conditions formulation is sufficiently general to encompass a wide variety of sources of identification. In Section 2.4, we discuss identification in detail and illustrate the construction of the moment function using an instrumental variable or a second measurement.
Under some additional regularity conditions, estimator $\hat{\beta}$ behaves as a standard GMM-type estimator: it is $\sqrt{n}$-consistent and asymptotically normal and unbiased. This result is formalized by the following theorem.
Theorem 2 (Asymptotic Normality).
Suppose that $\{(X_{i}^{\ast},S_{i}^{\prime},\varepsilon_{i})\}_{i=1}^{n}$ are i.i.d.. Then, under Assumptions 2.1, 2.2, 2.3, and A.1-A.3 in Appendix A,
[…]
$\Sigma\equiv(\Psi^{\prime}\Xi\Psi)^{-1}\Psi^{\prime}\Xi\Omega_{\psi\psi}\Xi\Psi(\Psi^{\prime}\Xi\Psi)^{-1}.$
Theorem 2 shows that the MERM approach addresses the EIV bias problem, and in particular provides a $\sqrt{n}$-consistent asymptotically normal and unbiased estimator $\hat{\theta}$, which can be used to conduct inference about the true parameters $\theta_{0}$. The asymptotic variance $\Sigma$ takes the standard sandwich form, with $\Psi\equiv\mathbb{E}\left[\nabla_{\beta}\psi_{i}(\beta_{0})\right]$, $\Omega_{\psi\psi}\equiv\mathbb{E}\left[\psi_{i}\left(\beta_{0}\right)\psi_{i}^{\prime}\left(\beta_{0}\right)\right]$, and $\hat{\Xi}\rightarrow_{p}\Xi$.
Remark 2.
Notice that the bias of naive estimators (such as a GMM estimator based on the original moment conditions) is $O(\tau_{n}^{2})$, so their rate of convergence is $O_{p}(\tau_{n}^{2}+n^{-1/2})$. The bias dominates sampling variability and naive estimators are not $\sqrt{n}$-consistent unless $\tau_{n}=O(n^{-1/4})$, i.e., unless the magnitude of the measurement error is rather small. At the same time, the MERM estimator remains $\sqrt{n}$-consistent for much larger values of $\tau_{n}$, up to $\tau_{n}=O(n^{-1/(2K+2)})$, whereas the rate of convergence of naive estimators is only $O_{p}(n^{-1/(K+1)})$ in this case.
Once the corrected moment condition $\psi$ is constructed, estimation of and inference about parameters $\beta_{0}$ can be performed using any standard software package for GMM estimation. In other words, the proposed estimator can be simply treated as a standard GMM estimator based on the corrected moment conditions $\psi$, and the conventional standard errors, tests, and confidence intervals are valid.
Remark 3.
Researchers may be interested in average effects of the form $\lambda_{0}\equiv\mathbb{E}\left[\lambda\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\right]$. In the NLR, one may be interested in the average partial effect $x$ (i.e., $\lambda_{0}\equiv\mathbb{E}\left[\nabla_{x}\rho\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\right]$) or another covariate. The naive average partial effect estimator $\hat{\lambda}_{\text{Naive}}\equiv\frac{1}{n}\sum_{i=1}^{n}\lambda(X_{i},S_{i},\hat{\theta})$ suffers from the EIV bias, unless function $\lambda$ is linear in $X_{i}^{\ast}$. Instead, one can use estimates $\hat{\gamma}$ to construct the bias-corrected estimator $\hat{\lambda}_{\text{MERM}}\equiv\frac{1}{n}\sum_{i=1}^{n}\left\{\lambda(X_{i},S_{i},\hat{\theta})-\sum_{k=2}^{K}\hat{\gamma}_{k}\lambda_{x}^{(k)}(X_{i},S_{i},\hat{\theta})\right\}$.
Remark 4.
[…]
Remark 5.
It is important to note that $\gamma_{0k}\neq\mathbb{E}\left[\varepsilon_{i}^{k}\right]\left/k!\right.$ for $k\geq 4$, contrary to what equation (10) might suggest. For example, $\gamma_{04}=\left(\mathbb{E}\left[\varepsilon_{i}^{4}\right]-6\sigma_{\varepsilon}^{4}\right)\left/24\right.$ is negative for many distributions, including normal. For instance, in the example of Remark 4, $\gamma_{04}\approx-0.0026$. The reason that generally $\gamma_{0k}\neq\mathbb{E}\left[\varepsilon_{i}^{k}\right]\left/k!\right.$ is that the estimators of the correction terms themselves need a correction, which is accounted for by the form of $\gamma_{0k}$. Since there is a one-to-one relationship between $\gamma_{0}$ and the moments $\mathbb{E}\left[\varepsilon_{i}^{\ell}\right]$, parameter space $\Gamma$ for $\gamma_{0}$ can incorporate restrictions that the moments must satisfy (e.g., $\sigma_{\varepsilon}^{2}\geq 0$ and $\mathbb{E}\left[\varepsilon_{i}^{4}\right]\geq\sigma_{\varepsilon}^{4}$). Such restrictions can increase the efficiency of the estimator and the power of tests.
Remark 6.
No parametric assumptions are imposed on the distribution of $\varepsilon_{i}$, i.e. the distribution of $\varepsilon_{i}$ is treated nonparametrically. The regularity conditions restrict only the magnitude of the moments of $\varepsilon_{i}$. The approach imposes no restrictions on the smoothness of the distributions of $X_{i}^{\ast}$ and $\varepsilon_{i}$, which are not even required to be continuous. Examples in which this can be useful include individual wages (whose distributions may have point masses at round numbers), and allowing the measurement error $\varepsilon_{i}$ to have a point mass at zero (a fraction of the population may have a zero measurement or recall error).
Remark 7.
Considering larger $K$ allows $\tau_{n}$converging to zero at a slower rate, which in finite samples corresponds to the asymptotics providing better approximations for larger magnitudes of measurement errors. On the other hand, taking a larger $K$ increases the dimension of the nuisance parameter $\gamma_{0}$ and thus typically increases the variance of $\hat{\theta}$.
Remark 8.
The formulas of the derivatives $g_{x}^{(k)}(\cdot)$ are typically easy to compute analytically or using symbolic algebra software. Alternatively, these derivatives can be computed using numerical differentiation. In either way, the corrected set of moments can be automatically produced for a generic moment function $g(\cdot)$ provided by the user.
2.4 Model Identification: Jacobian $\Psi$
Theorem 2 requires $\beta_{0}$ to be identified and the Jacobian matrix $\Psi$ to be full rank. Notably, the MERM framework encompasses many possible sources of identification at once, including instrumental variables, additional measurements, or nonlinearities of the functional form. The identifying information is incorporated in the moment functions. Essentially, our approach first characterizes in what directions the measurement errors can bias the moment conditions $\mathbb{E}\left[g\left(X_{i},S_{i},\theta\right)\right]$, and then uses the moments orthogonal to those directions for identification of $\theta_{0}$. To be more specific, we will now consider identification when in addition to the error-laden $X_{i}$ we have either (i) a general instrument $Z_{i}$ or (ii) a second measurement $Q_{i}$.
Identification Using A General Instrument $Z_{i}$
Many applications can be formulated as the following conditional moment restriction:
$\mathbb{E}\left[u\left(X_{i}^{\ast},S_{i},\theta\right)|X_{i}^{\ast}\right]=0\text{ iff }\theta=\theta_{0},$ (15)
for some moment function $u$. For example, consider the nonlinear regression model $\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=x\right]=\rho\left(x,\theta_{0}\right)$, then $u\left(x,y,\theta\right)=\rho\left(x,\theta\right)-y$ [^9] For simplicity of the exposition, in the expectation in equation (15) we only condition on $X_{i}^{\ast}$. The discussion applies in a straightforward way to the settings with additional correctly measured variables $W_{i}$ in the conditioning set, i.e., the model $\mathbb{E}\left[u\left(X_{i}^{\ast},S_{i},\theta_{0}\right)|X_{i}^{\ast},W_{i}\right]=0$. For example, in the nonlinear regression example with additional covariates $W_{i}$ we have $\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=x,W_{i}=w\right]=\rho\left(x,w,\theta_{0}\right)$, so $u\left(x,y,w,\theta\right)=\rho\left(x,w,\theta\right)-y$.
In applications, identification of the models with EIV would typically rely on an instrumental variable $Z_{i}$. Suppose the instrument satisfies the exclusion restriction $\mathbb{E}\left[u\left(X_{i}^{\ast},S_{i},\theta\right)|X_{i}^{\ast},Z_{i}\right]=\mathbb{E}\left[u\left(X_{i}^{\ast},S_{i},\theta\right)|X_{i}^{\ast}\right]$, i.e., conditional on the true $X_{i}^{\ast}$ the instrument has no further effect on the moment conditions $u$. Consider the moment functions $h\left(x,s,\theta\right)\equiv u\left(x,s,\theta\right)\otimes\varphi_{X}\left(x\right)$, where $\varphi_{X}\left(x\right)$ is a vector of functions of $x$, e.g., $\varphi_{X}\left(x\right)\equiv\left(1,x,\ldots,x^{J}\right)^{\prime}$. By the Law of Iterated Expectations, $\mathbb{E}\left[h\left(X_{i}^{\ast},S_{i},\theta_{0}\right)|Z_{i}=z\right]=0$ for all $z$. However, the same expectation with $X_{i}^{\ast}$ replaced by the observed $X_{i}$, $\mathbb{E}\left[h\left(X_{i},S_{i},\theta_{0}\right)|Z_{i}=z\right]$, will depend on $z$. One way to see this is to consider a Taylor expansion similar to equation (3):
$\mathbb{E}\left[h\left(X_{i},S_{i},\theta_{0}\right)|Z_{i}=z\right]=\underbrace{\mathbb{E}\left[h\left(X_{i}^{\ast},S_{i},\theta_{0}\right)|Z_{i}=z\right]}_{=0\text{ by the LIE}}+\gamma_{0}\mathbb{E}\left[h_{x}^{\left(2\right)}\left(X_{i}^{\ast},S_{i},\theta_{0}\right)|Z_{i}=z\right]+O\left(\tau_{n}^{3}\right),$
which shows that $\mathbb{E}\left[h\left(X_{i},S_{i},\theta_{0}\right)|Z_{i}=z\right]$ is zero for all $z$ (up to a negligible remainder) unless $\gamma_{0}\neq 0$, where $\gamma_{0}=\sigma^{2}/2$. Thus, $\mathbb{E}\left[h\left(X_{i},S_{i},\theta_{0}\right)|Z_{i}=z\right]$ varies with $z$ only because of the presence of the measurement error. Intuitively, the magnitude of this variation then identifies the nuisance parameters $\gamma_{0}$. Thus, one can rely on the original moment functions of the typical form $g\left(x,s,\theta\right)=h\left(x,s,\theta\right)\otimes\varphi_{Z}\left(z\right)$, where $\varphi_{Z}\left(z\right)$ is a vector of functions of $z$.
The above discussion provides the intuition for identification of the nonlinear moment condition models with EIV. It is important to note that identification of general nonlinear moment condition models is a complicated problem. Even in the settings without measurement errors, it is generally not possible to give low-level conditions guaranteeing that a specific set of nonlinear moment conditions identifies the parameter vector. The presence of EIV makes the question of identification even harder.
We attempt to address this concern and make the above intuitions more precise in two ways. First, in the following subsection we consider a specific (but frequently employed) kind of an instrument: a second measurement (possibly nonclassical). The specific form of the excluded variable allows us to provide more transparent identification conditions. Second, in Evdokimov and Zeleneev (2022) we study nonparametric regression model with EIV using the $\tau_{n}\rightarrow 0$ approximation. We show that the model is identified using an instrument (even a discrete one), and motivate the MERM approach from a nonparametric perspective. Finally, since MERM estimator is a standard GMM estimator, one can test the strength of identification of the model parameters, or conduct identification-robust inference using the standard methods (e.g., StockWright2000Ecta,Kleibergen2005Ecta,GuggenbergerSmith2005ET,GuggenbergerRamalhoSmith2012JoE,AndrewsMikusheva2016Ecta-ConditionalFunctionalNuisance,AndrewsI-2016-Ecta-CLC,AndrewsGuggenberger2019QE).
Identification Using A Second Measurement
Suppose we observe a second measurement
$Q_{i}=\alpha_{1}X_{i}^{\ast}+\varepsilon_{Q,i},$
where $\alpha_{1}$ may not be known. Assume that $\alpha_{1}\neq 0$ and $\mathbb{E}\left[\varepsilon_{Q,i}|X_{i}^{\ast},S_{i},\varepsilon_{i}\right]=0$. The variance of $\varepsilon_{Q,i}$ does not need to be small. Note that the measurement error in $Q_{i}$ can be non-classical: $Q_{i}-X_{i}^{\ast}$ and $X_{i}^{\ast}$ are correlated unless $\alpha_{1}=1$.
Consider the conditional moment restrictions (15). If $X_{i}^{\ast}$ were observed, we could have constructed the unconditional moments
$\mathbb{E}\left[h\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\right]=0,\text{\qquad}h\left(x,s,\theta\right)\equiv u\left(x,s,\theta\right)\times\left(1,x,\ldots,x^{J}\right)^{\prime},$
for some $J\geq\dim\left(\theta\right)-1$. Suppose that the model is identified if $X_{i}^{\ast}$ observed, which means that the Jacobian of these moment conditions has full rank:
$\mathop{\mathrm{R}k}\left(H^{\ast}\right)=\dim\left(\theta\right)\text{, where }H^{\ast}\equiv E\left[\nabla_{\theta}h\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\right].$
To deal with the error-laden $X_{i}$, consider the MERM estimator with $K=2$ based on the following moment function
$g\left(x,s,q,\theta\right)\equiv\left(\begin{array}[]{l}h\left(x,s,\theta\right)\\ u\left(x,s,\theta\right)q\times\left(1,x,\ldots,x^{J-1}\right)^{\prime}\end{array}\right).$ (16)
Here the total number of moments is $m=2J+1$. The additional $J$ moments added in equation (16) use $Q_{i}$, which will allow identifying $\gamma_{0}=\mathbb{E}[\varepsilon_{i}^{2}]/2$.
It turns out that in these settings there is a simple sufficient condition for Assumption 2.3(i) to hold. Appendix D demonstrates that $\Psi^{\ast}$ will have full rank if
$\mathbb{E}\left[u_{x}^{\left(1\right)}\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\times\left(1,X_{i}^{\ast},\ldots,\left(X_{i}^{\ast}\right)^{J-1}\right)^{\prime}\right]\neq 0.$ (17)
Condition (17) has a very simple interpretation: it essentially it means that $\mathbb{E}\left[\left.u_{x}^{\left(1\right)}\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\right|X_{i}^{\ast}\right]$ should not be identically zero. For example, in the nonlinear regression model $u\left(x,y,\theta\right)=\rho\left(x,\theta\right)-y$, and condition (17) is satisfied as long as $\mathbb{E}\left[\rho_{x}^{(1)}\left(X_{i}^{\ast},\theta_{0}\right)\left(X_{i}^{\ast}\right)^{j}\right]\neq 0$ for some $j\in\left\{0,\ldots,J-1\right\}$.
2.5 Assessing model validity
The standard $J$-test of overidentifying restrictions remains valid in the MERM settings, and can be used to check the model specification. The $J$-test jointly tests the following hypotheses: (i) $K$ is sufficiently large to correct the EIV bias; (ii) assumptions on the EIV are valid; and (iii) the original moment conditions $g$ are correctly specified so equation (1) holds, i.e., that the original economic model is correctly specified aside from the presence of the EIV in $X_{i}$. Issue (i) can be address by taking a larger $K$. In Section 4 we extend the framework to obtain corrected moments that are valid under weaker assumptions on the EIV, which can help addressing issue (ii).
It is possible to obtain an explicit bound on the higher-order EIV bias of the MERM estimator. Suppose the researcher is interested in estimating a linear combination of parameters $v^{\prime}\beta_{0}$ for some non-zero $v\in\mathbb{R}^{\dim(\beta)}$, e.g., the first component of $\theta_{0}$. Since $\hat{\beta}$ is a standard GMM estimator,
$\text{Bias}(v^{\prime}\hat{\beta})\approx-v^{\prime}(\Psi^{\prime}\Xi\Psi)^{-1}\Psi^{\prime}\Xi\;\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right].$ (18)
Here $\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right]\approx\gamma_{0\overline{K}}\mathbb{E}[g_{x}^{(\overline{K})}(X_{i},S_{i},\theta_{0})]$ with $\overline{K}=K+2$ if $K$ is even and $\varepsilon_{i}$ is symmetric, and $\overline{K}=K+1$ otherwise. Thus,
$\text{Bias}(v^{\prime}\hat{\beta})\approx\gamma_{0\overline{K}}b_{v},$
where $b_{v}$ can be easily estimated. The value of $\gamma_{0\overline{K}}$ is given by the recursive relationship (12) and determined by the unknown $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ (and by the lower order moments of $\varepsilon_{i}$, which can be recovered from $\hat{\gamma}$ with a sufficient precision). While $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ and, hence, $\gamma_{0\overline{K}}$ are generally unknown, one simple way to evaluate the magnitude of $\text{Bias}(v^{\prime}\hat{\beta})$ is to calculate $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ using the estimated $\sigma_{\varepsilon}^{2}$ and assuming normality of $\varepsilon$. Alternatively, one can also bound $\gamma_{0\overline{K}}$ and, subsequently, $\text{Bias}(v^{\prime}\hat{\beta})$ by imposing some (a priori) bounds on $\mathbb{E}[(\varepsilon_{i}/\sigma_{\varepsilon})^{\overline{K}}]$.
The bounds on Bias$(v^{\prime}\hat{\beta})$ can also be used to assess the appropriateness of Assumption 2.1 and the reliability of the inference based on the normal approximation in equation (14). If the (worst case) bias of $v^{\prime}\hat{\beta}$ is substantial, relative to its standard error, the researcher may want to account for this and consider choosing a higher $K$. Appendix E provides a further discussion of this issue and additional details on the calculation of Bias$(v^{\prime}\hat{\beta})$.
3 Numerical Evidence
3.1 Comparison with a Semi-Nonparametric Estimation Approach
We compare MERM estimator with the state-of-the-art semiparametric estimator of Schennach (2007, henceforth S07) for nonlinear regression models. The Monte Carlo designs are taken from S07, and include a polynomial, rational fraction, and Probit nonlinear regression models. Identification of the model is ensured by the availability of an instrument.
$Y_{i}=\rho(X_{i}^{*},\theta_{0})+U_{i},\quad X_{i}^{*}=\pi_{1}Z_{i}+V_{i},\quad X_{i}=X_{i}^{*}+\varepsilon_{i},$ (19)
$(Z_{i},V_{i},\varepsilon_{i})^{\prime}\sim N\left((0,0,0)^{\prime},\text{Diag}(1,1/4,1/4)\right)$, $\pi_{1}=1$, and $n=1000$. The conditional expectation function $\rho$, the true value of the parameter of interest $\theta_{0}$, and the conditional distribution of the regression error $U_{i}$ are design-specific and reported in Tables 1-3 below. In all designs, $\tau=\sigma_{\varepsilon}/\sigma_{X}^{*}\approx 0.45$, so the measurement error is “fairly large” (Schennach, 2007).
We report simulation results for the MERM estimator considering correction schemes with $K=2$ and $K=4$. The original moment function is
$g(x,y,z,\theta)=(y-\rho(x,\theta))\varphi(x,z),$
where we use $\varphi(x,z)=\left(1,x,z,x^{2},z^{2},x^{3},z^{3}\right)^{\prime}$ for $K=2$ and $\varphi(x,z)=\left(1,x,z,x^{2},xz,z^{2},x^{3},x^{2}z,xz^{2},z^{3}\right)^{\prime}$ for $K=4$.
The finite sample properties of the MERM estimators (evaluated based on 5,000 replications) are reported in Tables 1-3 below. For comparison, we also provide the same statistics for naive estimators (OLS/NLLS) and for the benchmark estimator of S07 (as reported in the original paper). For the polynomial model (Table 1), both $K=2$ and $K=4$ MERM estimators effectively remove the EIV bias. Component-wise, the MERM estimators perform similarly (for $\theta_{2}$ and $\theta_{4}$) or better (for $\theta_{1}$ and $\theta_{3}$) compared to the benchmark estimator of S07. For the rational fraction model (Table 2), both the MERM estimators are vastly superior to the benchmark estimator both in terms of the bias and the standard deviation. For the probit model (Table 3), the MERM estimator with $K=2$ removes a large fraction of the EIV bias compared to the NLLS estimator. However, the EIV bias remains non-negligible when this simplest correction scheme is used. Employing a higher order correction scheme with $K=4$ completely eliminates the remaining EIV bias, while at the same time having smaller standard deviations (than the benchmark estimator of S07) . Overall, in the considered designs, the MERM estimator with $K=4$ consistently outperforms the benchmark estimator. It also proves to be more effective in removing the EIV bias compared to the $K=2$ estimator, especially in the highly nonlinear settings of the considered probit design.
Table 1: Simulation results for the polynomial model of S07
Bias Std. Dev. RMSE
$\theta_{1}$ $\theta_{2}$ $\theta_{3}$ $\theta_{4}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ $\theta_{4}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ $\theta_{4}$ All
OLS -0.00 -0.43 0.00 0.21 0.07 0.13 0.06 0.04 0.07 0.45 0.06 0.22 0.51
S07 -0.05 -0.07 -0.02 0.05 0.17 0.19 0.24 0.05 0.17 0.20 0.24 0.07 0.36
$K=2$0.00 0.10 0.00 0.00 0.10 0.23 0.10 0.08 0.10 0.25 0.10 0.08 0.29
$K=4$0.00 0.00 0.00 0.02 0.09 0.21 0.10 0.08 0.09 0.21 0.10 0.08 0.27
  • The DGP is as in (19) with $\rho(x,\theta)=\theta_{1}+\theta_{2}x+\theta_{3}x^{2}+\theta_{4}x^{3}$, $\theta_{0}=(1,1,0,-0.5)^{\prime}$, and $U_{i}\sim N(0,1/4)$.
Table 2: Simulation results for the rational fraction model of S07
Bias Std. Dev. RMSE
$\theta_{1}$ $\theta_{2}$ $\theta_{3}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ All
OLS 0.339 -0.167 -0.644 0.040 0.020 0.076 0.341 0.168 0.648 0.752
S07 0.107 0.117 -0.150 0.146 0.139 0.328 0.181 0.182 0.361 0.443
$K=2$0.004 -0.018 0.014 0.062 0.026 0.139 0.062 0.032 0.139 0.156
$K=4$ 0.014 -0.002 -0.024 0.062 0.031 0.154 0.063 0.031 0.156 0.171
  • The DGP is as in (19) with $\rho(x,\theta)=\theta_{1}+\theta_{2}x+\frac{\theta_{3}}{(1+x^{2})^{2}}$, $\theta_{0}=(1,1,2)^{\prime}$, and $U_{i}\sim N(0,1/4)$.
Table 3: Simulation results for the Probit model of S07
Bias Std. Dev. RMSE
$\theta_{1}$ $\theta_{2}$ $\theta_{1}$ $\theta_{2}$ $\theta_{1}$ $\theta_{2}$ All
NLLS 0.38 -0.97 0.06 0.08 0.39 0.98 1.05
S07 0.05 -0.06 0.39 0.53 0.39 0.53 0.69
$K=2$ 0.11 -0.31 0.18 0.34 0.21 0.46 0.51
$K=4$0.01 -0.01 0.23 0.42 0.23 0.42 0.48
  • The DGP is as in (19) with $\rho(x,\theta)=\frac{1}{2}(1+\text{erf}(\theta_{1}+\theta_{2}x))$, $\theta_{0}=(-1,2)^{\prime}$, and $U_{i}=1-\rho(X_{i}^{*},\theta_{0})$ with probability $\rho(X_{i}^{*},\theta_{0})$ and $-\rho(X_{i}^{*},\theta_{0})$ otherwise.
3.2 Estimation and Inference in a Multinomial Choice Model
Consider the standard multinomial logit model, in which an agent chooses between 3 available options. For an agent $i$ with characteristics $(X_{i}^{*},W_{i})$, the utility of option $j$ is given by
$U_{ij}=\theta_{0j1}X_{i}^{*}+\theta_{0j2}W_{ij}+\theta_{0j3}+\epsilon_{ij}\quad\text{for }j\in\{1,2\},$
and $U_{i0}=\epsilon_{i0}$ for the outside option $j=0$, where $\epsilon_{ij}$ are i.i.d. (across $i$ and $j$) draws from a standard type-1 extreme value distribution. The researcher observes $\left\{(X_{i},W_{i},Y_{i1},Y_{i2},Y_{i0})\right\}_{i=1}^{n}$, where $Y_{ij}$ is a binary variable indicating whether agent $i$ chooses option $j$, i.e. $Y_{ij}=1$ if and only if $j=\operatorname*{\mathrm{arg}\!\max\limits}_{j^{\prime}\in\{0,1,2\}}U_{ij^{\prime}}$. In addition,
$X_{i}^{*}=V_{i1}Z_{i}+V_{i0},\quad X_{i}=X_{i}^{*}+\varepsilon_{i},\quad W_{ij}={\rho}X_{i}^{*}/\sigma_{X}^{*}+\sqrt{1-\rho^{2}}\nu_{ij},$
and $\left(V_{i1},V_{i0},Z_{i},\varepsilon_{i},\nu_{i1},\nu_{i2}\right)^{\prime}\sim N\left((1,0,0,0,0,0)^{\prime},\text{Diag}(\sigma_{V1}^{2},\sigma_{V0}^{2},\sigma_{Z}^{2},\sigma_{\varepsilon}^{2},\sigma_{\nu}^{2},\sigma_{\nu}^{2})\right)$. In all of the designs, we fix $(\theta_{011},\theta_{012},\theta_{013},\theta_{021},\theta_{022},\theta_{023},\rho,\sigma_{V1}^{2},\sigma_{V0}^{2},\sigma_{Z}^{2},\sigma_{\nu}^{2})=(1,0,0,0,0,0,0.7,1/2,1/2,1,1)$ and $n=2000$. We consider $\tau=\sigma_{\varepsilon}/\sigma_{X^{*}}\in\{1/4,1/2,3/4\}$. Setting $\sigma_{V1}=0$ would correspond to the additive control variable model. We omit such simulation results for brevity.
Similarly to Section 3.1, we report results for the MERM estimators with $K=2$ and $K=4$ based on the following original moment function
$g(x,w,y,z,\theta)$ $=\left(\left(y_{1}-p_{1}(x,w,\theta)\right)\varphi_{1}(x,z,w)^{\prime},\left(y_{2}-p_{2}(x,w,\theta)\right)\varphi_{2}(x,z,w)^{\prime}\right)^{\prime},$
$p_{j}(x,w,\theta)$ $=\frac{\exp(\theta_{j1}x+\theta_{j2}w_{j}+\theta_{j3})}{1+\exp(\theta_{11}x+\theta_{12}w_{1}+\theta_{13})+\exp(\theta_{21}x+\theta_{22}w_{2}+\theta_{23})},$
where $\varphi_{j}(x,z,w)=\left(1,x,z,x^{2},z^{2},x^{3},z^{3},w_{j}\right)^{\prime}$ for $K=2$ and $\varphi_{j}(x,z,w)=\left(1,x,z,x^{2},xz,z^{2},x^{3},x^{2}z,xz^{2},z^{3},w_{j}\right)^{\prime}$ for $K=4$.
We report the results on estimation and inference on the partial derivatives of the conditional choice probabilities $p_{j}(x,w_{1},w_{2})$ with respect to $x$, $w_{1}$, and $w_{2}$, evaluated at the population means.
Table LABEL:tab:_MC_mult_logit_2 reports the finite sample biases, standard deviations, and RMSE of the MERM estimators, as well as the sizes of the corresponding t-tests with nominal size of 5%. To illustrate the importance of dealing with EIV, we also report the same statistics for the standard (naive) MLE estimator that ignores the presence of the measurement errors.
In all designs, the MLE estimator is biased, and the corresponding t-tests over-reject. Note that failing to account for the EIV in the mismeasured variable $X_{i}^{*}$ generally biases estimators of all of the parameter, including those corresponding to the correctly measured variables $W_{i1}$ and $W_{i2}$. In particular, the t-tests may falsely reject true null hypotheses $\partial p_{j}/\partial w_{\ell}=0$ up to nearly $100\%$ of the time.
The MERM estimator with $K=2$ removes a large fraction of the EIV bias in all of the designs. While this proves to be enough to achieve accurate size control when the magnitude of the measurement error is moderate ($\tau=1/4$), the remaining EIV bias may still result in size distortions of the t-tests with larger measurement errors, especially $\tau=3/4$. Using the higher order correction scheme with $K=4$ effectively removes the EIV bias in all of the simulation designs for all of the parameters. Remarkably, the corresponding finite sample null rejection probabilities remain close to the nominal $5\%$ rate even when the standard deviation of the measurement error is as large as $75\%$ of the standard deviation of the mismeasured $X^{*}$.
3.3 Empirical Illustration: Choice of Transportation Mode
In this section, we illustrate the finite sample properties of the MERM estimator in the context of a classical multinomial choice application: choice of transportation mode (e.g., McFadden, 1974).
To calibrate the numerical experiment, we use the ModeCanada dataset, a survey of business travelers for the Montreal-Toronto corridor. We focus on the subset of travelers choosing between train, air, and car ($n=2769$), and estimate the conditional logit model with traveler $i$’s utilities given in the table below.
Mode Utility
Air $U_{i1}=\theta_{01}\thinspace Income_{i}^{*}+\theta_{02}\thinspace Urban_{i}+\theta_{03}+\theta_{07}\thinspace Price_{i1}+\theta_{08}\thinspace InTime_{i1}+\epsilon_{i1}$
Car $U_{i2}=\theta_{04}\thinspace Income_{i}^{*}+\theta_{05}\thinspace Urban_{i}+\theta_{06}+\theta_{07}\thinspace Price_{i2}+\theta_{08}\thinspace InTime_{i2}+\epsilon_{i2}$
Train $U_{i0}=\theta_{07}\thinspace Price_{i0}+\theta_{08}\thinspace InTime_{i0}+\epsilon_{i0}$
To generate the simulated samples, we randomly draw covariates from their joint empirical distribution. To generate the simulated outcomes, we draw $\epsilon_{ij}$ from the standard type-I extreme value distribution. The true value of $\theta_{0}$ is set to be the MLE estimate based on the original dataset. More details about this numerical experiment are given in Appendix H.
To evaluate the performance of the MERM estimator in these settings, we generate mismeasured $Income_{i}=Income_{i}^{*}+\varepsilon_{i}$. We focus on the individual income because it is often mismeasured.We report the results for $\tau=\sigma_{\varepsilon}/\sigma_{Income^{*}}\in\{1/4,1/2,3/4\}$.
Table 4 reports the simulation results for the (naive) MLE estimator and for the MERM estimators with $K=2$ and $K=4$. We focus on estimation of and inference on the income elasticities (evaluated at the population mean of the covariates). The MLE estimator is considerably biased for $\tau\in\{1/2,3/4\}$, which results in substantial size distortions of the MLE based t-tests. The MERM estimator with $K=4$ effectively eliminates the EIV bias and the corresponding t-tests provide accurate size control in all of the considered designs. The estimator with $K=2$ is more precise, while successfully removing the EIV bias for $\tau\leq 1/2$.
Overall, the MERM estimators perform well in the considered empirical context, providing a basis for estimation and inference even for quite large values of $\tau$.
Table 4: Simulation results for the empirically calibrated conditional logit model
MLE $K=2$ $K=4$
bias std rmse size bias std rmse size bias std rmse size
$\tau=1/4$
$\partial\ln p_{1}/\partial\ln I$0.07 0.12 0.14 9.00 0.01 0.14 0.14 5.68 0.02 0.19 0.19 7.02
$\partial\ln p_{2}/\partial\ln I$ 0.03 0.07 0.08 5.84 -0.00 0.08 0.08 5.68 -0.01 0.10 0.10 6.40
$\partial\ln p_{0}/\partial\ln I$ 0.05 0.13 0.13 6.10 0.00 0.14 0.14 5.42 -0.00 0.17 0.17 7.66
$\tau=1/2$
$\partial\ln p_{1}/\partial\ln I$0.24 0.11 0.27 61.84 -0.05 0.14 0.15 6.96 0.02 0.21 0.21 6.16
$\partial\ln p_{2}/\partial\ln I$ 0.09 0.07 0.11 24.76 0.02 0.09 0.09 5.96 -0.01 0.10 0.10 6.16
$\partial\ln p_{0}/\partial\ln I$ 0.16 0.12 0.20 25.36 0.04 0.15 0.15 6.06 -0.00 0.18 0.18 6.86
$\tau=3/4$
$\partial\ln p_{1}/\partial\ln I$0.43 0.09 0.44 99.50 -0.19 0.14 0.24 27.46 0.02 0.22 0.22 5.84
$\partial\ln p_{2}/\partial\ln I$ 0.16 0.06 0.17 71.88 0.07 0.08 0.11 13.78 -0.01 0.11 0.11 6.32
$\partial\ln p_{0}/\partial\ln I$ 0.29 0.11 0.31 73.20 0.12 0.15 0.19 13.40 0.00 0.19 0.19 6.32
  • This table reports the simulated finite sample bias, standard deviation, RMSE, and size of the MLE and the MERM estimators and the corresponding t-tests for the income elasticities $\partial\ln p_{j}(I,w,\theta_{0})/\partial\ln I$, $j\in\{1,2,0\}$, evaluated at the population mean. The true values of the income elasticities are $(\partial\ln p_{1}/\partial\ln I,\partial\ln p_{2}/\partial\ln I,\partial\ln p_{0}/\partial\ln I)=(1.11,-0.39,-0.82)$. The results are based on 5,000 replications.
4 Extensions
4.1 Multiple Mismeasured Variables
It is easy to use the MERM framework to deal with multiple mismeasured variables. This is useful in many applications, including not only settings with multiple mismeasured covariates, but also settings with serially correlated measurement errors, settings where repeated measurements are available, and panel data models. Using the MERM approach is particularly advantageous in such applications, since it avoids nonparametric estimation of multivariate unobserved distributions.
Suppose $X_{i}^{*}$, $\varepsilon_{i}$, and $X_{i}$ are $d\times 1$ vectors. Let $\tau_{n}\equiv\max_{j\leq d}\sigma_{\varepsilon_{j}}/\sigma_{X^{*}_{j}}$, where $\sigma_{\varepsilon_{j}}$ and $\sigma_{X^{*}_{j}}$ denote the standard deviations of the $j$-th components of $\varepsilon_{i}$ and $X_{i}^{*}$, so $\mathbb{E}\left[\left|\varepsilon_{ij}\right|^{k}\right]=O(\tau_{n}^{k})$ for $k\in\{1,\ldots,K\}$.
For a $d\times 1$ vector of non-negative integers $\kappa=(\kappa_{1},\ldots,\kappa_{d})\in\mathbb{Z}_{+}^{d}$, let
$\partial_{\kappa}\equiv\frac{\partial^{\left|\kappa\right|}}{\partial{x_{1}}^{\kappa_{1}}\ldots\partial{x_{d}}^{\kappa_{d}}},\quad\text{where }\left|\kappa\right|\equiv\sum_{j=1}^{d}\kappa_{j}.$
Also, for a positive integer $k$, let $\mathcal{K}_{k}=\{\kappa\in\mathbb{Z}_{+}^{d}:\left|\kappa\right|=k\}$. Then, we consider the following corrected moment function
$\psi(x,s,\theta,\gamma)=g(x,s,\theta)-\sum_{k=2}^{K}\sum_{\kappa\in\mathcal{K}_{k}}\gamma_{\kappa}\partial_{\kappa}g(x,s,\theta),$
where, with some abuse of notation, $\gamma$ is a collection of all $\gamma_{\kappa}$ with $\kappa\in\mathcal{K}_{k}$ and $k\in\{2,\dots,K\}$.
Under mild smoothness conditions
$\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0})\right]=\mathbb{E}\left[g(X_{i}^{*},S_{i},\theta_{0})\right]+O(\tau_{n}^{K+1})=o(n^{-1/2}),$
where the second equality holds provided that $O(\tau_{n}^{K+1})=o(n^{-1/2})$. Similarly to the scalar case, components of $\gamma_{0}$ are determined by the moments of $\varepsilon_{i}$. Specifically, let $\mu_{\kappa}\equiv\mathbb{E}\left[\varepsilon_{i1}^{\kappa_{1}}\dots\varepsilon_{id}^{\kappa_{d}}\right]$, then
$\gamma_{0\kappa}=\frac{\mu_{\kappa}}{\kappa!},\quad\text{for }\kappa\in\{\mathcal{K}_{2},\mathcal{K}_{3}\},$ (20)
where $\kappa!\equiv\kappa_{1}!\ldots\kappa_{d}!$. For $\left|\kappa\right|\geq 4$, the coefficients can be computed by the following formulas. For example, for $\kappa\in\mathcal{K}_{4}$, let $\mathcal{K}_{2,\kappa}=\{\tilde{\kappa}\in\mathcal{K}_{2}:\kappa-\tilde{\kappa}\in\mathcal{K}_{2}\}$. Then,
$\gamma_{0\kappa}=\frac{\mu_{\kappa}}{\kappa!}-\sum_{\tilde{\kappa}\in\mathcal{K}_{2,\kappa}}\frac{\mu_{\kappa-\tilde{\kappa}}}{(\kappa-\tilde{\kappa})!}\gamma_{0\tilde{\kappa}},\quad\text{for }\kappa\in\mathcal{K}_{4}.$
More generally, for $\kappa\in\mathcal{K}_{k}$ with $k\geq 4$, let $\mathcal{K}_{\ell,\kappa}=\{\tilde{\kappa}\in\mathcal{K}_{\ell},\kappa-\tilde{\kappa}\in\mathcal{K}_{\left|\kappa\right|-\ell}\}$ for $\ell\leq\left|\kappa\right|-2$. Then,
$\gamma_{0\kappa}=\frac{\mu_{\kappa}}{\kappa!}-\sum_{\ell=2}^{k-2}\sum_{\tilde{\kappa}\in\mathcal{K}_{\ell,\kappa}}\frac{\mu_{\kappa-\tilde{\kappa}}}{(\kappa-\tilde{\kappa})!}\gamma_{0\tilde{\kappa}}.$
Example (Bivariate $X$, $K=4$).
Suppose $X$ is bivariate (i.e., $d=2$) and $K=4$. For $\kappa\in\mathcal{K}_{2}=\{(2,0),(1,1),(0,2)\}$ and $\kappa\in\mathcal{K}_{3}=\{(3,0),(2,1),(1,2),(0,3)\}$, $\gamma_{0\kappa}$ is given by (20). For $\kappa\in\mathcal{K}_{4}$, $\gamma_{0\kappa}$ is given by
$\kappa$ $\gamma_{0\kappa}$
(4,0) $\left(\mathbb{E}[{\varepsilon_{i1}^{4}}]-6\mathbb{E}[{\varepsilon_{i1}^{2}}]^{2}\right)/24$
(3,1) $\left(\mathbb{E}[{\varepsilon_{i1}^{3}\varepsilon_{i2}}]-6\mathbb{E}[{\varepsilon_{i1}^{2}}]\mathbb{E}[{\varepsilon_{i1}\varepsilon_{i2}}]\right)/6$
(2,2) $\left(\mathbb{E}[{\varepsilon_{i1}^{2}\varepsilon_{i2}^{2}}]-2\mathbb{E}[{\varepsilon_{i1}^{2}}]\mathbb{E}[{\varepsilon_{i2}^{2}}]-4\mathbb{E}[{\varepsilon_{i1}\varepsilon_{i2}}]^{2}\right)/4$
(1,3) $\left(\mathbb{E}[{\varepsilon_{i1}\varepsilon_{i2}^{3}}]-6\mathbb{E}[{\varepsilon_{i2}^{2}}]\mathbb{E}[{\varepsilon_{i1}\varepsilon_{i2}}]\right)/6$
(0,4) $\left(\mathbb{E}[{\varepsilon_{i2}^{4}}]-6\mathbb{E}[{\varepsilon_{i2}^{2}}]^{2}\right)/24$
If in addition measurement errors $\varepsilon_{i1}$ and $\varepsilon_{i2}$ are independent, $\gamma_{0\kappa}=0$ for $\kappa\in\{(1,1),(2,1),(1,2),(3,1),(1,3)\}$. In this case, the total number of the nuisance parameters to be estimated is 6. $\blacksquare$
4.2 Weakly Classical and Non-Classical Measurement Errors
In some applications, assuming that the measurement error is classical, i.e., $\varepsilon_{i}$ is independent from $(X_{i}^{*},S_{i})$, may be too restrictive. In this section, we demonstrate how the MERM framework can be used to address this issue. To address general non-classical measurement errors it is convenient to proceed in two steps.
First, in Section 4.2.1, we consider estimation of models with weakly classical measurement errors satisfying $\mathbb{E}\left[\varepsilon_{i}|X_{i}^{\ast},S_{i}\right]=0$. This restriction on $\varepsilon_{i}$ is much weaker than the classical measurement error assumption. For example, the variance of the measurement error $\mathbb{E}\left[\varepsilon_{i}^{2}|X_{i}^{\ast},S_{i}\right]$ (and other moments) could depend on $X_{i}^{\ast}$ and $S_{i}$.
Second, in Section 4.2.2, we consider estimation with general non-classical measurement errors that may be correlated with the mismeasured variable.
4.2.1 Weakly Classical Measurement Errors
In this section, we relax Assumption 2.2 and assume that the measurement error is weakly classical. Specifically, assume that $\mathbb{E}\left[\varepsilon_{i}|X_{i}^{*},S_{i}\right]=0$ or, equivalently, $\mathbb{E}\left[X_{i}|X_{i}^{*},S_{i}\right]=X_{i}^{*}$.
Suppose that $\mathbb{E}\left[\left|\varepsilon_{i}\right|^{k}|X_{i}^{*},S_{i}\right]=O(\tau_{n}^{k})$ for $k\in\{2,\ldots,K\}$. Conditional on $X_{i}^{*}$ and $S_{i}$, the moment function can be expanded as
$\mathbb{E}[g(X_{i},S_{i},\theta)|X_{i}^{\ast},S_{i}]={g(X_{i}^{\ast},S_{i},\theta)}+\sum_{k=2}^{K}\frac{\mathbb{E}\left[\varepsilon_{i}^{k}|X_{i}^{*},S_{i}\right]}{k!}{g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)}+O(\tau_{n}^{K+1}).$ (21)
Suppose $\mathbb{E}\left[\varepsilon_{i}^{k}|X_{i}^{*},S_{i}\right]=v_{k}(X_{i}^{*},S_{i},\omega_{0})$, $k\in\{2,\ldots,K\}$ where $\omega_{0}\in\mathbb{R}^{\dim(\omega)}$ are unknown parameters. Typically, functions $v_{k}$ are assumed to depend only on some of the components of $(X_{i}^{*},S_{i})$. For example, in the nonlinear regression (2), it may be natural for $v_{k}$ to depend on $X_{i}^{*}$ and/or $W_{i}$, but not on $Y_{i}$. Equation (21) then implies that
$\mathbb{E}\left[g(X_{i}^{*},S_{i},\theta)\right]=\mathbb{E}\left[g(X_{i},S_{i},\theta)-\sum_{k=2}^{K}\frac{v_{k}(X_{i}^{*},S_{i},\omega_{0})}{k!}{g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)}\right]+O(\tau_{n}^{K+1}).$ (22)
Equation (22) is an analog of equation (10), which motivates correcting the moment conditions by estimating and subtracting the bias terms $\frac{v_{k}(X_{i}^{*},S_{i},\omega_{0})}{k!}{g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)}$. The corrected moment function takes the form
$\psi(X_{i},S_{i},\theta,\omega)=g(X_{i},S_{i},\theta)-\sum_{k=2}^{K}f_{k}(X_{i},S_{i},\theta,\omega),$ (23)
where
$f_{2}(x,s,\theta,\omega)=$ $\frac{v_{2}(x,s,\omega)}{2}g_{x}^{(2)}(x,s,\theta),\quad f_{3}(x,s,\theta,\omega)=\frac{v_{3}(x,s,\omega)}{6}g_{x}^{(3)}(x,s,\theta),$
$f_{k}(x,s,\theta,\omega)=$ $\frac{v_{k}(x,s,\omega)}{k!}g_{x}^{(k)}(x,s,\theta)-\sum_{\ell=2}^{k-2}\frac{v_{k-\ell}(x,s,\omega)}{(k-\ell)!}\frac{\partial^{k-\ell}}{\partial x^{k-\ell}}f_{\ell}(x,s,\theta,\omega),\text{ for $k\geq 4$}.$
Notice that for $K\leq 3$ the corrected moment functions are similar to the classical EIV case except for $v_{2}(x,s,\omega)$ and $v_{3}(x,s,\omega)$ taking the places of $\mathbb{E}\left[\varepsilon_{i}^{2}\right]$ and $\mathbb{E}\left[\varepsilon_{i}^{3}\right]$ respectively. For $K\geq 4$, we need to bias correct the bias correction terms such as $f_{2}(x,s,\theta,\omega)$ accounting for $v_{2}(X_{i},S_{i},\theta,\omega)$ being evaluated at $X_{i}$ instead of $X_{i}^{*}$. Also, when the conditional moments of $\varepsilon_{i}$ do not depend on $X_{i}^{*}$, the corrected moment function matches the classical EIV case with $v_{k}(s,\omega)$ replacing $\mathbb{E}\left[\varepsilon_{i}^{k}\right]$.
Under the regularity conditions
$\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\omega_{0})\right]=\mathbb{E}\left[g(X_{i}^{*},S_{i},\theta_{0})\right]+O(\tau_{n}^{K+1})=o(n^{-1/2}),$
provided that $O(\tau_{n}^{K+1})=o(n^{-1/2})$.
Thus, the model can be estimated using GMM estimator with the corrected moment function (23), where parameters $\theta_{0}$ and $\omega_{0}$ are estimated together, i.e., the estimator in equation (13) with $\beta\equiv(\theta^{\prime},\omega^{\prime})^{\prime}$. The following example provides a convenient parameterization of functions $v$.
Example 4.1 (Exponential Specification, $K=4$).
Suppose $\varepsilon_{i}=\exp(\omega_{01}X_{i}^{*})\zeta_{i}$, where $\zeta_{i}$ is mean zero and independent of $(X_{i}^{*},S_{i})$. In this case, $\mathbb{E}\left[\varepsilon_{i}^{k}|X_{i}^{*},S\right]=\mathbb{E}\left[\zeta_{i}^{k}\right]\exp(k\omega_{01}X_{i}^{*})$. Then, we can take $v_{k}(x,s,\omega)=\omega_{k}\exp(k\omega_{1}x)$, $k\in\{2,\ldots,4\}$, with $\omega_{0}=(\omega_{01},\mathbb{E}\left[\zeta_{i}^{2}\right],\mathbb{E}\left[\zeta_{i}^{3}\right],\mathbb{E}\left[\zeta_{i}^{4}\right])^{\prime}$. Functions $f_{k}$ in equation (23) are
$f_{2}(x,s,\theta,\omega)$ $=\frac{\omega_{2}}{2}e^{2\omega_{1}x}g_{x}^{(2)}(x,s,\theta),\quad f_{3}(x,s,\theta,\omega)=\frac{\omega_{3}}{6}e^{3\omega_{1}x}g_{x}^{(3)}(x,s,\theta),$
$f_{4}(x,s,\theta,\omega)$ $=\textstyle e^{4\omega_{1}x}\left(\frac{\omega_{4}-6\omega_{2}^{2}}{24}g_{x}^{(4)}(x,s,\theta)-\omega_{1}\omega_{2}^{2}g_{x}^{(3)}(x,s,\theta)-\omega_{1}^{2}\omega_{2}^{2}g_{x}^{(2)}(x,s,\theta)\right).\ \ \hfill\blacksquare$
Identifiability of parameters $\omega_{0}$ depends on the specific model and moment conditions $g$. For the nonlinear regression model with weakly classical EIV Evdokimov and Zeleneev (2022) establish nonparametric identification using instrumental variables.
4.2.2 Non-Classical Measurement Errors
Since $X_{i}^{\ast}$ is not observed, in the absence of some a priori information about its distribution, some assumptions such as $\mathbb{E}\left[\left.X_{i}\right|X_{i}^{\ast},S_{i}\right]=X_{i}^{\ast}$ appear to be necessary, and are routinely imposed in the EIV literature. For example, see Hu and Schennach (2008), Hu and Sasaki (2015), Freyberger (2021), and Schennach (2021) for recent discussions of such assumptions. In the absence of such assumptions one cannot determine even the scale of the true unobserved covariate.
To be specific, suppose the true covariate is denoted by $\mathcal{X}_{i}^{\ast}$ and is mismeasured according to the following model
$X_{i}=\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast}+\mathcal{E}_{i},\qquad\mathbb{E}\left[\left.\mathcal{E}_{i}\right|\mathcal{X}_{i}^{\ast},S_{i}\right]=0,$ (24)
so $\mathbb{E}\left[\left.X_{i}\right|\mathcal{X}_{i}^{\ast}\right]=\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast}$ for some unknown $\alpha_{0}$ and $\alpha_{1}>0$. Note that
$\mathbb{E}\left[\left.X_{i}-\mathcal{X}_{i}^{\ast}\right|\mathcal{X}_{i}^{\ast}\right]\neq 0,$
and the measurement error is non-classical. For concreteness, consider the nonlinear regression model:
$\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast},W_{i}\right]=\rho\left(\theta_{01}+\theta_{0\mathcal{X}}\mathcal{X}_{i}^{\ast}+\theta_{0W}^{\prime}W_{i}\right)$ (25)
for some known function $\rho$.
Since $\mathcal{X}_{i}^{\ast}$ is unobservable, we can write an observationally equivalent model in terms of another unobservable $X_{i}^{\ast}$ defined as
$X_{i}^{\ast}$ $\equiv\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast},$ (26)
$\mathbb{E}\left[Y_{i}|X_{i}^{\ast},W_{i}\right]$ $=\rho\left(\tilde{\theta}_{01}+\tilde{\theta}_{0X}{\large X}_{i}^{\ast}+\theta_{0W}^{\prime}W_{i}\right),$ (27)
where $\tilde{\theta}_{0X}\equiv\theta_{0\mathcal{X}}/\alpha_{1},\quad\tilde{\theta}_{01}\equiv\theta_{01}-\tilde{\theta}_{0X}\alpha_{0}$. Thus, parameters $\left(\theta_{01},\theta_{0\mathcal{X}}\right)$ and $\left(\alpha_{0},\alpha_{1}\right)$ are not separately identified without additional information about the location and scale of $\mathcal{X}_{i}^{\ast}$. Note that neither instrumental variables nor multiple measurements of the form (24) provide such information.
Equation (24) and the definition (26) of $X_{i}^{\ast}$ imply that condition $\mathbb{E}\left[\left.X_{i}\right|X_{i}^{\ast},S_{i}\right]=X_{i}^{\ast}$ holds for the observationally equivalent model (27). This observationally equivalent model can be identified and estimated using the approach of Section 4.2.1. This then allows identification of many parameters of interest in the original model (25) because they do not depend on the (unidentified) scale of $\mathcal{X}_{i}^{\ast}$.
First, parameters $\theta_{0W}$ are the same in both models. Moreover, we have $\left(\mathcal{X}_{i}^{\ast}-\mu_{\mathcal{X}^{\ast}}\right)/\sigma_{\mathcal{X}^{\ast}}=\left(X_{i}^{\ast}-\mu_{X^{\ast}}\right)/\sigma_{X^{\ast}}$, and hence the following average structural function
$m(t,w)\equiv\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast}=\mu_{\mathcal{X}^{\ast}}+t\sigma_{\mathcal{X}^{\ast}},W_{i}=w\right]$
is identified, since $m(t,w)=\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=\mu_{X^{\ast}}+t\sigma_{X^{\ast}},W_{i}=w\right]$ is identified. As a result, we can identify marginal effects such as $\frac{\partial}{\partial w}\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast}=\mu_{\mathcal{X}^{\ast}},W_{i}=w\right]=\frac{\partial}{\partial w}m(0,w)$, or the effect of increasing $\mathcal{X}_{i}^{\ast}$ from $\mu_{\mathcal{X}^{\ast}}$ to $\mu_{\mathcal{X}^{\ast}}+\lambda\sigma_{\mathcal{X}^{\ast}}$, i.e., $m(\lambda,w)-m(0,w)$. The corresponding marginal effects averaged with respect to the distribution of $\left(\mathcal{X}_{i}^{\ast},W_{i}\right)$ are also identified and can be estimated following Remark 3. Thus, in this model, the MERM estimator from Section 4.2.1 allows estimation of many economically meaningful marginal effects, even though the true measurement model is given by equation (24), and the measurement error $X_{i}-\mathcal{X}_{i}^{\ast}$ is non-classical. [^10] In the constructed observationally equivalent model (27) we have $\varepsilon_{i}=\mathcal{E}_{i}$, and the MERM approach is justified when $\sigma_{\mathcal{E}}^{2}/\sigma_{X}^{2}$ is moderate. The true measurement error $X_{i}-\mathcal{X}_{i}^{\ast}$ does not need to be small.
At the same time, identification of the true values of $\theta_{01}$ and $\theta_{0\mathcal{X}}$ or some marginal effects such as $\frac{\partial}{\partial\varkappa}\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast}=\varkappa,W_{i}=w\right]$ requires identifying $\alpha_{0}$ and $\alpha_{1}$.
Using the summary statistics for $\mathcal{X}^{\ast}$
In some applications the researchers may have information about the marginal distribution of $\mathcal{X}^{\ast}$ such as its mean $\mu_{\mathcal{X}^{\ast}}$ and variance $\sigma_{\mathcal{X}^{\ast}}^{2}$, e.g., from an administrative dataset. [^11] It is important that $\mathcal{X}_{j}^{\ast}$ and $\left(X_{i},S_{i}\right)$ are drawn from the same population., [^12] Knowing $\sigma_{\mathcal{X}^{\ast}}^{2}$ does not identify even the unconditional variance $\sigma_{\varepsilon}^{2}$, since $\sigma_{X}^{2}=\alpha_{1}^{2}\sigma_{\mathcal{X}^{\ast}}^{2}+\sigma_{\varepsilon}^{2}$. Then, the MERM approach can be used to estimate $\alpha_{0}$ and $\alpha_{1}$ together with parameters $\theta_{0}$. For any original moment condition $\mathbb{E}\left[g\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)\right]=0$ consider the augmented moment condition $\mathbb{E}\left[g_{A}\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)\right]=0$, where
$g_{A}\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)\equiv\left(g\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)^{\prime},\mathcal{X}_{i}^{\ast}-\mu_{\mathcal{X}^{\ast}},\left(\mathcal{X}_{i}^{\ast}-\mu_{\mathcal{X}^{\ast}}\right)^{2}-\sigma_{\mathcal{X}^{\ast}}^{2}\right)^{\prime}.$
Define $X_{i}^{\ast}\equiv\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast}$, $\tilde{\theta}_{0}\equiv\left(\theta_{0}^{\prime},\alpha_{0},\alpha_{1}\right)^{\prime}$, and $\tilde{g}_{A}(X_{i}^{\ast},S_{i},\tilde{\theta}_{0})\equiv g_{A}\left(\left.\left(X_{i}^{\ast}-\alpha_{0}\right)\right/\alpha_{1},S_{i},\theta_{0}\right)$. Then one can take $\mathbb{E}[\tilde{g}_{A}(X_{i}^{\ast},S_{i},\tilde{\theta}_{0})]=0$ as the new “original” moment condition (1). Since the condition $\mathbb{E}\left[X_{i}|X_{i}^{\ast},S_{i}\right]=X_{i}^{\ast}$ holds, $\tilde{\theta}_{0}$ can be estimated using the MERM approach and the corrected moments (23) applied to $\tilde{g}_{A}$ in place of $g$.
Evdokimov and Zeleneev (2022) show how an identification argument based on the assumption of weakly classical measurement errors can be used for nonparametric identification in the settings with non-classical measurement errors. The linearity of equation (24) simplifies the estimation procedure, but the underlying approach is valid nonparametrically. Using a parsimonious model, this section provides a practical way of dealing with non-classical measurement errors in empirical applications, avoiding any nonparametric estimation.
Appendix A Regularity Conditions
Notation. Let $\mathcal{X}\subseteq\mathbb{R}$ be some closed convex set containing the union of the supports of $X_{i}^{*}$ and $X_{i}$, and $\mathcal{S}=\text{supp}\left(S_{i}\right)$.
Assumption A.1.
(Moment function) Suppose that the moment restrictions (1) are satisfied and the following conditions hold:
  1. (i)
    For all $s\in\mathcal{S}$ and $\theta\in\Theta$, $g_{x}^{(K)}(x,s,\theta)$ exists and is continuous on $\mathcal{X}$. Moreover, there exist functions $b_{1},b_{2}:\mathcal{X}\times\mathcal{S}\times\Theta\rightarrow\mathbb{R}_{+}$ and integer $M\geqslant K+1$ such that for all $x,x^{\prime}\in\mathcal{X}$, $s\in\mathcal{S}$, and $\theta\in\Theta$,
    $\left\|g^{(K)}_{x}(x^{\prime},s,\theta)-g^{(K)}_{x}(x,s,\theta)\right\|\leqslant b_{1}(x,s,\theta)|x^{\prime}-x|+b_{2}(x,s,\theta)|x^{\prime}-x|^{M-K};$ (A.1)
  2. (ii)
    Assumption 2.1 holds with $L\geqslant M$;
  3. (iii)
    $\mathbb{E}\left[g_{x}^{(k)}(X_{i}^{*},S_{i},\theta_{0})\right]$, $k\in\{1,\dots,K\}$, and $\mathbb{E}\left[b_{j}(X_{i}^{*},S_{i},\theta_{0})\right]$, $j\in\{1,2\}$, exist and are bounded.
Assumption A.1 allows us to bound the remainder of the Taylor expansion of $g(X_{i},S_{i},\theta)$ around $X_{i}^{*}$ by a polynomial in $\left|X_{i}-X_{i}^{*}\right|=\left|\varepsilon_{i}\right|$. Combined with Assumption 2.1 (which bounds the moments of $\varepsilon_{i}$), it ensures that this remainder is $o(n^{-1/2})$, which is crucial for establishing validity of the corrected moment function $\psi$ (Lemma 1).
Notice that if $\mathcal{X}$ is compact, condition (A.1) is satisfied if $g_{x}^{(K+1)}(x,s,\theta)$ is bounded on $\mathcal{X}$ (for all $s\in\mathcal{S}$ and $\theta\in\Theta$). If $\mathcal{X}$ is unbounded, condition (A.1) is satisfied if for some $J$, such that $K<J\leqslant M$, $\sup_{x\in\mathcal{X}}\left\|g_{x}^{(J)}(x,s,\theta)\right\|\leqslant B(s,\theta)$ for some function $B(s,\theta)$. Also notice that condition (A.1) is stronger than the standard Lipschitz continuity because in applications $\left\|g_{x}^{(K)}(x,s,\theta)\right\|$ may behave like a polynomial in $x$ for large $x$.
Assumption A.2.
(Parameter space)
  1. (i)
    $\Theta\subset\mathbb{R}^{\dim(\theta)}$ and $\Gamma\subset\mathbb{R}^{K-1}$ are compact, $\theta_{0}\in\text{int}\left(\Theta\right)$ and $\gamma_{0n}\in\Gamma$;
  2. (ii)
    $0_{K-1}\in\text{int}\left(\Gamma\right)$.
Assumption A.3.
(Regularity and smoothness conditions)
  1. (i)
    For all $s\in\mathcal{S}$, $G_{x}^{(K)}(x,s,\theta)$ exists and is continuous on $\mathcal{X}\times\Theta$; moreover, there exist functions $b_{G1},b_{G2}:\mathcal{X}\times\mathcal{S}\times\Theta\rightarrow\mathbb{R}_{+}$ and $\delta>0$ and for all $x,x^{\prime}\in\mathcal{X}$, $s\in\mathcal{S}$, and $\theta\in B_{\delta}(\theta_{0})$
    $\left\|G^{(K)}_{x}(x^{\prime},s,\theta)-G^{(K)}_{x}(x,s,\theta)\right\|\leqslant b_{G1}(x,s,\theta)|x^{\prime}-x|+b_{G2}(x,s,\theta)|x^{\prime}-x|^{M-K}$
  2. (ii)
    $\mathbb{E}\left[\left\|g_{x}^{(k)}(X_{i}^{*},S_{i},\theta_{0})\right\|^{2}\right]$, $\mathbb{E}\left[\sup_{\theta\in\Theta}\left\|g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)\right\|\right]$, for $k\in\{0,\dots,K\}$, and $\mathbb{E}\left[{b_{j}(X_{i}^{*},S_{i},\theta_{0})}^{2}\right]$, $\mathbb{E}\left[\sup_{\theta\in\Theta}{b_{j}(X_{i}^{*},S_{i},\theta)}\right]$, for $j\in\{1,2\}$, are bounded;
  3. (iii)
    for some $\delta>0$, $\mathbb{E}\left[\sup_{\theta\in B_{\delta}(\theta_{0})}\left\|G_{x}^{(k)}(X_{i}^{*},S_{i},\theta)\right\|\right]$, for $k\in\{0,\dots,K\}$, and $\mathbb{E}\left[\sup_{\theta\in B_{\delta}(\theta_{0})}b_{Gj}(X_{i}^{*},S_{i},\theta)\right]$, for $j\in\{1,2\}$, are bounded;
  4. (iv)
[…]
  5. (v)
    Assumption 2.1 holds with $L\geqslant 2M$.
Assumptions A.2 and A.3 include basic regularity conditions, which help to ensure $\sqrt{n}$-consistency and asymptotic normality of the suggested estimator $\hat{\theta}$. Specifically, Assumption A.3 (i) is a counterpart of Assumption A.1 (i) applied to the Jacobian function. It ensures that the effect of the measurement error on the Jacobian is localized and allows us to establish $G\rightarrow G^{*}$, so $\Psi\rightarrow\Psi^{*}$. As a result, local identification of $\theta_{0}$ and the asymptotic properties of $\hat{\theta}$ are controlled by $G^{*}$ (and $\Psi^{*}$), the Jacobian associated with the correctly measured variables (see Assumption 2.3 (i)).
Appendix B Proof of Lemma 1
To stress that in our asymptotic approximation the variance and the higher moments of $\varepsilon_{i}$ depend on $n$, we will use $\sigma_{n}^{2}\equiv\mathbb{E}\left[\varepsilon_{i}^{2}\right]$, $\gamma_{0n}\equiv\gamma_{0}$.
Making use of Assumption A.1 (i) , we expand $g(X_{i},S_{i},\theta_{0})$ around $X_{i}^{*}$ as
$g(X_{i},S_{i},\theta_{0})=$ $g(X_{i}^{*},S_{i},\theta_{0})+g_{x}^{(1)}(X_{i}^{*},S_{i},\theta_{0})\varepsilon_{i}+\sum_{k=2}^{K}\frac{1}{k!}g_{x}^{(k)}(X_{i}^{*},S_{i},\theta_{0})\varepsilon_{i}^{k}$
$+\frac{1}{K!}\left(g_{x}^{(K)}(\tilde{X}_{i},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K},$ (B.1)
where $\tilde{X}_{i}$ lies between $X_{i}^{*}$ and $X_{i}$ (and hereafter $\tilde{X}_{i}$ is allowed to be component specific). Similarly, for $k^{\prime}\in\{2,\dots,K\}$, we have
$g_{x}^{(k)}(X_{i},S_{i},\theta_{0})=$ $g_{x}^{(k)}(X_{i}^{*},S_{i},\theta_{0})+\sum_{\ell=k+1}^{K}\frac{1}{(\ell-k)!}g_{x}^{(\ell)}(X_{i}^{*},S_{i},\theta_{0})\varepsilon_{i}^{\ell}$
$+\frac{1}{(K-k)!}\left(g_{x}^{(K)}(\tilde{X}_{ki},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K-k},$ (B.2)
where $\tilde{X}_{ki}$ lies between $X_{i}^{*}$ and $X_{i}$. Hence, combining these expressions and rearranging the terms, we obtain
$\psi(X_{i},S_{i},\theta_{0},\gamma)=$ $g(X_{i},S_{i},\theta_{0})-\sum_{k=2}^{K}\gamma_{k}g_{x}^{(k)}(X_{i},S_{i},\theta_{0})$
$=$ $g(X_{i}^{*},S_{i},\theta_{0})+g_{x}^{(1)}(X_{i}^{*},S_{i},\theta_{0})\varepsilon_{i}$
$+\sum_{k=2}^{K}g_{x}^{(k)}(X_{i}^{*},S_{i},\theta_{0})\left(\dfrac{1}{k!}\varepsilon_{i}^{k}-\sum_{\ell=2}^{k}\dfrac{1}{(k-\ell)!}\varepsilon_{i}^{k-\ell}\gamma_{\ell}\right)$
$+\dfrac{1}{K!}\left(g_{x}^{(K)}(\tilde{X}_{i},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K}$
$-\sum_{k=2}^{K}\dfrac{\gamma_{k}}{(K-k)!}\left(g_{x}^{(K)}(\tilde{X}_{ki},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K-k}.$ (B.3)
We want to show that for a properly chosen $\gamma=\gamma_{0n}$, $\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right]=o(n^{-1/2})$. Note that the first two terms in (B.3) are mean zero, i.e. we have
$\mathbb{E}\left[g(X_{i}^{*},S_{i},\theta_{0})\right]$ $=0,\quad\mathbb{E}\left[g_{x}^{(1)}(X_{i}^{*},S_{i},\theta_{0})\varepsilon_{i}\right]=0,$ (B.4)
where the latter is guaranteed by Assumptions 2.2.
Second, we argue that for a properly chosen $\gamma=\gamma_{0n}$, we have
$\mathbb{E}\left[\dfrac{1}{k!}\varepsilon_{i}^{k}-\sum_{\ell=2}^{k}\dfrac{1}{(k-\ell)!}\varepsilon_{i}^{k-\ell}\gamma_{0\ell n}\right]=0,$ (B.5)
for all $k\in\{2,\dots,K\}$. Let us reparameterize $\gamma_{0n}=(\gamma_{02n},\dots,\gamma_{0Kn})^{\prime}$ using $\gamma_{0kn}=\sigma_{n}^{k}a_{kn}$. Then, (B.5) can be rewritten as
$\mathbb{E}\left[\dfrac{1}{k!}(\varepsilon_{i}/\sigma_{n})^{k}-\sum_{\ell=2}^{k}\dfrac{1}{(k-\ell)!}(\varepsilon_{i}/\sigma_{n})^{k-\ell}a_{\ell n}\right]=0,$
which can also be represented as
$B_{n}a_{n}=c_{n}$ (B.6)
where $a_{n}=(a_{2n},\dots,a_{Kn})^{\prime}$, and
$B_{n}=\begin{bmatrix}1&0&\ldots&0&0\\ \mathbb{E}[\varepsilon_{i}/\sigma_{n}]&1&\ldots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \frac{\mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K-3}]}{(K-3)!}&\frac{\mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K-4}]}{(K-4)!}&\ldots&1&0\\ \frac{\mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K-2}]}{(K-2)!}&\frac{\mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K-3}]}{(K-3)!}&\ldots&\mathbb{E}[(\varepsilon_{i}/\sigma_{n})]&1\\ \end{bmatrix},\quad c_{n}=\begin{bmatrix}\mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{2}]/2!\\ \mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{3}]/3!\\ \vdots\\ \mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K-1}]/(K-1)!\\ \mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K}]/K!\end{bmatrix}.$
[…]
$\mathbb{E}\left[\sum_{k=2}^{K}g_{x}^{(k)}(X_{i}^{*},S_{i},\theta_{0})\left(\dfrac{1}{k!}\varepsilon_{i}^{k}-\sum_{\ell=2}^{k}\dfrac{1}{(k-\ell)!}\varepsilon_{i}^{k-\ell}\gamma_{0\ell n}\right)\right]=0.$ (B.7)
To complete the proof of $\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right]=o_{n}(n^{-1/2})$, it is sufficient to show that
$\mathbb{E}\left[\left(g_{x}^{(K)}(\tilde{X}_{i},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K}\right]$ $=o(n^{-1/2}),$ (B.8)
$\gamma_{0kn}\mathbb{E}\left[\left(g_{x}^{(K)}(\tilde{X}_{ki},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K-k}\right]$ $=o(n^{-1/2})$ (B.9)
for $k\in\{2,\dots,K\}$. We start with (B.8). Using Assumption A.1 (i), we obtain
$\left\|\left(g_{x}^{(K)}(\tilde{X}_{i},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K}\right\|\leqslant b_{1}(X_{i}^{*},S_{i},\theta_{0})\left|\varepsilon_{i}\right|^{K+1}+b_{2}(X_{i}^{*},S_{i},\theta_{0})\left|\varepsilon_{i}\right|^{M}.$ (B.10)
Hence, using Assumption 2.2, and the fact $\left|\tilde{X}_{i}-X_{i}^{*}\right|\leqslant\varepsilon_{i}$, we get
$\mathbb{E}\left[\left(g_{x}^{(K)}(\tilde{X}_{i},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K}\right]$
$\leqslant\sigma_{n}^{K+1}{\mathbb{E}\left[b_{1}(X_{i}^{*},S_{i},\theta_{0})\right]\mathbb{E}\left[\left|\varepsilon_{i}/\sigma_{n}\right|^{K+1}\right]}+\sigma_{n}^{M}{\mathbb{E}\left[b_{2}(X_{i}^{*},S_{i},\theta_{0})\right]\mathbb{E}\left[\left|\varepsilon_{i}/\sigma_{n}\right|^{M}\right]}.$
Since (i) the expectations above are bounded (Assumptions 2.1, A.1 (ii), and A.1 (iii)) and (ii) $\sigma_{n}^{K+1}=o(n^{-1/2})$ and $\sigma_{n}^{M}=o(n^{-1/2})$ (Assumption 2.1), this implies that (B.8) holds. To inspect (B.9), recall that $\gamma_{0kn}=\sigma_{n}^{k}a_{kn}$. As a result, using Assumptions A.1 (i) and 2.2, and $\left|\tilde{X}_{ki}-X_{i}^{*}\right|\leqslant\varepsilon_{i}$ again, we also have
$\gamma_{0kn}\mathbb{E}\left[\left(g_{x}^{(K)}(\tilde{X}_{ki},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K-k}\right]$
$\leqslant a_{kn}\left(\sigma_{n}^{K+1}{\mathbb{E}\left[b_{1}(X_{i}^{*},S_{i},\theta_{0})\right]\mathbb{E}\left[\left|\varepsilon_{i}/\sigma_{n}\right|^{K+1-k}\right]}+\sigma_{n}^{M}{\mathbb{E}\left[b_{2}(X_{i}^{*},S_{i},\theta_{0})\right]\mathbb{E}\left[\left|\varepsilon_{i}/\sigma_{n}\right|^{M-k}\right]}\right).$
[…]
Combining (B.3) with (B.4), and (B.7)-(B.9), we conclude that $\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right]=o(n^{-1/2})$.
Finally, we want to verify the recursive expressions for the components of $\gamma_{0n}$ using (B.5). First, $\gamma_{02n}=\mathbb{E}\left[\varepsilon_{i}^{2}\right]/2$ and $\gamma_{03n}=\mathbb{E}\left[\varepsilon_{i}^{3}\right]/6$ (since $\mathbb{E}\left[\varepsilon_{i}\right]=0$). For $k\geqslant 4$, suppose that $\gamma_{0\ell n}$ are known for $\ell\in\{2,\dots,k-1\}$. Then $\gamma_{0kn}$ can be directly computed from (B.5):
$\sum_{\ell=2}^{k}\frac{\mathbb{E}\left[\varepsilon_{i}^{k-\ell}\right]}{(k-\ell)!}\gamma_{0\ell n}=\frac{\mathbb{E}\left[\varepsilon_{i}^{k}\right]}{k!}.$
Plugging $\mathbb{E}\left[\varepsilon_{i}\right]=0$ and rearranging the terms give the expession in (12). Q.E.D.
Appendix C Proof of Theorem 2
Notation. To stress that in our asymptotic approximation the variance and the higher moments of $\varepsilon_{i}$ depend on $n$, we will use $\sigma_{n}^{2}\equiv\mathbb{E}\left[\varepsilon_{i}^{2}\right]$, $\gamma_{0n}\equiv\gamma_{0}$, and $\beta_{0n}\equiv\beta_{0}\equiv(\theta_{0}^{\prime},\gamma_{0n}^{\prime})^{\prime}$.
All vectors are columns. For some generic parameter vector $\alpha$ and a vector (or matrix) valued function $a(x,s,\alpha)$ and , let $a_{i}(\beta)\equiv a(X_{i},S_{i},\alpha)$, $\overline{a}(\alpha)\equiv n^{-1}\sum_{i=1}^{n}a_{i}(\alpha)$, $a(\alpha)\equiv\mathbb{E}[a_{i}(\alpha)]$. Similarly, we let $a_{i}^{*}(\alpha)\equiv a(X_{i}^{*},S_{i},\alpha)$, $\overline{a}^{*}(\alpha)\equiv n^{-1}\sum_{i=1}^{n}a_{i}^{*}(\alpha)$, $a^{*}(\alpha)\equiv\mathbb{E}[a_{i}^{*}(\alpha)]$.
For the true value of the parameter $\alpha_{0}$, we often write $a_{i}\equiv a(\alpha_{0})$, $\overline{a}\equiv\overline{a}(\alpha_{0})$, $a\equiv a(\alpha_{0})$, $a_{i}^{*}\equiv a(\alpha_{0})$, $\overline{a}^{*}\equiv\overline{a}^{*}(\alpha_{0})$, $a^{*}\equiv a^{*}(\alpha_{0})$.
C.1 Auxiliary lemmas
Lemma C.1.
Suppose that $\{(X_{i}^{*},S_{i}^{\prime},\varepsilon_{i})\}_{i=1}^{n}$ are i.i.d.. Then, under Assumptions 2.1, 2.2, A.1, A.2 (i), and A.3 (i)-(iii), we have
  1. (i)
    $\sup_{\theta\in\Theta}\left\|\overline{g}_{x}^{(k)}(\theta)-g_{x}^{(k)*}(\theta)\right\|=o_{p}(1)$
    and $g_{x}^{(k)*}(\theta)$ is continuous on $\Theta$ for $k\in\{0,\dots,K\}$;
  2. (ii)
    for some $\delta>0$,
    $\sup_{\theta\in B_{\delta}(\theta_{0})}\left\|\overline{G}_{x}^{(k)}(\theta)-G_{x}^{(k)*}(\theta)\right\|=o_{p}(1),$
    and $G_{x}^{(k)*}(\theta)$ is continuous on $B_{\delta}(\theta_{0})$ for $k\in\{0,\dots,K\}$.
Proof of Lemma C.1.
First, we show
$\sup_{\theta\in\Theta}\left\|\overline{g}(\theta)-g^{*}(\theta)\right\|=o_{p}(1).$
By the triangle inequality,
$\sup_{\theta\in\Theta}\left\|\overline{g}(\theta)-g^{*}(\theta)\right\|\leqslant\sup_{\theta\in\Theta}\left\|\overline{g}(\theta)-\overline{g}^{*}(\theta)\right\|+\sup_{\theta\in\Theta}\left\|\overline{g}^{*}(\theta)-g^{*}(\theta)\right\|.$
Then, it is sufficient to show that both terms on the right hand side of the inequality above are $o_{p}(1)$. Expanding $g(X_{i},S_{i},\theta_{0})$ around $X_{i}^{*}$ as in (B.1) and invoking Assumption A.1 (i),
$\sup_{\theta\in\Theta}\left\|\overline{g}(\theta)-\overline{g}^{*}(\theta)\right\|=$ $\sup_{\theta\in\Theta}\left\|\sum_{k=1}^{K-1}\frac{1}{k!}\frac{1}{n}\sum_{i=1}^{n}g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)\varepsilon_{i}^{k}+\dfrac{1}{K!}\frac{1}{n}\sum_{i=1}^{n}g_{x}^{(K)}(\tilde{X}_{i}^{*},S_{i},\theta)\varepsilon_{i}^{K}\right\|$
$\leqslant$ $\sum_{k=1}^{K}\frac{1}{k!}\underbrace{\frac{1}{n}\sum_{i=1}^{n}\sup_{\theta\in\Theta}\left\|g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)\right\||\varepsilon_{i}|^{k}}_{o_{p}(1)}$
$+\dfrac{1}{K!}\underbrace{\frac{1}{n}\sum_{i=1}^{n}\sup_{\theta\in\Theta}b_{1}(X_{i}^{*},S_{i},\theta)|\varepsilon_{i}|^{K+1}}_{o_{p}(1)}$
$+\dfrac{1}{K!}\underbrace{\frac{1}{n}\sum_{i=1}^{n}\sup_{\theta\in\Theta}b_{2}(X_{i}^{*},S_{i},\theta)|\varepsilon_{i}|^{M}}_{o_{p}(1)},$
where $\tilde{X}_{i}$ lies in between of $X_{i}^{*}$ and $X_{i}$. Now observe that all the terms following the inequality sign are $o_{p}(1)$. Indeed, this is guaranteed by Markov’s inequality paired with Assumptions 2.1, 2.2, and A.3 (ii). Hence, $\sup_{\theta\in\Theta}\left\|\overline{g}(\theta)-\overline{g}^{*}(\theta)\right\|=o_{p}(1)$, and we are left to show $\sup_{\theta\in\Theta}\left\|\overline{g}^{*}(\theta)-g^{*}(\theta)\right\|=o_{p}(1)$. This, in turn, follows from the standard ULLN (e.g., Lemma 2.4 in Newey and McFadden, 1994), which also ensures continuity of $g^{*}(\theta)$ on $\Theta$. Hence, we conclude that the assertion of the lemma holds for $g$.
Applying nearly identical arguments, one can also establish the desired results for $g_{x}^{(k)}$ for $k\in\{1,\dots,K\}$ and for $G_{x}^{(k)}$ for $k\in\{0,\dots,K\}$ (for the latter, Assumptions A.3 (i) and (iii) take the places of Assumptions A.1 (i) and A.3 (ii), respectively).
Lemma C.2.
[…]
Proof of Lemma C.2.
First, we prove the assertions of the lemma for $g_{x}^{(k)}$. Note that, by the standard expansion of $g_{x}^{(k)}(X_{i},S_{i},\theta_{0})$ around $X_{i}^{*}$ (see Eq. (B.2) above), we have
$\left\|g_{x}^{(k)}-g_{x}^{(k)^{*}}\right\|\leqslant$ $\mathbb{E}\left[\left\|g(X_{i},S_{i},\theta_{0})-g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)\right\|\right]$
$\leqslant$ $\sum_{\ell=k+1}^{K}\frac{1}{(\ell-k)!}\mathbb{E}\left[\left\|g_{x}^{(\ell)}(X_{i}^{*},S_{i},\theta_{0})\right\|\left|\varepsilon_{i}\right|^{\ell}\right]$
$+\frac{1}{(K-k)!}\mathbb{E}\left[\left\|\left(g_{x}^{(K)}(\tilde{X}_{i},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\right\|\left|\varepsilon_{i}\right|^{K-k}\right].$
By Assumptions 2.1, 2.2, and A.3 (ii), $\mathbb{E}\left[\left\|g_{x}^{(\ell)}(X_{i}^{*},S_{i},\theta_{0})\right\|\left|\varepsilon_{i}\right|^{\ell}\right]\rightarrow 0$ for all $\ell\in\{1,\dots,K\}$. Next, using Assumptions A.1 (i) and 2.2,
$\mathbb{E}\left[\left\|\left(g_{x}^{(K)}(\tilde{X}_{i},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\right\|\left|\varepsilon_{i}\right|^{K-k}\right]$
$\leqslant{\mathbb{E}\left[b_{1}(X_{i}^{*},S_{i},\theta_{0})\right]\mathbb{E}\left[\left|\varepsilon_{i}\right|^{K+1-k}\right]}+{\mathbb{E}\left[b_{2}(X_{i}^{*},S_{i},\theta_{0})\right]\mathbb{E}\left[\left|\varepsilon_{i}\right|^{M-k}\right]}\rightarrow 0,$
where the convergence follows from Assumptions 2.1, A.1 (ii) and A.3 (ii). Hence, we conclude $g_{x}^{(k)}\rightarrow g_{x}^{(k)*}$.
[…]
$\left\|\overline{g}_{x}^{(k)}(\hat{\theta})-g_{x}^{(k)*}\right\|\leqslant\sup_{\theta\in B_{\delta}(\theta_{0})}\left\|\overline{g}_{x}^{(k)}(\theta)-g_{x}^{(k)*}(\theta)\right\|+\left\|g_{x}^{(k)*}(\hat{\theta})-g_{x}^{(k)*}(\theta_{0})\right\|,$
[…]
A nearly identical argument, can be invoked to establish the same results for $G_{x}^{(k)}$ for $k\in\{0,\dots,K\}$, with Assumptions A.3 (i) and (iii) taking the places of Assumptions A.1 (i) and A.3 (ii), respectively.
Lemma C.3.
[…]
Proof of Lemma C.3.
First, we argue that $\sup_{\beta\in\mathcal{B}}\left\|\overline{\psi}(\beta)-\psi^{*}(\beta)\right\|=o_{p}(1)$. Notice that, by the triangle inequality,
$\sup_{\beta\in\mathcal{B}}\left\|\overline{\psi}(\beta)-\psi^{*}(\beta)\right\|\leqslant\sup_{\theta\in\Theta}\left\|\overline{g}(\theta)-g^{*}(\theta)\right\|+\sum_{k=2}^{K}\left|\gamma_{k}\right|\sup_{\theta\in\Theta}\left\|\overline{g}_{x}^{(k)}(\theta)-g_{x}^{(k)*}(\theta)\right\|=o_{p}(1),$ (C.1)
where the equality follows from Lemma C.1 (i) and boundedness of $\gamma$ (Assumption A.2 (i)). Moreover, Lemma C.1 (i) also ensures that $\psi^{*}(\beta)$ is continuous on compact $\mathcal{B}$ and, consequently, is bounded.
[…]
Lemma C.4.
Suppose that $\{(X_{i}^{*},S_{i}^{\prime},\varepsilon_{i})\}_{i=1}^{n}$ are i.i.d.. Then, under Assumptions 2.1, 2.2, A.1, and A.3 (ii) and (v), we have
[…]
where $\Omega_{gg}^{*}\equiv\mathbb{E}\left[g(X_{i},S_{i},\theta_{0})g(X_{i},S_{i},\theta_{0})^{\prime}\right]$.
Proof of Lemma C.4.
Using expansion (B.3), we obtain
$n^{1/2}\overline{\psi}(\beta_{0n})=$ $n^{-1/2}\sum_{i=1}^{n}g(X_{i}^{*},S_{i},\theta_{0})+n^{-1/2}\sum_{i=1}^{n}g_{x}^{(1)}(X_{i}^{*},S_{i},\theta_{0})\varepsilon_{i}$
$+\sum_{k=2}^{K}n^{-1/2}\sum_{i=1}^{n}g_{x}^{(k)}(X_{i}^{*},S_{i},\theta_{0})\left(\dfrac{1}{k!}\varepsilon_{i}^{k}-\sum_{\ell=2}^{k}\dfrac{1}{(k-\ell)!}\varepsilon_{i}^{k-\ell}\gamma_{0kn}\right)$
$+\dfrac{1}{K!}n^{-1/2}\sum_{i=1}^{n}\left(g_{x}^{(K)}(\tilde{X}_{i},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K}$
$-\sum_{k=2}^{K}\dfrac{\gamma_{0kn}}{(K-k)!}n^{-1/2}\sum_{i=1}^{n}\left(g_{x}^{(K)}(\tilde{X}_{ki},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K-k}.$ (C.2)
[…]
$n^{-1/2}\sum_{i=1}^{n}g_{x}^{(1)}(X_{i}^{*},S_{i},\theta_{0n})\varepsilon_{i}=o_{p}(1)$
Next, (B.7) ensures that we can similarly apply Chebyshev’s inequality (combined with Assumptions 2.1, 2.2, A.3 (ii) and (v)) to ensure that for $k\in\{2,\ldots,K\}$
$n^{-1/2}\sum_{i=1}^{n}g_{x}^{(k)}(X_{i}^{*},S_{i},\theta_{0})\left(\dfrac{1}{k!}\varepsilon_{i}^{k}-\sum_{\ell=2}^{k}\dfrac{1}{(k-\ell)!}\varepsilon_{i}^{k-\ell}\gamma_{0kn}\right)=o_{p}(1).$
Next, using (B.10),
$\left\|n^{-1/2}\sum_{i=1}^{n}\left(g_{x}^{(K)}(\tilde{X}_{i},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K}\right\|$
$\leqslant$ $n^{-1/2}\sum_{i=1}^{n}b_{1}(X_{i}^{*},S_{i},\theta_{0})\left|\varepsilon_{i}\right|^{K+1}+n^{-1/2}\sum_{i=1}^{n}b_{2}(X_{i}^{*},S_{i},\theta_{0})\left|\varepsilon_{i}\right|^{M}$
$\leqslant$ $\underbrace{n^{1/2}\sigma_{n}^{K+1}}_{\rightarrow 0}\underbrace{\left(n^{-1}\sum_{i=1}^{n}b_{1}(X_{i}^{*},S_{i},\theta_{0})\left|\varepsilon_{i}/\sigma_{n}\right|^{K+1}\right)}_{O_{p}(1)}$
$+\underbrace{n^{1/2}\sigma_{n}^{M}}_{\rightarrow 0}\underbrace{\left(n^{-1}\sum_{i=1}^{n}b_{2}(X_{i}^{*},S_{i},\theta_{0})\left|\varepsilon_{i}/\sigma_{n}\right|^{M}\right)}_{O_{p}(1)}=o_{p}(1),$
where both $n^{1/2}\sigma_{n}^{K+1}$ and $n^{1/2}\sigma_{n}^{M}$ converge to zero by Assumption 2.1, and the terms in the brackets are $O_{p}(1)$ by Markov’s inequality (ensured by Assumptions 2.1, 2.2, A.1 (ii) and A.3 (ii)). Recall that in the proof of Lemma 1, we have demonstrated that $\gamma_{0kn}=\sigma_{n}^{k}a_{kn}$, where $a_{kn}$ are bounded, for $k\in\{2,\dots,K\}$. Hence, similarly, we have
$\left\|\gamma_{0kn}n^{-1/2}\sum_{i=1}^{n}\left(g_{x}^{(K)}(\tilde{X}_{i},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K-k}\right\|$
$\leqslant$ $a_{kn}\sigma_{n}^{k}\left[n^{-1/2}\sum_{i=1}^{n}b_{1}(X_{i}^{*},S_{i},\theta_{0})\left|\varepsilon_{i}\right|^{K-k+1}+n^{-1/2}\sum_{i=1}^{n}b_{2}(X_{i}^{*},S_{i},\theta_{0})\left|\varepsilon_{i}\right|^{M-k}\right]$
$\leqslant$ $a_{kn}\underbrace{n^{1/2}\sigma_{n}^{K+1}}_{\rightarrow 0}\underbrace{\left(n^{-1}\sum_{i=1}^{n}b_{1}(X_{i}^{*},S_{i},\theta_{0})\left|\varepsilon_{i}/\sigma_{n}\right|^{K-k+1}\right)}_{O_{p}(1)}$
$+a_{kn}\underbrace{n^{1/2}\sigma_{n}^{M}}_{\rightarrow 0}\underbrace{\left(n^{-1}\sum_{i=1}^{n}b_{2}(X_{i}^{*},S_{i},\theta_{0})\left|\varepsilon_{i}/\sigma_{n}\right|^{M-k}\right)}_{O_{p}(1)}=o_{p}(1).$
Hence, we have demonstrated that all the remaining terms in (C.2) are $o_{p}(1)$, i.e. we have
$n^{1/2}\overline{\psi}(\beta_{0n})=$ $n^{-1/2}\sum_{i=1}^{n}g(X_{i}^{*},S_{i},\theta_{0})+o_{p}(1)$
[…]
which completes the proof.
C.2 Proof of Theorem 2
Equipped with Lemmas C.1-C.4, we are ready to prove Theorem 2.
Proof of Theorem 2.
Since (i) $\hat{\theta}$ and $\hat{\gamma}$ are consistent for $\theta_{0}$ and $\gamma_{0n}$, respectively (Lemma C.3) and (ii) both $\theta_{0}$ and $\gamma_{0n}\rightarrow 0$ (Assumption 2.1) are bounded away from the boundaries of $\Theta$ and $\Gamma$ respectively (Assumption A.2), the standard GMM FOC is satisfied with probability approaching one, i.e., we have (with probability approaching one)
$\overline{\Psi}(\hat{\beta})^{\prime}\hat{\Xi}\overline{\psi}(\hat{\beta})=0.$
Expanding $\overline{\psi}(\hat{\beta})$ around $\overline{\psi}(\beta_{0n})$ gives
$\overline{\Psi}(\hat{\beta})^{\prime}\hat{\Xi}\left(\overline{\psi}(\beta_{0n})+\overline{\Psi}(\tilde{\beta})(\hat{\beta}-\beta_{0})\right)=0,$ (C.3)
[…]
$\overline{\Psi}(\hat{\beta})=\left[\overline{G}(\hat{\theta})-\sum_{k=2}^{K}\hat{\gamma}_{k}\overline{G}_{x}^{(k)}(\hat{\theta}),-\overline{g}_{x}^{(2)}(\hat{\theta}),\dots,-\overline{g}_{x}^{(K)}(\hat{\theta})\right].$
[…]
$n^{1/2}(\hat{\beta}-\beta_{0n})$ $=-\left(\overline{\Psi}(\hat{\beta})^{\prime}\hat{\Xi}\overline{\Psi}(\tilde{\beta})\right)^{-1}\overline{\Psi}(\hat{\beta})^{\prime}\hat{\Xi}n^{1/2}\overline{\psi}(\beta_{0n})$
$=-\left(\Psi^{*\prime}\Xi\Psi^{*}\right)^{-1}\Psi^{*\prime}\Xi n^{1/2}\overline{\psi}(\beta_{0n})+o_{p}(1),$
[…]
where
$\Sigma^{*}=\left(\Psi^{*\prime}\Xi\Psi^{*}\right)^{-1}\Psi^{*\prime}\Xi\Omega_{gg}^{*}\Psi^{*}\Xi\left(\Psi^{*\prime}\Xi\Psi^{*}\right)^{-1}.$
To complete the proof, we need to show that $\Sigma\rightarrow\Sigma^{*}$. First, note that, by Lemma C.2 and $\gamma_{0n}\rightarrow 0$ (Assumption 2.1)
$\Psi=\left[G-\sum_{k=2}^{K}\gamma_{0kn}G_{x}^{(k)},-g_{x}^{(2)},\dots,-g_{x}^{(K)}\right]\rightarrow\left[G^{*},-g_{x}^{(2)*},\dots,-g_{x}^{(K)*}\right]=\Psi^{*}.$
Next, we want to argue that $\Omega_{\psi\psi}\rightarrow\Omega_{gg}^{*}$. Observe that
$\Omega_{\psi\psi}=\mathbb{E}\left[\left(g_{i}-\sum_{k=2}^{K}\gamma_{0kn}g_{xi}^{(k)}\right)\left(g_{i}-\sum_{k=2}^{K}\gamma_{0kn}g_{xi}^{(k)}\right)^{\prime}\right]=\mathbb{E}\left[g_{i}g_{i}^{\prime}\right]+o(1),$
where the equality follows since (i) $\gamma_{0kn}\rightarrow 0$ for all $k\in\{2,\dots,K\}$ (Assumption 2.1) and (ii) $\mathbb{E}\left[g_{xi}^{(k)}\left(g_{xi}^{(k^{\prime})}\right)^{\prime}\right]$ is bounded for all $k,k^{\prime}\in\{0,\dots,K\}$. In particular, (ii) can be inspected by expanding $g_{x}^{{(k)}}(X_{i},S_{i},\theta_{0})$ and $g_{x}^{{(k^{\prime})}}(X_{i},S_{i},\theta_{0})$ around $X_{i}^{*}$ as in (B.2) and bounding the expectations as in the proof of Lemma C.2 (using Assumptions 2.1, 2.2, A.1 (i), A.3 (ii), and A.3 (v)). Similarly, by expanding $g(X_{i},S_{i},\theta_{0})$ around $X_{i}^{*}$ and bounding the residual terms as in the proof of Lemma C.2 (again, using Assumptions 2.1, 2.2, A.1 (i), A.3 (ii), and A.3 (v)), we verify that $\mathbb{E}\left[g_{i}g_{i}^{\prime}\right]\rightarrow\mathbb{E}\left[g_{i}^{*}g_{i}^{*\prime}\right]=\Omega_{gg}^{*}$. Hence, $\Omega_{\psi\psi}\rightarrow\Omega_{gg}^{*}$ and, consequently, we verified that $\Sigma\rightarrow\Sigma^{*}$. Finally, we conclude
$n^{1/2}\Sigma^{-1/2}(\hat{\beta}-\beta_{0n})\rightarrow N(0,I_{\dim(\theta)+K-1}),$
which completes the proof.
Appendix D Details on the sufficient condition (17)
In this section, we demonstrate that the Jacobian $\Psi^{\ast}$ associated with the moment conditions (16) in Section 2.4 is guaranteed to have a full rank (for $K=2$) when the sufficient condition (17) holds.
It is convenient to define $\varrho_{i}\left(x,\theta\right)\equiv u\left(x,S_{i},\theta\right)\times\left(1,x,\ldots,x^{J-1}\right)^{\prime}$ and $u_{i}\left(x,\theta\right)\equiv\left(x,S_{i},\theta\right)$. Then
$g\left(x,S_{i},q,\theta\right)$$\equiv$ $\left(\begin{array}[]{c}u_{i}\left(x,\theta\right)\\ \varrho_{i}\left(x,\theta\right)x\\ \varrho_{i}\left(x,\theta\right)q\end{array}\right),\ \text{so}$
$g_{x}^{(2)}\left(x,S_{i},q,\theta\right)$ $=$ $\left(\begin{array}[]{l}u_{x,i}^{(2)}\left(x,\theta\right)\\ \varrho_{x,i}^{(2)}\left(x,\theta\right)x+2\varrho_{x,i}^{(1)}\left(x,\theta\right)\\ \varrho_{x,i}^{(2)}\left(x,\theta\right)q\end{array}\right).$
First, notice that
$\Psi^{\ast}=\mathbb{E}\begin{pmatrix}u_{\theta,i}\left(X_{i}^{\ast},\theta\right)&-u_{x,i}^{(2)}\left(X_{i}^{\ast},\theta\right)\\ \varrho_{\theta,i}\left(X_{i}^{\ast},\theta_{0}\right)X_{i}^{\ast}&-\varrho_{x,i}^{(2)}\left(X_{i}^{\ast},\theta_{0}\right)X_{i}^{\ast}-2\varrho_{x,i}^{(1)}\left(X_{i}^{\ast},\theta_{0}\right)\\ \varrho_{\theta,i}\left(X_{i}^{\ast},\theta_{0}\right)\left(\alpha_{1}X_{i}^{\ast}\right)&-\varrho_{x,i}^{(2)}\left(X_{i}^{\ast},\theta_{0}\right)\left(\alpha_{1}X_{i}^{\ast}\right)\end{pmatrix},$
where we used $\mathbb{E}\left[\varepsilon_{Q,i}|X_{i}^{\ast},S_{i},\varepsilon_{i}\right]=0$. Dividing moments $J+2,\ldots,2J+1$ by $\alpha_{1}$ and subtracting them from the moments $12,\ldots,J+$ 1, we obtain
$\mathop{\mathrm{R}k}\left(\Psi^{\ast}\right)$ $=$ $\mathop{\mathrm{R}k}\mathbb{E}\begin{pmatrix}u_{\theta,i}\left(X_{i}^{\ast},\theta\right)&-u_{x,i}^{(2)}\left(X_{i}^{\ast},\theta\right)\\ 0&-2\varrho_{x,i}^{(1)}\left(X_{i}^{\ast},\theta_{0}\right)\\ \varrho_{\theta,i}\left(X_{i}^{\ast},\theta_{0}\right)X_{i}^{\ast}&-\varrho_{x,i}^{(2)}\left(X_{i}^{\ast},\theta_{0}\right)X_{i}^{\ast}\end{pmatrix}$
$=$ $\mathop{\mathrm{R}k}\mathbb{E}\begin{pmatrix}u_{\theta,i}\left(X_{i}^{\ast},\theta\right)&-u_{x,i}^{(2)}\left(X_{i}^{\ast},\theta\right)\\ \varrho_{\theta,i}\left(X_{i}^{\ast},\theta_{0}\right)X_{i}^{\ast}&-\varrho_{x,i}^{(2)}\left(X_{i}^{\ast},\theta_{0}\right)X_{i}^{\ast}\\ 0&-2\varrho_{x,i}^{(1)}\left(X_{i}^{\ast},\theta_{0}\right)\end{pmatrix}$
$=$ $\mathop{\mathrm{R}k}\begin{pmatrix}H^{\ast}&\mathbb{E}\begin{pmatrix}u_{x,i}^{(2)}\left(X_{i}^{\ast},\theta\right)\\ \varrho_{x,i}^{(2)}\left(X_{i}^{\ast},\theta_{0}\right)X_{i}^{\ast}\end{pmatrix}\\ 0&2\mathbb{E}\left[\varrho_{x,i}^{(1)}\left(X_{i}^{\ast},\theta_{0}\right)\right]\end{pmatrix}.$
Here $\mathop{\mathrm{R}k}\left(H^{\ast}\right)=\dim\left(\theta\right)$, because this is the rank identification condition for $\theta_{0}$ in the model without EIV. Thus, for $\Psi^{\ast}$ to have full rank $\dim\left(\theta\right)+1$, it is sufficient to have $\mathbb{E}\left[\varrho_{x,i}^{(1)}\left(X_{i}^{\ast},\theta_{0}\right)\right]\neq 0$. Note that since $\mathbb{E}\left[u\left(X_{i}^{\ast},S_{i},\theta_{0}\right)|X_{i}^{\ast}\right]=0$, we have
$\mathbb{E}\left[\varrho_{x,i}^{(1)}\left(X_{i}^{\ast},\theta_{0}\right)\right]=\mathbb{E}\left[u_{x}^{(1)}\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\left(1,X_{i}^{\ast},\ldots,\left(X_{i}^{\ast}\right)^{J-1}\right)^{\prime}\right]\neq 0,$
where the last equality follows from (17). $\blacksquare$
Appendix E Evaluating the remaining bias of the MERM estimator
In this section, we provide additional details on the derivation of the (higher-order) bias of the MERM estimator of $v^{\prime}\hat{\beta}$ and discuss potential implications for inference.
The derivation of the expression for $\text{Bias}(v^{\prime}\hat{\beta})$ provided in equation (18) is based on the standard expansion of the GMM first order conditions around $\beta_{0}$. Following the lines of the proof of Lemma 1, we can consider a higher order expansion of $\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})$ to the order $o(\tau^{\overline{K}})$, which implies
$\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right]\approx\gamma_{0\overline{K}}\mathbb{E}\left[g_{x}^{(\overline{K})}(X_{i},S_{i},\theta_{0})\right],$
where $\overline{K}=K+2$ if $K$ is even and $\varepsilon_{i}$ is symmetric, and $\overline{K}=K+1$ otherwise. Hence,
$\text{Bias}(v^{\prime}\hat{\beta})\approx-v^{\prime}(\Psi^{\prime}\Xi\Psi)^{-1}\Psi^{\prime}\Xi\;\mathbb{E}\left[g_{x}^{(\overline{K})}(X_{i},S_{i},\theta_{0})\right]\gamma_{0\overline{K}}.$ (E.1)
Notice that, except for $\gamma_{0\overline{K}}$, all the remaining elements on the right hand side of (E.1) are either known or can be consistently estimated. The unknown $\gamma_{0\overline{K}}$ is given by the recursive relationship (12), and the only unknown determinant of $\gamma_{0\overline{K}}$ is $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ because the lower order moments of $\varepsilon_{i}$ can be recovered from $\hat{\gamma}$ with a required precision. Hence, the ranges of the possible values of $\gamma_{0\overline{K}}$ and, subsequently, of $\text{Bias}(v^{\prime}\hat{\beta})$ can be determined based on an a priori range of possible values of $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ (or, equivalently, of $\mathbb{E}[(\varepsilon_{i}/\sigma_{\varepsilon})^{\overline{K}}]$ combined with the estimated $\sigma_{\varepsilon}^{2}$) specified by the researcher.
Example (Evaluating the bias with $K=2$).
Suppose the researcher uses the MERM estimator with $K=2$, and believes that $\mathbb{E}\left[\varepsilon_{i}^{3}\right]=0$. Then, the bound on the bias depends on the $\mathbb{E}\left[\varepsilon_{i}^{4}\right]$. The variance $\sigma_{\varepsilon}^{2}$ is estimated by the MERM estimator, hence it is convenient to bound the kurtosis of $\varepsilon_{i}$ as $\mathbb{E}\left[\varepsilon_{i}^{4}\right]/\sigma_{\varepsilon}^{4}\in[1,\overline{\kappa}]$ for some $\overline{\kappa}$. For example, one could take $\overline{\kappa}=10$ as a conservative bound for the kurtosis. [^13] For Gaussian $\varepsilon_{i}$ the kurtosis $\kappa$ is $3$. Student’s $t(\nu)$ distribution has $\kappa<10$ for all $\nu\geq 5$ (the kurtosis and the $4$th moment do not exist for $\nu\leq 4$). Since $\hat{\gamma}$ estimates $\gamma_{0}=\gamma_{02}=\sigma_{\varepsilon}^{2}/2$, the range of possible values of $\gamma_{04}$ can be approximately bounded using
$\gamma_{04}=\frac{\mathbb{E}\left[\varepsilon_{i}^{4}\right]-6\sigma_{\varepsilon}^{4}}{24}\in\left[-\frac{5}{24}\sigma_{\varepsilon}^{4},\;\frac{\overline{\kappa}-6}{24}\sigma_{\varepsilon}^{4}\right]=\left[-\frac{5}{6}\gamma_{02}^{2},\;\frac{\overline{\kappa}-6}{6}\gamma_{02}^{2}\right].{}$
Hence, the range of possible values of $\text{Bias}(v^{\prime}\hat{\beta})$ can be approximated as
$\text{Bias}(v^{\prime}\hat{\beta})\in\begin{cases}\left[-5\overline{b}\hat{\gamma}^{2}/6,\;\overline{b}\hat{\gamma}^{2}(\overline{\kappa}-6)/6\right],&\text{if $\overline{b}\geq 0$},\\ \left[\overline{b}\hat{\gamma}^{2}(\overline{\kappa}-6)/6,\;-5\overline{b}\hat{\gamma}^{2}/6\right],&\text{otherwise},\end{cases}$
where
$\overline{b}\equiv-v^{\prime}(\hat{\Psi}^{\prime}\hat{\Xi}\hat{\Psi})^{-1}\hat{\Psi}^{\prime}\hat{\Xi}\;\left(\frac{1}{n}\sum_{i=1}^{n}g_{x}^{(4)}(X_{i},S_{i},\hat{\theta})\right).$
Notice that if $\overline{\kappa}\leq 6$, we have $\gamma_{04}\leq 0$, and the sign of $\text{Bias}(v^{\prime}\hat{\beta})$ becomes known.
If (the absolute value of the worst case) $\text{Bias}(v^{\prime}\hat{\beta})$ is sufficiently small relative to its standard error $\sqrt{v^{\prime}\hat{\Sigma}v/n}$, the estimator $v^{\prime}\hat{\beta}$ is approximately unbiased and the inference on $v^{\prime}\beta_{0}$ can be based on equation (14). Otherwise, the researcher could consider using a higher $K$ and/or possibly employing inference tools that take into account the remaining bias (e.g., ArmstrongKolesar2021QE).
Appendix F MERM derivation when $\sigma_{\varepsilon}$ is not small
Note that $\tau$ can be small without $\sigma_{\varepsilon}$ being small in absolute magnitude. For example, suppose $\sigma_{\varepsilon}=10$ and $\sigma_{X^{\ast}}=100$. Then $\tau=0.1$, so the measurement error is quite small relative to $\sigma_{X^{\ast}}$, and relying on the approximation $\tau\rightarrow 0$ is reasonable. At the same time, approximation $\sigma_{\varepsilon}\rightarrow 0$ may not be suitable for this example.
In this Appendix we show that the corrected moment conditions and the MERM estimator are valid without assuming that $\sigma_{\varepsilon}$ is small in absolute magnitude. In Section 2 we used Taylor expansions in $\varepsilon_{i}$ around $\varepsilon_{i}=0$, with the remainder of order $\mathbb{E}\left[\left|\varepsilon_{i}\right|^{K+1}\right]$. When $\sigma_{\varepsilon}>1$, term $O\left(\mathbb{E}\left[\left|\varepsilon_{i}\right|^{K+1}\right]\right)$ in equation (9) cannot be viewed as a negligible remainder, because $\mathbb{E}\left[\left|\varepsilon_{i}\right|^{K+1}\right]>1$ and, moreover, terms $\mathbb{E}\left[\left|\varepsilon_{i}\right|^{k}\right]$ increase rather than decrease with $k$.
In Section 2, to simplify the exposition, we have assumed that $X^{\ast}$ is scaled so that $\sigma_{X^{\ast}}$ is of order one. This in particular ensures that $\mathbb{E}\left[\left|\varepsilon_{i}\right|^{k}\right]$ decrease with $k$. We will now show that this assumption about the scale of $X^{\ast}$ is not necessary, and that the procedure remains valid without any such scaling.
We will show that by rescaling the Taylor expansions in Section 2 can be written in terms of powers of $\tau^{k}$, which necessarily decrease with $k$ when $\tau<1$.
Remember the model of Section 2:
$\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta_{0})]=0,\quad X_{i}=X_{i}^{\ast}+\varepsilon_{i},\quad\mathbb{E}[\varepsilon_{i}]=0.$ (F.1)
Let $\xi_{i}$ denote a random variable with $\mathbb{E}\left[\xi_{i}\right]=0$ and $\mathbb{E}\left[\xi_{i}^{2}\right]=1$, $\mathbb{E}\left[\left|\xi_{i}\right|^{L+1}\right]$ is bounded, and $\varepsilon_{i}\equiv\sigma_{\varepsilon}\xi_{i}$. Also, let us denote
$\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}},\quad\widetilde{X}_{i}\equiv X_{i}/\sigma_{X^{\ast}},\quad\widetilde{X}_{i}^{\ast}\equiv X_{i}^{\ast}/\sigma_{X^{\ast}},\quad\widetilde{g}(\widetilde{x},s,\theta)\equiv g(\sigma_{X^{\ast}}\widetilde{x},s,\theta).$
Then, we can rewrite equation (F.1) as
$\mathbb{E}[\widetilde{g}(\widetilde{X}_{i}^{\ast},S_{i},\theta_{0})]=0,\quad\widetilde{X}_{i}=\widetilde{X}_{i}^{\ast}+\tau\xi_{i},\quad\mathbb{E}[\xi_{i}]=0.$
Expand $\widetilde{g}(\widetilde{X}_{i},S_{i},\theta)=\widetilde{g}(\widetilde{X}_{i}^{\ast}+\tau\xi_{i},S_{i},\theta)$ around $\tau=0$ to obtain
$\mathbb{E}[\widetilde{g}(\widetilde{X}_{i},S_{i},\theta)]=\mathbb{E}[\widetilde{g}(\widetilde{X}_{i}^{\ast},S_{i},\theta)]+\sum_{k=2}^{K}\frac{\tau^{k}\mathbb{E}\left[\xi_{i}^{k}\right]}{k!}\mathbb{E}\left[\widetilde{g}_{x}^{(k)}(\widetilde{X}_{i}^{\ast},S_{i},\theta)\right]+O(\tau^{K+1}),$
which is similar to equation (9), except $\mathbb{E}\left[\varepsilon_{i}^{k}\right]$ is replaced by $\tau^{k}\mathbb{E}\left[\xi_{i}^{k}\right]$, and $\widetilde{X}_{i}$, $\widetilde{X}_{i}^{\ast}$, $\widetilde{g}$ are replaced by $X_{i}$, $X_{i}^{\ast}$, $g$. Then, the corrected moment condition has the form
$\widetilde{\psi}(\widetilde{X}_{i},S_{i},\theta,\widetilde{\gamma})=\widetilde{g}(\widetilde{X}_{i},S_{i},\theta)-\sum_{k=2}^{K}\widetilde{\gamma}_{k}\widetilde{g}_{x}^{(k)}(\widetilde{X}_{i},S_{i},\theta),$ (F.2)
where true parameter values $\widetilde{\gamma}_{0}$ are $\widetilde{\gamma}_{02}=\tau^{2}\mathbb{E}\left[\xi_{i}^{2}\right]/2=\tau^{2}/2$, $\widetilde{\gamma}_{03}=\tau^{3}\mathbb{E}\left[\xi_{i}^{3}\right]/6$, and $\widetilde{\gamma}_{0k}=\frac{\tau^{k}\mathbb{E}\left[\xi_{i}^{k}\right]}{k!}-\sum_{\ell=2}^{k-2}\frac{\tau^{k-\ell}\mathbb{E}\left[\xi_{i}^{k-\ell}\right]}{(k-\ell)!}\widetilde{\gamma}_{0\ell}$ for $k\geq 4$.
We will now show that
$\gamma_{0k}=\sigma_{X^{\ast}}^{k}\widetilde{\gamma}_{0k}\text{ for all }k\geq 2\text{.}$
First, $\gamma_{02}=\mathbb{E}\left[\varepsilon_{i}^{2}\right]/2=\mathbb{E}\left[\left(\sigma_{\varepsilon}\xi_{i}\right)^{2}\right]/2=\sigma_{X^{\ast}}^{2}\widetilde{\gamma}_{02}$, $\gamma_{03}=\mathbb{E}\left[\varepsilon_{i}^{3}\right]/6=\sigma_{X^{\ast}}^{3}\widetilde{\gamma}_{03}$ by definition. Then, for $k\geq 4$, by induction we have
$\gamma_{0k}$ $=$ $\frac{\mathbb{E}\left[\varepsilon_{i}^{k}\right]}{k!}-\sum_{\ell=2}^{k-2}\frac{\mathbb{E}\left[\varepsilon_{i}^{k-\ell}\right]}{(k-\ell)!}\gamma_{0\ell}$
$=$ $\sigma_{X^{\ast}}^{k}\left(\frac{\left(\sigma_{\varepsilon}/\sigma_{X^{\ast}}\right)^{k}\mathbb{E}\left[\xi_{i}^{k}\right]}{k!}-\sum_{\ell=2}^{k-2}\frac{\left(\sigma_{\varepsilon}/\sigma_{X^{\ast}}\right)^{k-\ell}\mathbb{E}\left[\xi_{i}^{k-\ell}\right]}{(k-\ell)!}\frac{\gamma_{0\ell}}{\sigma_{X^{\ast}}^{\ell}}\right)$
$=$ $\sigma_{X^{\ast}}^{k}\left(\frac{\tau^{k}\mathbb{E}\left[\xi_{i}^{k}\right]}{k!}-\sum_{\ell=2}^{k-2}\frac{\tau^{k-\ell}\mathbb{E}\left[\xi_{i}^{k-\ell}\right]}{(k-\ell)!}\widetilde{\gamma}_{0\ell}\right)=\sigma_{X^{\ast}}^{k}\widetilde{\gamma}_{0k}\text{.}$
Finally, let us now show that moment condition $\widetilde{\psi}$ in equation (F.2) is numerically identical to $\psi$ in equation (11) with $\gamma_{k}=\sigma_{X^{\ast}}^{k}\widetilde{\gamma}_{k}$. Note that for $\widetilde{x}=x/\sigma_{X^{\ast}}$ we have $\widetilde{g}_{\widetilde{x}}^{(k)}(\widetilde{x},s,\theta)\equiv\nabla_{\widetilde{x}}^{k}g(\sigma_{X^{\ast}}\widetilde{x},s,\theta)=\sigma_{X^{\ast}}^{k}g_{a}^{(k)}(a,s,\theta)|_{a=\sigma_{X^{\ast}}\widetilde{x}}=\sigma_{X^{\ast}}^{k}g_{x}^{(k)}(x,s,\theta)$, and hence
$\widetilde{\psi}(\widetilde{X}_{i},S_{i},\theta,\widetilde{\gamma})$ $=$ $g(\sigma_{X^{\ast}}\widetilde{X}_{i},S_{i},\theta)-\sum_{k=2}^{K}\left(\widetilde{\gamma}_{k}\sigma_{X^{\ast}}^{k}\right)g_{x}^{(k)}(\sigma_{X^{\ast}}\widetilde{X}_{i},S_{i},\theta)$
$=$ $g(X_{i},S_{i},\theta)-\sum_{k=2}^{K}\left(\widetilde{\gamma}_{k}\sigma_{X^{\ast}}^{k}\right)g_{x}^{(k)}(X_{i},S_{i},\theta)$
$=$ $\psi(X_{i},S_{i},\theta,\gamma)\text{.}$
Appendix G Some Implementation Details
Numerical Optimization
Since $\overline{\psi}\left(\theta,\gamma\right)$ is a linear function of $\gamma$ it can be profiled out of the quadratic form $\hat{Q}(\theta,\gamma)$. Thus, the criterion function only needs to be minimized numerically over $\theta$.
Choice of the weighting matrix $\hat{\Xi}$
As for the standard GMM estimator, the optimal weighting matrix can be estimated by
$\hat{\Xi}_{\text{eff}}\equiv\hat{\Omega}_{\psi\psi}^{-1}(\tilde{\theta},\tilde{\gamma}),$
where $\tilde{\theta}$ and $\tilde{\gamma}$ are some preliminary estimators of $\theta_{0}$ and $\gamma_{0}$, and $\hat{\Omega}_{\psi\psi}(\theta,\gamma)\equiv n^{-1}\sum_{i=1}^{n}\psi_{i}(\theta,\gamma)\psi_{i}(\theta,\gamma)^{\prime}.$ One example of such a preliminary estimator would be the 1-step (GMM-)MERM estimator using $\hat{\Xi}_{\text{GMM}1}\equiv\hat{\Omega}_{\psi\psi}^{-1}(\hat{\theta}_{\text{Naive}},0)$ as the first-step GMM weighting matrix, where $\hat{\theta}_{\text{Naive}}$ is a naive estimator of $\theta_{0}$ that ignores EIV. Note that $\hat{\Omega}_{\psi\psi}(\hat{\theta}_{\text{Naive}},0)=$ $\hat{\Omega}_{gg}\left(\hat{\theta}_{\text{Naive}}\right)$, where $\hat{\Omega}_{gg}(\theta)\equiv n^{-1}\sum_{i=1}^{n}g_{i}(\theta)g_{i}(\theta)^{\prime}$.
One may also consider the regularized version of the efficient weighting matrix estimator $\hat{\Xi}_{\text{eff,R}}\equiv\hat{\Omega}_{\psi\psi}^{-1}(\tilde{\theta},0)$. Since $\gamma_{0}\rightarrow 0$, using the regularized version $\hat{\Xi}_{\text{eff,R}}$ does not lead to a loss of efficiency. Moreover, our simulation studies suggest that using the regularized weighting matrix $\hat{\Xi}_{\text{eff,R}}$ results in better finite sample performance of the MERM estimator and, hence, is recommended in practice.
Although not indicated by the notation in equation (13), the weighting matrix $\hat{\Xi}\equiv\hat{\Xi}(\theta,\gamma)$ is allowed to be a function of $\theta$ and $\gamma$. For example, Continuously Updating GMM Estimator (CUE) corresponds to taking $\hat{\Xi}_{\text{CUE}}\left(\theta,\gamma\right)\equiv\hat{\Omega}_{\psi\psi}^{-1}(\theta,\gamma)$. Similarly to $\hat{\Xi}_{\text{eff,R}}$, one may also consider $\hat{\Xi}_{\text{CUE,R}}\left(\theta,\gamma\right)\equiv\hat{\Omega}_{\psi\psi}^{-1}(\theta,0)$ without introducing any loss of efficiency. In contrast to the criterion function of the CUE estimator, criterion function of $\hat{Q}_{\text{CUE,R}}(\theta,\gamma)$ is quadratic in $\gamma$. This implies that $\gamma$ can be profiled out analytically. This simplifies the numerical optimization problem reducing it to minimizing $\hat{Q}_{\text{CUE,R}}\left(\theta,\hat{\gamma}\left(\theta\right)\right)$ over $\theta\in\Theta$. Then, the dimension of the optimization parameter $\theta$ for the corrected moment condition problem remains the same as for the original (naive) estimation problem without the EIV correction.
Estimation of the asymptotic variance $\Sigma$
Theorem 2 shows that the MERM estimator $\hat{\beta}=(\hat{\theta}^{\prime},\hat{\gamma}^{\prime})^{\prime}$ behaves like a standard GMM estimator based on the corrected moment function $\psi(\theta,\gamma)$. The researcher can rely on the standard GMM inference procedures. The asymptotic variance of $\hat{\beta}$ can be consistently estimated by
$\hat{\Sigma}\equiv(\hat{\Psi}^{\prime}\hat{\Xi}\hat{\Psi})^{-1}\hat{\Psi}^{\prime}\hat{\Xi}\hat{\Omega}_{\psi\psi}\hat{\Xi}\hat{\Psi}(\hat{\Psi}^{\prime}\hat{\Xi}\hat{\Psi})^{-1},$
where, $\hat{\Xi}$ is the chosen weighting matrix, and $\hat{\Psi}\equiv\overline{\Psi}(\hat{\theta},\hat{\gamma})=n^{-1}\sum_{i=1}^{n}\Psi_{i}(\hat{\theta},\hat{\gamma})$ and $\hat{\Omega}_{\psi\psi}=\hat{\Omega}_{\psi\psi}(\hat{\theta},\hat{\gamma})$ are estimators of $\Psi$ and $\Omega_{\psi\psi}$.
Appendix H Implementation Details of the Empirical Illustration
In this section, we provide additional details on the implementation of the numerical experiment in Section 3.3.
Data
The original dataset is the ModeCanada dataset supplied with the R package mlogit. This dataset has been extensively used in transportation research. For a detailed description of the dataset see, for example, Koppelman and Wen (2000), Wen and Koppelman (2001), and Hansen (2022). As in Koppelman and Wen (2000), we use only the subset of travelers who chose train, air, or car (and had all of those alternatives available for them), which leaves $n=2769$ observations.
Monte-Carlo design
We choose $\theta_{0}$ to be the MLE estimates using the considered dataset, which are reported in the table bellow.
$\theta_{1}$ $\theta_{2}$ $\theta_{3}$ $\theta_{4}$ $\theta_{5}$ $\theta_{6}$ $\theta_{7}$ $\theta_{8}$
Estimates 0.0355 0.2976 -2.0891 0.0079 -0.9900 1.8794 -0.0223 -0.0149
Std. Err. 0.0036 0.0844 0.4674 0.0036 0.0876 0.2037 0.0038 0.0008
To generate the simulated samples, we randomly draw the covariates (with replacement) from their joint empirical distribution. To ensure identification of the model, we also generate an instrumental variable $Z_{i}$ as
$Z_{i}=\kappa\thinspace Income_{i}^{*}/\sigma_{Income^{*}}+\sqrt{1-\kappa^{2}}\zeta_{i},$
where $\sigma_{Income^{*}}\approx 17.5$ is the standard deviation of $Income^{*}$, $\kappa=0.5$, and $\zeta_{i}$ are i.i.d. draws from $N(0,1)$ (which are also independent from all the over variables). Note that the instrument $Z_{i}$ is “caused by $X_{i}^{*}$”. For example, $Z_{i}$ can be some (noisy) measure of individual consumption.
Moments
To simplify the notation, let $X_{i}^{*}\equiv Income_{i}^{*}$, $X_{i}\equiv Income_{i}$, $R_{i}\equiv Urban_{i}$, $R_{ij}\equiv(Price_{ij},InTime_{ij})^{\prime}$ for $j\in\{0,1,2\}$, and $W_{i}\equiv(R_{i},R_{i1}^{\prime},R_{i2}^{\prime},R_{i0}^{\prime})^{\prime}$. Also let $Y_{ij}\equiv\mathbbm{1}\{j=\operatorname*{\mathrm{arg}\!\max\limits}_{j^{\prime}\in\{0,1,2\}}U_{ij^{\prime}}\}$ for $j\in\{0,1,2\}$, $Y_{i}\equiv(Y_{i1},Y_{i2},Y_{i0})^{\prime}$, and $p_{j}(x,w,\theta)\equiv\mathbb{P}(Y_{ij}=1|X_{i}^{*}=x,W_{i}=w;\theta)$ with $w\equiv(r,r_{1}^{\prime},r_{2}^{\prime},r_{0}^{\prime})$, so
$p_{1}(x,w,\theta)$ $=\frac{e^{\theta_{1}x+\theta_{2}r+\theta_{3}+(\theta_{7},\theta_{8})r_{1}}}{e^{\theta_{1}x+\theta_{2}r+\theta_{3}+(\theta_{7},\theta_{8})r_{1}}+e^{\theta_{4}x+\theta_{5}r+\theta_{6}+(\theta_{7},\theta_{8})r_{2}}+e^{(\theta_{7},\theta_{8})r_{0}}},$
$p_{2}(x,w,\theta)$ $=\frac{e^{\theta_{4}x+\theta_{5}r+\theta_{6}+(\theta_{7},\theta_{8})r_{2}}}{e^{\theta_{1}x+\theta_{2}r+\theta_{3}+(\theta_{7},\theta_{8})r_{1}}+e^{\theta_{4}x+\theta_{5}r+\theta_{6}+(\theta_{7},\theta_{8})r_{2}}+e^{(\theta_{7},\theta_{8})r_{0}}},$
and $p_{0}(x,w,\theta)=1-p_{1}(x,w,\theta)-p_{2}(x,w,\theta)$. Then, the original moment function takes the form of
$g(x,w,y,z,\theta)=\left(\left(y_{1}-p_{1}(x,w,\theta)\right)\varphi_{1}(x,z,w)^{\prime},\left(y_{2}-p_{2}(x,w,\theta)\right)\varphi_{2}(x,z,w)^{\prime}\right)^{\prime}.$
and $\varphi_{j}(x,z,w)=\left(1,x,z,x^{2},z^{2},x^{3},z^{3},r,(r_{j}-r_{0})^{\prime}\right)^{\prime}$ for $K=2$ and $\varphi_{j}(x,z,w)=\left(1,x,z,x^{2},xz,z^{2},x^{3},x^{2}z,xz^{2},z^{3},r,(r_{j}-r_{0})^{\prime}\right)^{\prime}$ for $K=4$.
Income Elasticities
In Section 3.3, we focus on estimation of and inference on the income elasticities
$\frac{\partial\ln p_{j}}{\partial\ln x}(x,w,\theta)=\frac{x}{p_{j}(x,w,\theta)}\frac{\partial p_{j}(x,w,\theta)}{\partial x}.$
We report the results are for the income elasticities evaluated at the sample mean of $X^{*}$ and $W$ in the original sample.
Estimation of and Inference on the $\theta_{0}$
In Table 5 below, we also report the estimation and inference results for the vector of parameters $\theta_{0}$ underlying the reported results about elasticities.
Table 5: Simulation results for the empirically calibrated conditional logit model
MLE $K=2$ $K=4$
bias std rmse size bias std rmse size bias std rmse size
$\tau=1/4$
$\theta_{1}$0.0021 0.0035 0.0041 8.70 0.0001 0.0042 0.0042 5.48 0.0005 0.0057 0.0058 7.38
$\theta_{2}$ 0.0047 0.0932 0.0933 5.10 0.0028 0.0957 0.0957 5.36 0.0022 0.0960 0.0960 5.40
$\theta_{3}$ 0.1152 0.4452 0.4599 6.00 -0.0048 0.4821 0.4821 5.94 -0.0251 0.5336 0.5342 6.68
$\theta_{4}$0.0004 0.0031 0.0031 4.52 -0.0001 0.0034 0.0034 5.32 -0.0001 0.0036 0.0036 6.86
$\theta_{5}$0.0023 0.0894 0.0895 5.18 -0.0088 0.0918 0.0922 5.54 -0.0113 0.0922 0.0929 5.96
$\theta_{6}$ 0.0232 0.1821 0.1836 4.64 0.0250 0.1982 0.1998 5.72 0.0329 0.2089 0.2115 6.74
$\theta_{7}$0.0001 0.0035 0.0035 5.58 -0.0002 0.0036 0.0036 6.24 -0.0003 0.0036 0.0037 6.04
$\theta_{8}$0.0001 0.0007 0.0007 4.82 -0.0001 0.0007 0.0007 5.48 -0.0002 0.0007 0.0007 5.58
$\tau=1/2$
$\theta_{1}$0.0073 0.0032 0.0080 60.08 -0.0016 0.0043 0.0046 6.86 0.0005 0.0061 0.0061 6.60
$\theta_{2}$ 0.0109 0.0930 0.0936 5.18 0.0050 0.0959 0.0960 5.54 0.0026 0.0964 0.0965 5.36
$\theta_{3}$ 0.4080 0.4452 0.6039 17.12 0.0936 0.4874 0.4963 6.58 -0.0263 0.5475 0.5481 6.38
$\theta_{4}$0.0012 0.0029 0.0031 6.46 -0.0003 0.0035 0.0035 5.22 -0.0002 0.0038 0.0038 6.52
$\theta_{5}$0.0006 0.0894 0.0894 5.16 -0.0083 0.0919 0.0923 5.52 -0.0110 0.0924 0.0930 5.92
$\theta_{6}$ 0.0655 0.1752 0.1870 6.22 0.0348 0.2035 0.2064 5.86 0.0326 0.2158 0.2183 6.42
$\theta_{7}$0.0003 0.0035 0.0036 5.64 -0.0003 0.0036 0.0037 6.34 -0.0003 0.0037 0.0037 6.06
$\theta_{8}$0.0001 0.0007 0.0007 5.06 -0.0001 0.0007 0.0007 5.54 -0.0002 0.0007 0.0007 5.48
$\tau=3/4$
$\theta_{1}$0.0132 0.0029 0.0135 99.34 -0.0056 0.0043 0.0071 25.12 0.0003 0.0065 0.0065 6.06
$\theta_{2}$ 0.0180 0.0923 0.0940 5.36 0.0102 0.0961 0.0966 5.76 0.0033 0.0973 0.0973 5.44
$\theta_{3}$ 0.7336 0.4496 0.8604 41.66 0.3203 0.4859 0.5820 12.00 -0.0130 0.5666 0.5667 6.20
$\theta_{4}$0.0024 0.0026 0.0035 14.00 -0.0009 0.0035 0.0036 5.94 -0.0002 0.0041 0.0041 5.76
$\theta_{5}$ 0.0021 0.0890 0.0891 5.08 -0.0071 0.0921 0.0924 5.68 -0.0109 0.0926 0.0932 5.82
$\theta_{6}$ 0.1204 0.1654 0.2046 9.76 0.0648 0.2048 0.2148 6.56 0.0334 0.2294 0.2318 5.98
$\theta_{7}$0.0004 0.0036 0.0036 6.00 -0.0004 0.0036 0.0037 6.34 -0.0003 0.0037 0.0037 6.06
$\theta_{8}$0.0001 0.0007 0.0007 5.54 -0.0002 0.0007 0.0008 6.00 -0.0002 0.0007 0.0008 5.42
  • This table reports the simulated finite sample bias, standard deviation, RMSE, and size of the MLE and the MERM estimators and the corresponding t-tests for the components of $\theta_{0}$. The true value of the parameters of interest are $\theta_{0}=(0.0355,0.2976,-2.0891,0.0079,-0.9900,1.8794,-0.0223,-0.0149)^{\prime}$. The results are based on 5,000 replications.

### v4

@@
 This is true even when the EIV are so large that naive estimators (that ignore the EIV problem) are heavily biased with their confidence intervals having 0% coverage.
-Our approach involves no nonparametric estimation, which is especially important for applications with many covariates, and settings with multivariate or non-classical EIV.
+Our approach involves no nonparametric estimation, which is especially important for applications with many covariates and settings with multivariate EIV.
 In particular, the approach makes it easy to use instrumental variables to address EIV in nonlinear models.
-Keywords: errors-in-variables, nonstandard asymptotic approximation, nonclassical measurement errors, nonparametric identification, instrumental variables
+Keywords: errors-in-variables, nonstandard asymptotic approximation, nonparametric identification, instrumental variables
 1 Introduction
@@
 Addressing the Errors-In-Variables (EIV) bias in nonlinear models requires elaborate strategies.
-[^1] See Hausman et al.
-(1991, 1995); Newey (2001); Schennach (2007); Li (2002); Schennach (2004); Chen et al.
-(2005); Hu and Schennach (2008); Schennach (2014); Wilhelm (2019), among others.
+[^1] See HINP1991JoE; HausmanNeweyPowell1995JoE; Newey2001REStat; Schennach2007Ecta; Li2002JoE; Schennach2004Ecta; ChenHongTamer2005ReStud; HuSchennach2008Ecta; Schennach2014Ecta-ELVIS; Wilhelm2019WP-TestingForME, among others.
 Despite the fundamental theoretical progress in identification and estimation of nonlinear models with EIV, the problem of EIV is still rarely addressed in empirical work outside of linear specifications.
@@
 $X_{i}=X_{i}^{\ast}+\varepsilon_{i}.$
-The measurement error can be classical or non-classical.
 We will refer to $g\left(\cdot\right)$ as the original moment function, since it would have been valid had the researcher observed $X_{i}^{\ast}$.
@@
 Even in this well-studied example of nonlinear regression, estimation in the presence of the EIV is a difficult problem.
-Importantly, nonlinear instrumental variable regression estimator cannot be used, since it is inconsistent in the presence of EIV (Amemiya, 1985).
+Importantly, nonlinear instrumental variable regression estimator cannot be used, since it is inconsistent in the presence of EIV (Amemiya1985JoE).
 The existing approaches typically require nonparametric estimation that can be impractical in many empirical applications.
@@
 In particular, our approach makes it easy to use instrumental variables to address EIV in nonlinear models.
-To provide a practical estimation approach for the general class of models (1), we focus on the empirical settings in which the researcher believes the variability of the measurement error to be at most a fraction of the variability of the mismeasured variable, i.e., the noise-to-signal ratio $\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}}$ to be moderate, e.g., $\tau\lesssim 0.5$ (we later discuss how the appropriateness of this assumption can be checked).
+To provide a practical estimation approach for the general class of models (1), we focus on empirical settings in which the researcher believes the variability of the measurement error to be at most a fraction of the variability of the mismeasured variable, i.e., the noise-to-signal ratio $\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}}$ to be moderate.
 The absolute magnitude of the measurement error $\sigma_{\varepsilon}$ does not need to be small.
+Existing validation studies provide insights into the magnitude of $\tau$ for some key economic variables and datasets.
+BoundKrueger1991JoLaborEcon consider log-earnings in the Current Population Survey (CPS) data matched to the Social Security payroll records.
+Their estimates of the variance of the measurement errors correspond to $\tau$ of approximately $0.47$ and $0.30$ for the subsamples of men and women, respectively.
+BoundBrownDuncanRodgers1994JoLaborEcon consider a validation study of Panel Study of Income Dynamics (PSID).
+Their estimates imply $\tau=0.39-0.66$ for log-earnings and $\tau=0.63-0.76$ for hours worked.
+Pischke1995JBES estimates correspond to $\tau=0.39-0.50$ for the log-earnings in PSID.
+AshenfelterKrueger1994AER assess the mismeasurement in the years of education; their estimates correspond to $\tau=0.30-0.37$.
 Focusing on these settings allows us to isolate the most important aspects of the problem and, as result, to develop a simple estimator, which does not require any nonparametric estimation or simulation.
@@
 Although the concerns about measurement errors are often raised, the majority of applied work does not explicitly correct the EIV bias in nonlinear models, and instead implicitly or explicitly argues or conjectures that the EIV bias is likely not to be too large.
-See also the review of Bound et al.
-(2001).
 By considering $\tau$ as drifting towards zero with the sample size, our approach provides a better guidance on construction of EIV robust estimators with good finite sample properties when $\tau$ is small or moderate.
 [^4] Nonstandard asymptotic approximations with drifting parameters are often used to obtain better approximations of the finite sample behavior of estimators and tests.
-For example, in the instrumental variable regression settings, to consider the settings with relatively small first stage coefficients, Staiger and Stock (1997) model them as shrinking with $n$.
+For example, in the instrumental variable regression settings, to consider the settings with relatively small first stage coefficients, StaigerStock1997 model them as shrinking with $n$.
 It is important to keep in mind that such nonstandard asymptotic approximations are merely mathematical tools.
@@
 In particular, the simulation results show that despite the simplicity of implementation, the MERM estimators can compete with and outperform semi-nonparametric estimators.
-The MERM estimator is structurally different from the existing approaches that require nonparametric estimation of some nuisance parameters, for example, of the density $f_{X_{i}^{\ast}|W}$.
+The MERM estimator is structurally different from the existing approaches that require nonparametric estimation of some nuisance parameters, for example, of the density $f_{X^{\ast}|Z,W}$.
 Avoiding nonparametric estimation has at least two advantages.
-First, since the majority of empirical applications include at least a handful of additional covariates $W_{i}$, nonparametric estimation is often infeasible due to the curse of dimensionality.
+First, since the majority of empirical applications include additional covariates $W_{i}$, nonparametric estimation is often infeasible due to the curse of dimensionality.
 Because the MERM estimator does not involve any nonparametric estimation, it can be used in applications with a relatively large number of additional covariates $W_{i}$, and remains feasible even in the more complicated settings, including multi-equation and structural models, and applications with multiple mismeasured variables $X_{i}$.
@@
 In contrast, having a discrete instrument is sufficient for the MERM approach because the nuisance parameter $\gamma_{0}$ is finite-dimensional.
+For example, in Section 3.3 we consider estimation of the model of multinomial choice among three modes of transportation.
+A leading alternative approach to the errors-in-variables problem in this model is the semi-nonparametric sieve-MLE estimator advocated by HuSchennach2008Ecta; CarrollChenHu2010JoNS, among others.
+This approach requires, among other things, estimating the conditional density $f_{X^{\ast}|Z,W}$ of $X_{i}^{\ast}$ given the instrument $Z_{i}$ and covariates $W_{i}$.
+In this empirical example, $W_{i}$ includes four continuously distributed covariates (two continuously distributed characteristics per choice) and a discrete one, while scalar $X_{i}^{\ast}$ and $Z_{i}$ are also continuous.
+Thus, $f_{X^{\ast}|Z,W}$ is a function of six continuous and one discrete variable.
+Hence, for typical sample sizes, estimating $f_{X^{\ast}|Z,W}$ in this example is infeasible due to the curse of dimensionality.
+In contrast, as the results of Section 3.3 demonstrate, the MERM approach is practical and effective in this application, in part because it avoids estimation of the high-dimensional nuisance functions like $f_{X^{\ast}|Z,W}$ altogether.
 The simplicity and practicality of the MERM approach do come at a cost: there is a limit on the magnitude of the measurement errors it can handle.
@@
 We view the MERM approach as providing a bridge between the settings in which the measurement errors are guaranteed to be absent or negligible, and the settings where the measurement errors are so large that one has to use the relatively more complicated estimators from the earlier literature (if they exist at all for the model of interest).
-Related Literature Chen et al. (2011), Schennach (2016), and Schennach (2020) provide excellent overviews of the measurement error literature.
-The existing semiparametric approaches to estimation and inference in models with EIV involve nonparametric estimation of infinite-dimensional nuisance parameters (e.g., Chesher2000WP,Li2002JoE,Schennach2004Ecta,Schennach2007Ecta,HuSchennach2008Ecta,SchennachHu2013JASA,Song2015JoE), simulation (e.g., Schennach2014Ecta-ELVIS), or both (e.g., Newey2001REStat,WangHsiao2011JoE)))).
-The exceptions include models with linear and polynomial regression functions (see HINP1991JoE,HausmanNeweyPowell1995JoE)))), and Gaussian control variable models such as Probit and Tobit with endogeneity (see SmithBlundell1986Ecta,RiversVuong1988JoE).
+Related Literature ChenHongNekipelov2011JEL, Schennach2016AnnRev, and Schennach2020HB-ME provide excellent overviews of the measurement error literature.
+The existing semiparametric approaches to estimation and inference in models with EIV involve nonparametric estimation of infinite-dimensional nuisance parameters (e.g., Chesher2000WP,Li2002JoE,Schennach2004Ecta,Schennach2007Ecta,HuSchennach2008Ecta,SchennachHu2013JASA,Song2015JoE), simulation (e.g., Schennach2014Ecta-ELVIS), or both (e.g., Newey2001REStat,WangHsiao2011JoE)).
+The exceptions include models with linear and polynomial regression functions (see HINP1991JoE,HausmanNeweyPowell1995JoE)), and Gaussian control variable models such as Probit and Tobit with endogeneity (see SmithBlundell1986Ecta,RiversVuong1988JoE).
 To the best of our knowledge, this paper is the first to provide an approach for $\sqrt{n}$-consistent and asymptotically normal and unbiased estimation of general GMM models with EIV that does not require any nonparametric estimation (or simulation).
@@
 Modeling the variance of the measurement error as shrinking to zero with the sample size is a popular approach in Statistics.
-The method has been proposed by Wolter and Fuller (1982), who used it to construct an approximate MLE estimator of a nonlinear regression model with Gaussian errors.
+The method has been proposed by WolterFuller1982AS, who used it to construct an approximate MLE estimator of a nonlinear regression model with Gaussian errors.
 Following their approach, the Statistics literature has mainly focused on the settings where the moments of the EIV needed to bias correct the estimators are either known or can be directly estimated from the available data such as repeated measurements (e.g., CarrollStefanski1990JASA,CarrollEtAl2006Book-ME).
 In Economics, such data are relatively rare.
-The use of approximations with shrinking variance of measurement errors in Econometrics literature has been pioneered by Kadane (1971), Amemiya (1985), and Chesher (1991).
-Such approximations have been used to check the sensitivity of naive estimators to the EIV by considering how the estimates change as the unknown moments of the measurement errors vary within some set of plausible values, e.g., see Chesher and Schluter (2002), Chesher et al.
-(2002), Battistin and Chesher (2014), Chesher (2017), and Hong and Tamer (2003).
+The use of approximations with shrinking variance of measurement errors in Econometrics literature has been pioneered by Kadane1971Ecta, Amemiya1985JoE, and Chesher1991Biomet.
+Such approximations have been used to check the sensitivity of naive estimators to the EIV by considering how the estimates change as the unknown moments of the measurement errors vary within some set of plausible values, e.g., see ChesherSchluter2002ReStud, ChesherDumanganeSmith2002JoE, BattistinChesher2014JoE, Chesher2017JoE, and HongTamer2003JoE.
+BoundBrownMathiowetz2001HBoE review a broad list of validation studies matching standard economic dataset to administrative records.
+The estimates they report suggest that the measurement errors of moderate magnitude are typical for empirical applications.
+This suggests that the approach developed in this paper could prove valuable for a wide range of applied work.
 This paper differs from the earlier literature in several ways.
@@
 Third, the MERM approach allows the measurement errors to have larger magnitudes than most of the papers in the earlier literature; this is achieved by the MERM approach recursively bias correcting the bias correction terms.
-Fourth, our approach allows relaxing the assumption of classical measurement errors.
-The most widespread approach to identification of the EIV models in economic applications is to use instrumental variables, e.g., see Hausman et al.
-(1991); Newey (2001); Schennach (2007); Wang and Hsiao (2011).
-In a recent paper, Hahn et al.
-(2021) reconsider the regression model in Amemiya (1990) using a bias correction similar to ours.
-When proper excluded variables are not available, researchers have considered using higher moments of $X_{i}$ as instruments, e.g., see Reiersøl (1950); Lewbel (1997); Erickson and Whited (2002); Schennach and Hu (2013); Ben-Moshe et al.
-(2017).
-When available, repeated measurements can also be used to identify the model, e.g., see Hausman et al.
-(1991); Li and Vuong (1998); Li (2002); Schennach (2004).
+The most widespread approach to identification of the EIV models in economic applications is to use instrumental variables, e.g., see HINP1991JoE; Newey2001REStat; Schennach2007Ecta; WangHsiao2011JoE.
+In a recent paper, HahnHausmanKim2021EL reconsider the regression model in Amemiya1990JoE using a bias correction similar to ours.
+When proper excluded variables are not available, researchers have considered using higher moments of $X_{i}$ as instruments, e.g., see Reiersol1950Ecta; Lewbel1997Ecta; EricksonWhited2002ET; SchennachHu2013JASA; BenMosheDHaultfeuilleLewbel2017JoE.
+When available, repeated measurements can also be used to identify the model, e.g., see HINP1991JoE; LiVuong1998JoE; Li2002JoE; Schennach2004Ecta.
 The MERM estimator accommodates these identification approaches within a unified estimation framework.
@@
 We discuss identification in Section 2.4.
-In addition, in a related paper Evdokimov and Zeleneev (2022) study nonparametric regression with EIV using the $\tau\rightarrow 0$ approximation, and demonstrate that the MERM approach can also be motivated from a nonparametric perspective.
-Kitamura et al.
-(2013); Andrews et al.
-(2017); Armstrong and Kolesár (2021); Bonhomme and Weidner (2022), among others, develop tools for estimation and inference in GMM, which are robust to general perturbation or misspecification of the true data generating process.
+In addition, in a related paper EvdokimovZeleneev2022WP-NPID study nonparametric regression with EIV using the $\tau\rightarrow 0$ approximation, and demonstrate that the MERM approach can also be motivated from a nonparametric perspective.
+KitamuraOtsuEvdokimov2013Ecta; AndrewsGentzkowShapiro2017QJE; ArmstrongKolesar2021QE; BonhommeWeidner2021QE, among others, develop tools for estimation and inference in GMM, which are robust to general perturbation or misspecification of the true data generating process.
 They focus on the settings in which these perturbations are sufficiently small, so that naive estimators remain $\sqrt{n}$-consistent, and their biases are of the same order of magnitude as their standard errors.
@@
 The MERM approach also provides a useful foundation for dealing with EIV in more complicated settings.
-Evdokimov and Zeleneev (2018) utilize the MERM framework to address an issue of nonstandard inference, which turns out to arise generally when EIV models are identified using instrumental variables.
-Evdokimov and Zeleneev (2019) extend the analysis of this paper to long panel and network settings.
+EvdokimovZeleneev2018WP-Inference utilize the MERM framework to address an issue of nonstandard inference, which turns out to arise generally when EIV models are identified using instrumental variables.
+EvdokimovZeleneev2019WP-Panel extend the analysis of this paper to long panel and network settings.
 Organization of the paper Section 2 introduces the Moderate Measurement Error framework and the proposed MERM estimator.
@@
 Section 4 considers several extensions of the framework.
+A supplementary appendix contains all proofs and additional results for the numerical and empirical illustrations.
 2 Moderate Measurement Errors Framework
-To present the main ideas we first consider the case of univariate $X_{i}^{\ast}$ and classical measurement error $\varepsilon_{i}$.
-Later we consider multivariate $X_{i}^{\ast}$ and non-classical measurement errors $\varepsilon_{i}$.
+To present the main ideas we first consider the case of univariate $X_{i}^{\ast}$.
+We will consider multivariate $X_{i}^{\ast}$ later.
+We assume that the measurement error is classical, i.e., that $\varepsilon_{i}$ is independent of $X_{i}^{*}$ and $S_{i}$; later we will discuss how this assumption can be relaxed.
+Following the rest of the literature, we assume that $\mathbb{E}\left[\varepsilon_{i}\right]=0$.
+[^5] A location normalization such as $\mathbb{E}\left[\varepsilon_{i}\right]=0$ is usually necessary because it is not possible to separately identify the means $\mathbb{E}\left[X_{i}^{\ast}\right]$ and $\mathbb{E}\left[\varepsilon_{i}\right]$.
 To develop a practical estimation approach for general moment condition models we focus on the settings in which $\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}}$ is small or moderate.
@@
 Likewise, the extent of the EIV problem does not change with such rescaling.
-[…]
+For simplicity of exposition, it is convenient to assume that $X_{i}^{\ast}$ is scaled so that $\sigma_{X^{\ast}}$ is of order one and, correspondingly, the moments $\mathbb{E}[\left|\varepsilon_{i}\right|^{k}]\propto\tau_{n}^{k}$ decrease with $k$ when $\tau_{n}<1$.
+For example, this could be ensured by normalizing observed $X_{i}$ to have $\sigma_{X}=1$.
+Let us stress that this normalization is used only to simplify the exposition; as we show in Appendix G, the proposed MERM estimator does not require any normalizations in practice.
 2.1 Special Case: Quadratic Expansion
@@
 $=\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta)]+\dfrac{\mathbb{E}[\varepsilon_{i}^{2}]}{2}\mathbb{E}[g_{x}^{(2)}(X_{i}^{\ast},S_{i},\theta)]+O(\tau_{n}^{3}),$ (3)
-where the second equality holds because because $\varepsilon_{i}$ and $\left(X_{i}^{\ast},S_{i}\right)$ are independent, and $\mathbb{E}\left[\varepsilon_{i}\right]=0$.
-This expansion implies that $\mathbb{E}[g(X_{i},S_{i},\theta_{0})]=O\left(\sigma_{\varepsilon}^{2}\right)=O(\tau_{n}^{2})$.
+where the second equality holds because $\varepsilon_{i}$ and $\left(X_{i}^{\ast},S_{i}\right)$ are independent, and $\mathbb{E}\left[\varepsilon_{i}\right]=0$}.
+Evaluating the expansion above at $\theta=\theta_{0}$ gives $\mathbb{E}[g(X_{i},S_{i},\theta_{0})]=O\left(\sigma_{\varepsilon}^{2}\right)=O(\tau_{n}^{2})$, because $\mathbb{E}[g(X_{i}^{*},S_{i},\theta_{0})]=0$.
 As a result, a naive estimator that ignores the EIV and uses $X_{i}$ in place of $X_{i}^{\ast}$ has EIV bias of order $\tau_{n}^{2}$.
 [^6] For example, consider a linear regression with a scalar mismeasured regressor.
-The bias of the naive OLS estimator of the slope parameter $\theta_{01}$ is $-\theta_{01}\frac{\tau_{n}^{2}}{1+\tau_{n}^{2}}=-\theta_{01}\tau_{n}^{2}+O\left(\tau_{n}^{4}\right)$.
-Bias of the naive estimator should be compared with its standard error, which is of order $n^{-1/2}$.
-Bias of the naive estimator is not negligible, unless the measurement error is rather small (theoretically, unless $\tau_{n}^{2}=o\left(n^{-1/2}\right)$).
+The bias of the naive OLS estimator of the slope parameter $\theta_{01}$ is $\theta_{01}\frac{\tau_{n}^{2}}{1+\tau_{n}^{2}}=-\theta_{01}\tau_{n}^{2}+O\left(\tau_{n}^{4}\right)$.
+The bias of the naive estimator should be compared with its standard error, which is of order $n^{-1/2}$.
+Thus, the bias of the naive estimator is not negligible, unless the measurement error is rather small (theoretically, unless $\tau_{n}^{2}=o\left(n^{-1/2}\right)$).
 In particular, tests and confidence intervals based on the naive estimator are invalid and can provide highly misleading results.
@@
 The idea behind this substitution is that the bias of order $O(\tau_{n}^{2})$ in $\mathbb{E}[g_{x}^{(2)}(X_{i},S_{i},\theta)]$ can be ignored because it is multiplied by $E\left[\varepsilon_{i}^{2}\right]=O\left(\tau_{n}^{2}\right)$.
-[^7] Such substitutions of $X^{\ast}$ with $X$ have been used in other contexts, e.g., Chesher and Schluter (2002).
+[^7] Such substitutions of $X^{\ast}$ with $X$ have been used in other contexts, e.g., ChesherSchluter2002ReStud.
 With the substitution, we can rearrange equation (4) and write it as
@@
 Hence, the corrected moment conditions $\psi$ can be used to jointly estimate the true parameters $\theta_{0}$ and $\gamma_{02}$ by a GMM estimator.
-[^8] In the moment condition settings, having $o\left(n^{-1/2}\right)$ is equivalent to having 00 on the right-hand side of equation (8).
+[^8] In the moment condition settings, having $o\left(n^{-1/2}\right)$ is equivalent to having 0 on the right-hand side of equation (8).
 Remark 1.
@@
 $\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta)]=\mathbb{E}[g(X_{i},S_{i},\theta)]-\sum_{k=2}^{K}\frac{\mathbb{E}[\varepsilon_{i}^{k}]}{k!}\mathbb{E}\left[g_{x}^{(k)}(X_{i}^{\ast},S_{i},\theta)\right]+o(n^{-1/2}).$ (10)
-This equation is the general expansion analog of equation (4). The summation on the right hand side is the bias correction term, which we use to construct the MERM estimator.
+[…]
 It turns out that for $K\geq 4$, estimation of $\mathbb{E}[g_{x}^{(k)}(X_{i}^{\ast},S_{i},\theta)]$ is more intricate than in the case of $K=2$, and the substitution we made in equation (6) no longer works.
@@
 $\psi(X_{i},S_{i},\theta,\gamma)\equiv g(X_{i},S_{i},\theta)-\sum_{k=2}^{K}\gamma_{k}g_{x}^{(k)}(X_{i},S_{i},\theta),$ (11)
-where $\gamma=(\gamma_{2},\dots,\gamma_{K})^{\prime}$ is a $K-1$ dimensional vector of parameters.
+where $\gamma=(\gamma_{2},\dots,\gamma_{K})^{\prime}$ is a K1K-1 dimensional vector of parameters.
 Let $\gamma_{0}\equiv(\gamma_{02},\dots,\gamma_{0K})^{\prime}$ denote the vector of true parameters $\gamma_{0k}$, defined as
@@
 Assumption CME.
-$\varepsilon_{i}$ is independent from $(X_{i}^{\ast},S_{i})$ and $\mathbb{E}[\varepsilon_{i}]=0$.
-Assumption 2.2 is the classical measurement error assumption.
-We relax this assumption later in Section 4.2.
+(Classical Measurement Error) $\varepsilon_{i}$ is independent from $(X_{i}^{\ast},S_{i})$ and $\mathbb{E}[\varepsilon_{i}]=0$.
 The following lemma establishes validity of the corrected moment conditions under Assumptions 2.1, 2.2, and some mild regularity conditions provided in Appendix A.
@@
 For example, the overidentifying restrictions can be constructed by using an instrumental variable; we discuss this in more detail below.
+Remark 2.
+Construction of the corrected moment conditions $\psi$ requires the original moment function $g(x,s,\theta)$ to have a sufficient number of derivatives with respect to $x$.
+Thus, the proposed correction method does not apply to settings with non-differentiable moment functions, for example, those arising in the instrumental variable quantile regression (IVQR).
 2.3 Measurement Error Robust Moments (MERM) estimator
 The MERM estimator jointly estimates the parameters $\theta_{0}$ and $\gamma_{0}$ using moment conditions $\psi$. It is convenient to define the joint vector of parameters
-$\beta\equiv(\theta^{\prime},\gamma^{\prime})^{\prime},\ \beta_{0}\equiv(\theta_{0}^{\prime},\gamma_{0}^{\prime})^{\prime},\text{ }\hat{\beta}\equiv(\hat{\theta}^{\prime},\hat{\gamma}^{\prime})^{\prime},$
+[…]
 and the parameter space $\mathcal{B}\equiv\Theta\times\Gamma$, where $\Theta$ and $\Gamma$ are the parameter spaces for $\theta$ and $\gamma$.
@@
 Suppose that $\{(X_{i}^{\ast},S_{i}^{\prime},\varepsilon_{i})\}_{i=1}^{n}$ are i.i.d.. Then, under Assumptions 2.1, 2.2, 2.3, and A.1-A.3 in Appendix A,
-[…]
+$n^{1/2}\Sigma^{-1/2}(\hat{\beta}-\beta_{0})\overset{d}{\rightarrow}N(0,I_{\dim(\beta)}),\text{ where}$ (14)
 $\Sigma\equiv(\Psi^{\prime}\Xi\Psi)^{-1}\Psi^{\prime}\Xi\Omega_{\psi\psi}\Xi\Psi(\Psi^{\prime}\Xi\Psi)^{-1}.$
@@
 In other words, the proposed estimator can be simply treated as a standard GMM estimator based on the corrected moment conditions $\psi$, and the conventional standard errors, tests, and confidence intervals are valid.
-Remark 3.
-Researchers may be interested in average effects of the form $\lambda_{0}\equiv\mathbb{E}\left[\lambda\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\right]$.
-In the NLR, one may be interested in the average partial effect $x$ (i.e., $\lambda_{0}\equiv\mathbb{E}\left[\nabla_{x}\rho\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\right]$) or another covariate.
-The naive average partial effect estimator $\hat{\lambda}_{\text{Naive}}\equiv\frac{1}{n}\sum_{i=1}^{n}\lambda(X_{i},S_{i},\hat{\theta})$ suffers from the EIV bias, unless function $\lambda$ is linear in $X_{i}^{\ast}$.
-Instead, one can use estimates $\hat{\gamma}$ to construct the bias-corrected estimator $\hat{\lambda}_{\text{MERM}}\equiv\frac{1}{n}\sum_{i=1}^{n}\left\{\lambda(X_{i},S_{i},\hat{\theta})-\sum_{k=2}^{K}\hat{\gamma}_{k}\lambda_{x}^{(k)}(X_{i},S_{i},\hat{\theta})\right\}$.
-Remark 4.
 […]
 Remark 4.
-It is important to note that $\gamma_{0k}\neq\mathbb{E}\left[\varepsilon_{i}^{k}\right]\left/k!\right.$ for $k\geq 4$, contrary to what equation (10) might suggest.
-For example, $\gamma_{04}=\left(\mathbb{E}\left[\varepsilon_{i}^{4}\right]-6\sigma_{\varepsilon}^{4}\right)\left/24\right.$ is negative for many distributions, including normal.
-For instance, in the example of Remark 4, $\gamma_{04}\approx-0.0026$.
-The reason that generally $\gamma_{0k}\neq\mathbb{E}\left[\varepsilon_{i}^{k}\right]\left/k!\right.$ is that the estimators of the correction terms themselves need a correction, which is accounted for by the form of $\gamma_{0k}$.
+The standard $J$-test of overidentifying restrictions remains valid in the MERM settings, and can be used to check the model specification.
+The $J$-test jointly tests the following hypotheses: (i) $K$ is sufficiently large to correct the EIV bias; (ii) assumptions on the EIV are valid; and (iii) the original moment conditions $g$ are correctly specified so equation (1) holds, i.e., that the original economic model is correctly specified aside from the presence of the EIV in $X_{i}$.
+Thus, if the $J$-test rejects the validity of the corrected moment conditions $\psi$, the researcher might want to (i) consider taking a larger $K$; (ii) employ a different correction method; or (iii) consider an alternative specification of the original moments $g$.
+Remark 5.
+Considering larger $K$ allows for $\tau_{n}\$converging to zero at a slower rate, which in finite samples corresponds to the asymptotics providing better approximations for larger magnitudes of measurement errors.
+On the other hand, taking a larger $K$ increases the dimension of the nuisance parameter $\gamma_{0}$ and thus typically increases the variance of $\hat{\theta}$.
+We consider this issue in more detail and provide a data-driven method for choosing $K$ in Section 4.1.
+Remark 6.
+The MERM framework can be extended to the case of non-classical measurement errors; see EvdokimovZeleneev2022WP-NPID for details and a fully nonparametric analysis.
+In this paper, we focus on the classical measurement errors, developing a practical bias correction approach, which can be easily implemented in a wide range of economic applications.
+Even when Assumption 2.2 is violated, the deviations from it are often limited in magnitude, so the corrections based on the 2.2 assumption remove most of the EIV bias.
+Thus, in practice, using the estimator designed for classical measurement errors is typically preferable to ignoring mismeasurement altogether.
+Remark 7.
+It is important to note that $\gamma_{0k}\neq\mathbb{E}\left[\varepsilon_{i}^{k}\right]\left/k!\right$.
+for $k\geq 4$, contrary to what equation (10) might suggest.
+For example, $\gamma_{04}=\left(\mathbb{E}\left[\varepsilon_{i}^{4}\right]-6\sigma_{\varepsilon}^{4}\right)\left/24\right$.
+is negative for many distributions, including normal.
+The reason that generally $\gamma_{0k}\neq\mathbb{E}\left[\varepsilon_{i}^{k}\right]\left/k!\right$.
+is that the estimators of the correction terms themselves need a correction, which is accounted for by the form of $\gamma_{0k}$.
 Since there is a one-to-one relationship between $\gamma_{0}$ and the moments $\mathbb{E}\left[\varepsilon_{i}^{\ell}\right]$, parameter space $\Gamma$ for $\gamma_{0}$ can incorporate restrictions that the moments must satisfy (e.g., $\sigma_{\varepsilon}^{2}\geq 0$ and $\mathbb{E}\left[\varepsilon_{i}^{4}\right]\geq\sigma_{\varepsilon}^{4}$).
@@
 Remark 9.
-Considering larger $K$ allows $\tau_{n}$converging to zero at a slower rate, which in finite samples corresponds to the asymptotics providing better approximations for larger magnitudes of measurement errors.
-On the other hand, taking a larger $K$ increases the dimension of the nuisance parameter $\gamma_{0}$ and thus typically increases the variance of $\hat{\theta}$.
-Remark 8.
 The formulas of the derivatives $g_{x}^{(k)}(\cdot)$ are typically easy to compute analytically or using symbolic algebra software.
 Alternatively, these derivatives can be computed using numerical differentiation.
-In either way, the corrected set of moments can be automatically produced for a generic moment function $g(\cdot)$ provided by the user.
+Thus, the corrected set of moments can be automatically produced for a generic moment function $g(\cdot)$ provided by the user.
 2.4 Model Identification: Jacobian $\Psi$
@@
 for some moment function $u$.
-For example, consider the nonlinear regression model $\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=x\right]=\rho\left(x,\theta_{0}\right)$, then $u\left(x,y,\theta\right)=\rho\left(x,\theta\right)-y$ [^9] For simplicity of the exposition, in the expectation in equation (15) we only condition on $X_{i}^{\ast}$.
+For example, consider the nonlinear regression model $\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=x\right]=\rho\left(x,\theta_{0}\right)$, then $u\left(x,y,\theta\right)=\rho\left(x,\theta\right)-y$.
+[^10] For simplicity of the exposition, in the expectation in equation (15) we only condition on $X_{i}^{\ast}$.
 The discussion applies in a straightforward way to the settings with additional correctly measured variables $W_{i}$ in the conditioning set, i.e., the model $\mathbb{E}\left[u\left(X_{i}^{\ast},S_{i},\theta_{0}\right)|X_{i}^{\ast},W_{i}\right]=0$.
@@
 We attempt to address this concern and make the above intuitions more precise in two ways.
-First, in the following subsection we consider a specific (but frequently employed) kind of an instrument: a second measurement (possibly nonclassical).
+First, in the following subsection we consider a specific (but frequently employed) kind of an instrument: a second measurement (possibly non-classical).
 The specific form of the excluded variable allows us to provide more transparent identification conditions.
-Second, in Evdokimov and Zeleneev (2022) we study nonparametric regression model with EIV using the $\tau_{n}\rightarrow 0$ approximation.
+Second, in EvdokimovZeleneev2022WP-NPID we study nonparametric regression model with EIV using the $\tau_{n}\rightarrow 0$ approximation.
 We show that the model is identified using an instrument (even a discrete one), and motivate the MERM approach from a nonparametric perspective.
@@
 Here the total number of moments is $m=2J+1$. The additional $J$ moments added in equation (16) use $Q_{i}$, which will allow identifying $\gamma_{0}=\mathbb{E}[\varepsilon_{i}^{2}]/2$.
-It turns out that in these settings there is a simple sufficient condition for Assumption 2.3(i) to hold. Appendix D demonstrates that $\Psi^{\ast}$ will have full rank if
+It turns out that in these settings there is a simple sufficient condition for Assumption 2.3(i) to hold. Appendix E demonstrates that $\Psi^{\ast}$ will have full rank if
 $\mathbb{E}\left[u_{x}^{\left(1\right)}\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\times\left(1,X_{i}^{\ast},\ldots,\left(X_{i}^{\ast}\right)^{J-1}\right)^{\prime}\right]\neq 0.$ (17)
@@
 For example, in the nonlinear regression model $u\left(x,y,\theta\right)=\rho\left(x,\theta\right)-y$, and condition (17) is satisfied as long as $\mathbb{E}\left[\rho_{x}^{(1)}\left(X_{i}^{\ast},\theta_{0}\right)\left(X_{i}^{\ast}\right)^{j}\right]\neq 0$ for some $j\in\left\{0,\ldots,J-1\right\}$.
-2.5 Assessing model validity
-The standard $J$-test of overidentifying restrictions remains valid in the MERM settings, and can be used to check the model specification.
-The $J$-test jointly tests the following hypotheses: (i) $K$ is sufficiently large to correct the EIV bias; (ii) assumptions on the EIV are valid; and (iii) the original moment conditions $g$ are correctly specified so equation (1) holds, i.e., that the original economic model is correctly specified aside from the presence of the EIV in $X_{i}$.
-Issue (i) can be address by taking a larger $K$.
-In Section 4 we extend the framework to obtain corrected moments that are valid under weaker assumptions on the EIV, which can help addressing issue (ii).
-It is possible to obtain an explicit bound on the higher-order EIV bias of the MERM estimator.
-Suppose the researcher is interested in estimating a linear combination of parameters $v^{\prime}\beta_{0}$ for some non-zero $v\in\mathbb{R}^{\dim(\beta)}$, e.g., the first component of $\theta_{0}$.
-Since $\hat{\beta}$ is a standard GMM estimator,
-$\text{Bias}(v^{\prime}\hat{\beta})\approx-v^{\prime}(\Psi^{\prime}\Xi\Psi)^{-1}\Psi^{\prime}\Xi\;\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right].$ (18)
-Here $\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right]\approx\gamma_{0\overline{K}}\mathbb{E}[g_{x}^{(\overline{K})}(X_{i},S_{i},\theta_{0})]$ with $\overline{K}=K+2$ if $K$ is even and $\varepsilon_{i}$ is symmetric, and $\overline{K}=K+1$ otherwise.
-Thus,
-$\text{Bias}(v^{\prime}\hat{\beta})\approx\gamma_{0\overline{K}}b_{v},$
-where $b_{v}$ can be easily estimated.
-The value of $\gamma_{0\overline{K}}$ is given by the recursive relationship (12) and determined by the unknown $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ (and by the lower order moments of $\varepsilon_{i}$, which can be recovered from $\hat{\gamma}$ with a sufficient precision).
-While $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ and, hence, $\gamma_{0\overline{K}}$ are generally unknown, one simple way to evaluate the magnitude of $\text{Bias}(v^{\prime}\hat{\beta})$ is to calculate $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ using the estimated $\sigma_{\varepsilon}^{2}$ and assuming normality of $\varepsilon$.
-Alternatively, one can also bound $\gamma_{0\overline{K}}$ and, subsequently, $\text{Bias}(v^{\prime}\hat{\beta})$ by imposing some (a priori) bounds on $\mathbb{E}[(\varepsilon_{i}/\sigma_{\varepsilon})^{\overline{K}}]$.
-The bounds on Bias$(v^{\prime}\hat{\beta})$ can also be used to assess the appropriateness of Assumption 2.1 and the reliability of the inference based on the normal approximation in equation (14).
-If the (worst case) bias of $v^{\prime}\hat{\beta}$ is substantial, relative to its standard error, the researcher may want to account for this and consider choosing a higher $K$.
-Appendix E provides a further discussion of this issue and additional details on the calculation of Bias$(v^{\prime}\hat{\beta})$.
 3 Numerical Evidence
 3.1 Comparison with a Semi-Nonparametric Estimation Approach
-We compare MERM estimator with the state-of-the-art semiparametric estimator of Schennach (2007, henceforth S07) for nonlinear regression models.
+We compare MERM estimator with the state-of-the-art semiparametric estimator of Schennach2007Ecta for nonlinear regression models.
 The Monte Carlo designs are taken from S07, and include a polynomial, rational fraction, and Probit nonlinear regression models.
@@
 The conditional expectation function $\rho$, the true value of the parameter of interest $\theta_{0}$, and the conditional distribution of the regression error $U_{i}$ are design-specific and reported in Tables 1-3 below.
-In all designs, $\tau=\sigma_{\varepsilon}/\sigma_{X}^{*}\approx 0.45$, so the measurement error is “fairly large” (Schennach, 2007).
+In all designs, $\tau=\sigma_{\varepsilon}/\sigma_{X}^{*}\approx 0.45$, so the measurement error is “fairly large” (Schennach2007Ecta).
 We report simulation results for the MERM estimator considering correction schemes with $K=2$ and $K=4$. The original moment function is
@@
 S07 -0.05 -0.07 -0.02 0.05 0.17 0.19 0.24 0.05 0.17 0.20 0.24 0.07 0.36
-$K=2$0.00 0.10 0.00 0.00 0.10 0.23 0.10 0.08 0.10 0.25 0.10 0.08 0.29
-$K=4$0.00 0.00 0.00 0.02 0.09 0.21 0.10 0.08 0.09 0.21 0.10 0.08 0.27
+$K=2$ -0.00 0.10 0.00 0.00 0.10 0.23 0.10 0.08 0.10 0.25 0.10 0.08 0.29
+$K=4$ -0.00 0.00 0.00 0.02 0.09 0.21 0.10 0.08 0.09 0.21 0.10 0.08 0.27
   • The DGP is as in (18) with $\rho(x,\theta)=\theta_{1}+\theta_{2}x+\theta_{3}x^{2}+\theta_{4}x^{3}$, $\theta_{0}=(1,1,0,-0.5)^{\prime}$, and $U_{i}\sim N(0,1/4)$.
@@
 S07 0.107 0.117 -0.150 0.146 0.139 0.328 0.181 0.182 0.361 0.443
-$K=2$0.004 -0.018 0.014 0.062 0.026 0.139 0.062 0.032 0.139 0.156
+$K=2$ -0.004 -0.018 0.014 0.062 0.026 0.139 0.062 0.032 0.139 0.156
 $K=4$ 0.014 -0.002 -0.024 0.062 0.031 0.154 0.063 0.031 0.156 0.171
@@
 $K=2$ 0.11 -0.31 0.18 0.34 0.21 0.46 0.51
-$K=4$0.01 -0.01 0.23 0.42 0.23 0.42 0.48
-  • The DGP is as in (19) with $\rho(x,\theta)=\frac{1}{2}(1+\text{erf}(\theta_{1}+\theta_{2}x))$, $\theta_{0}=(-1,2)^{\prime}$, and $U_{i}=1-\rho(X_{i}^{*},\theta_{0})$ with probability $\rho(X_{i}^{*},\theta_{0})$ and $-\rho(X_{i}^{*},\theta_{0})$ otherwise.
+$K=4$ -0.01 -0.01 0.23 0.42 0.23 0.42 0.48
+  • The DGP is as in (18) with $\rho(x,\theta)=\frac{1}{2}(1+\text{erf}(\theta_{1}+\theta_{2}x))$, $\theta_{0}=(-1,2)^{\prime}$, and $U_{i}=1-\rho(X_{i}^{*},\theta_{0})$ with probability $\rho(X_{i}^{*},\theta_{0})$ and $\rho(X_{i}^{*},\theta_{0})$ otherwise.
 3.2 Estimation and Inference in a Multinomial Choice Model
@@
 Remarkably, the corresponding finite sample null rejection probabilities remain close to the nominal $5\%$ rate even when the standard deviation of the measurement error is as large as $75\%$ of the standard deviation of the mismeasured $X^{*}$.
+To further check the limits of applicability of our method, we also consider larger values of $\tau\in\{1,3/2,2\}$.
+The numerical results analogous to the ones reported in Table LABEL:tab:_MC_mult_logit_2 are provided in Table 6 in Appendix F.1.
+Specifically, we find that inference results based on the correction scheme with $K=4$ remain accurate even for $\tau=1$.
+Unsurprisingly, inference becomes less reliable for bigger $\tau=3/2$ and $\tau=2$.
+At the same time, while the $K=4$ correction scheme fails to entirely eliminate the EIV bias in these designs, it still removes a big fraction of the bias and greatly improves on MLE in terms of the RMSE.
+Thus, while inference based on our estimator might be less reliable in extreme settings when the measurement error overwhelms the signal, the MERM estimator using $K=4$ appears to be sufficiently accurate over a wide range of $\tau$.
 3.3 Empirical Illustration: Choice of Transportation Mode
-In this section, we illustrate the finite sample properties of the MERM estimator in the context of a classical multinomial choice application: choice of transportation mode (e.g., McFadden, 1974).
+In this section, we illustrate the finite sample properties of the MERM estimator in the context of a classical multinomial choice application: choice of transportation mode (e.g., McFadden1974JPubE).
 To calibrate the numerical experiment, we use the ModeCanada dataset, a survey of business travelers for the Montreal-Toronto corridor.
@@
 The true value of $\theta_{0}$ is set to be the MLE estimate based on the original dataset.
-More details about this numerical experiment are given in Appendix H.
+More details about this numerical experiment are given in Appendix I.
 To evaluate the performance of the MERM estimator in these settings, we generate mismeasured $Income_{i}=Income_{i}^{*}+\varepsilon_{i}$.
@@
 $\tau=1/4$
-$\partial\ln p_{1}/\partial\ln I$0.07 0.12 0.14 9.00 0.01 0.14 0.14 5.68 0.02 0.19 0.19 7.02
+$\partial\ln p_{1}/\partial\ln I$ -0.07 0.12 0.14 9.00 0.01 0.14 0.14 5.68 0.02 0.19 0.19 7.02
 $\partial\ln p_{2}/\partial\ln I$ 0.03 0.07 0.08 5.84 -0.00 0.08 0.08 5.68 -0.01 0.10 0.10 6.40
@@
 $\tau=1/2$
-$\partial\ln p_{1}/\partial\ln I$0.24 0.11 0.27 61.84 -0.05 0.14 0.15 6.96 0.02 0.21 0.21 6.16
+$\partial\ln p_{1}/\partial\ln I$ -0.24 0.11 0.27 61.84 -0.05 0.14 0.15 6.96 0.02 0.21 0.21 6.16
 $\partial\ln p_{2}/\partial\ln I$ 0.09 0.07 0.11 24.76 0.02 0.09 0.09 5.96 -0.01 0.10 0.10 6.16
@@
 $\tau=3/4$
-$\partial\ln p_{1}/\partial\ln I$0.43 0.09 0.44 99.50 -0.19 0.14 0.24 27.46 0.02 0.22 0.22 5.84
+$\partial\ln p_{1}/\partial\ln I$ -0.43 0.09 0.44 99.50 -0.19 0.14 0.24 27.46 0.02 0.22 0.22 5.84
 $\partial\ln p_{2}/\partial\ln I$ 0.16 0.06 0.17 71.88 0.07 0.08 0.11 13.78 -0.01 0.11 0.11 6.32
@@
 4 Extensions
+4.1 Data-driven choice of $K$
+Making an appropriate choice of the expansion order $K$ is important for the estimation procedure.
+One has to be cautious not to take $K$ too small, as this may result in an estimator that only partially removes the EIV bias.
+On the other hand, picking a larger $K$ than needed might inflate standard errors and result in less powerful inference.
+In this section, we address this issue by providing a data-dependent procedure for selecting $K$.
+We demonstrate that our procedure has desirable theoretical properties.
+We also find that the procedure has good finite sample properties in a set of Monte Carlo simulation experiments across different values of $\tau$ and sample sizes.
+Consider two alternative values of the expansion order: $L$ and $K$, where $2\leq L<K$.
+In practice, even-order biases tend to dominate, so to reduce the set of choices, it is useful to focus on even values of the expansion orders $L$ and $K$.
+For example, one would typically be interested in choosing between $L=2$ and $K=4$.
+Let $\hat{\beta}_{L}$ and $\hat{\beta}_{K}$ denote the corresponding MERM estimators.
+The estimator $\hat{\beta}_{L}$ should be preferred as having smaller asymptotic variance provided that its remaining EIV bias is negligible relative to its standard error.
+Otherwise, the more conservative $\hat{\beta}_{K}$ should be used instead.
+Note that Lemma 1 suggests that the remaining asymptotic bias of $\hat{\beta}_{L}$ (due to the additional terms accounted for when the expansion of higher order $K$ is used) is given by (up to an $o(n^{-1/2})$ remainder)
+$\text{AsyB}(\hat{\beta}_{L})\equiv B\sum_{k=L+1}^{K}\gamma_{0k}\mathbb{E}[g_{x}^{(k)}(X_{i},S_{i},\theta_{0})],\quad B\equiv-(\Psi^{\prime}\Xi\Psi)^{-1}\Psi^{\prime}\Xi,$
+where matrix $B$ is based on the moments used for estimation of $\hat{\beta}_{L}$.
+Importantly, $\gamma_{0k}=O\left(\sigma_{\varepsilon}^{k}\right)$, and hence for $k>2$ we can estimate a bound on $\sqrt{n}\gamma_{0k}$ sufficiently quickly to provide a valid procedure for choosing $K$.
+To this end, we first estimate the model using the larger $K$.
+Let $\hat{\beta}_{K}=(\hat{\theta}^{\prime},\hat{\gamma}^{\prime})^{\prime}$ and $\hat{\sigma}_{\varepsilon}^{2}\equiv 2\hat{\gamma}_{2}$, where we dropped the additional subscripts $K$ for notation simplicity.
+Then, we can estimate $\sigma_{\varepsilon}^{K}$ by $\hat{\sigma}_{\varepsilon}^{K}\equiv(\hat{\sigma}_{\varepsilon}^{2})^{K/2}$.
+Using $\hat{\sigma}_{\varepsilon}^{2}=\sigma_{\varepsilon}^{2}+O_{p}(n^{-1/2})$, in the appendix we show that
+$\sqrt{n}\left(\hat{\sigma}_{\varepsilon}^{K}-\sigma_{\varepsilon}^{K}\right)=O_{p}\left(\sigma_{\varepsilon}^{K-2}+n^{-\left(K-2\right)/4}\right)=o_{p}\left(1\right).$ (19)
+Next, consider a sequence $\varkappa_{n}\rightarrow 0$, which we will specify precisely later, and let
+$\delta_{n}=1\left\{\max_{1\leq\ell\leq\dim(\beta)}\left|\hat{\Sigma}_{\ell\ell}^{-1/2}\sqrt{n}\hat{\sigma}_{\varepsilon}^{K}\hat{B}_{\ell\cdot}\overline{g}_{x}^{(K)}(\hat{\theta})\right|\leq c_{K}\varkappa_{n}\right\},$ (20)
+where $\hat{\Sigma}$ and $\hat{B}$ are consistent estimators of the asymptotic variance of $\hat{\beta}_{L}$ and of $B$, with $\hat{\Sigma}_{\ell\ell}$ denoting its $\ell$-th diagonal element, $\hat{B}_{\ell\cdot}$ denoting its $\ell$-th row, ${\overline{g}_{x}^{(K)}(\hat{\theta})\equiv n^{-1}\sum_{i=1}^{n}g^{(K)}(X_{i},S_{i},\hat{\theta})}$, and $c_{K}>0$ is a constant that we will calibrate below after we state the main theoretical result of this section.
+If $\delta_{n}=1$, the researchers should select $\hat{\beta}_{L}$.
+Otherwise, $\hat{\beta}_{K}$ should be used.
+The following lemma demonstrates that the proposed selection procedure has desirable theoretical properties.
+Lemma 3.
+Suppose the hypotheses of Theorem 2 hold for some even $K\geq 4$, and consider $L$ satisfying $2\leq L<K$.
+Suppose $\varkappa_{n}n^{\left(K-L-1\right)/\left(2L+2\right)}\rightarrow 0$.
+Then $\sqrt{n}\text{AsyB}(\hat{\beta}_{L})\delta_{n}=o_{p}\left(1\right)$.
+Moreover, consider $\varkappa_{n}=n^{-\left(K-L-1\right)/\left(2L+2\right)}\left(\ln n\right)^{-a}$ for any $a>0$.
+Then the criterion is consistent, in the sense that if $\tau_{n}=o(n^{-\frac{1}{2L+2}-\epsilon})$ for any $\epsilon>0$, the criterion will choose $\hat{\beta}_{L}$ with probability approaching one.
+The first part of Lemma 3 shows that, provided that $\varkappa_{n}$ goes to zero sufficiently fast, $\hat{\beta}_{L}$ is selected (i.e., $\delta_{n}=1$) only when its asymptotic bias is negligible.
+Next, note that $\hat{\beta}_{L}$ is asymptotically unbiased as long as $\tau_{n}=(n^{-\frac{1}{2L+2}})$.
+The second part of the lemma shows that, for the suggested choices of $\varkappa_{n}$, the criterion is non-vacuous, i.e., that it does select the MERM estimator with a smaller expansion order $L$ when this is appropriate.
+Typically, one would pick L=K2L=K-2, so the lemma suggests taking $\varkappa_{n}=n^{-1/\left(2K-2\right)}\left(\ln n\right)^{-a}$ in this case.
+In practice, it is important to pick an appropriate constant $c_{K}$ used in the construction of $\delta_{n}$ in equation (20).
+When constructing $\delta_{n}$, we used $\left|\gamma_{0K}\right|\propto\sigma_{\varepsilon}^{K}$, which motivates choosing $c_{K}=\sigma_{\varepsilon}^{K}/\left|\gamma_{0K}\right|$.
+It is convenient to use a rule-of-thumb approach, using a reference distribution to determine $c_{K}$.
+It turns out that the normal distribution is not only convenient, but also sufficiently conservative (note that the bigger $\gamma_{0K}$ is, the smaller $c_{K}$ is, resulting in a more conservative selection procedure).
+For example, suppose $K=4$, and consider using Student’s $t(\nu)$ distribution as the reference distribution.
+Then, for $\nu\geq 5$, the biggest $\left|\gamma_{04}\right|/\sigma_{\varepsilon}^{4}$ and the smallest $c_{4}$ correspond to $\nu=\infty$ matching the normal distribution.
+[^11] Note that Student’s $t(\nu)$ distribution does not have a finite $5$-th moment $\nu\leq 5$.
+Since for the normal distribution we have $\gamma_{0K}=\sigma_{\varepsilon}^{K}/K!!$ for even $K$, we recommend using $c_{K}=K!!$ in equation (20).
+Finally, while we recommend using normal distribution as the reference distribution, we also stress that the results of Lemma 3 hold even if the measurement error is not normal or if it is skewed.
+We summarize the proposed procedure in the following suggested algorithm.
+Algorithm 1.
+  1. 1.
+    Pick a large enough even $K\geq 4$ and compute $\hat{\beta}_{K}=(\hat{\theta}^{\prime},\hat{\gamma}^{\prime})^{\prime}$ and $\hat{\sigma}_{\varepsilon}^{2}=2\hat{\gamma}_{2}$.
+  2. 2.
+    Compute $\delta_{n}$ in (20) with L=K2L=K-2, $c_{K}=K!!$, and $\varkappa_{n}=\left(n\ln n\right)^{-1/\left(2K-2\right)}$.
+  3. 3.
+    Pick the length of expansion L=K2L=K-2 if $\delta_{n}=1$, otherwise keep the initial $K$.
+This procedure can be iterated if desired.
+To illustrate the performance of the algorithm provided above, we revisit the numerical experiment considered in Section 3.2.
+We focus on choosing between $K=4$ and $L=2$, and, as in Section 3.2, we report results for $n=2000$ in Table 5 below.
+To ensure that the proposed algorithm performs well in a variety of sample sizes, we also consider $n=1000$ and $n=4000$ and report the corresponding results in Tables 7 and 8 in Appendix F.2.
+We report the finite sample bias and RMSE for the naive MLE estimator, as well as for the MERM estimators using $K=2$ and $K=4$, and for the adaptive MERM estimator using data-driven $K$ following Algorithm 1.
+Notice that for the considered sample sizes, $K=2$ is preferred when $\tau=1/4$, and $K=4$ is preferred when $\tau=3/4$.
+We find that in both of these regimes and for all sample sizes, the adaptive estimator using data-driven $K$ is essentially equivalent to the preferred estimators, i.e., our procedure selects the appropriate $K$.
+Interestingly, in the intermediate regime with $\tau=1/2$, the adaptive estimator also has the smallest bias among the considered estimators, coming at the cost of a slightly bigger RMSE compared to the MERM estimator using $K=4$.
+Thus, the considered numerical experiment suggests that our algorithm has good finite sample properties supporting the findings of Lemma 3.
+Table 5: Choice of $K$ simulation results for the multinomial logit model, $n=2000$
+MLE $K=2$ $K=4$ data-driven $K$
+bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$
+$\tau=1/4$
+$\partial p_{1}/\partial x$ -3.24 3.51 0.74 2.74 1.13 2.95 0.75 2.75
+$\partial p_{1}/\partial w_{1}$ 2.32 2.84 -0.11 2.30 -0.31 2.31 -0.11 2.30
+$\partial p_{1}/\partial w_{2}$ 0.48 0.90 -0.04 0.87 -0.08 0.87 -0.04 0.87
+$\partial p_{2}/\partial x$ 1.96 2.28 -0.40 1.92 -0.63 2.03 -0.41 1.93
+$\partial p_{2}/\partial w_{1}$ -1.16 1.42 0.06 1.15 0.15 1.16 0.06 1.15
+$\partial p_{2}/\partial w_{2}$ -0.96 1.78 0.09 1.74 0.17 1.75 0.09 1.74
+$\partial p_{0}/\partial x$ 1.28 1.63 -0.34 1.47 -0.50 1.56 -0.34 1.48
+$\partial p_{0}/\partial w_{1}$ -1.16 1.42 0.06 1.15 0.15 1.16 0.06 1.15
+$\partial p_{0}/\partial w_{2}$ 0.48 0.88 -0.05 0.87 -0.09 0.88 -0.05 0.87
+$\tau=1/2$
+$\partial p_{1}/\partial x$ -8.97 9.04 -1.69 3.10 0.97 3.05 0.88 3.15
+$\partial p_{1}/\partial w_{1}$ 6.44 6.62 1.39 2.81 -0.21 2.42 -0.15 2.48
+$\partial p_{1}/\partial w_{2}$ 1.28 1.47 0.29 0.94 -0.06 0.92 -0.05 0.93
+$\partial p_{2}/\partial x$ 5.22 5.31 1.05 2.16 -0.53 2.15 -0.48 2.20
+$\partial p_{2}/\partial w_{1}$ -3.21 3.30 -0.69 1.40 0.10 1.21 0.08 1.24
+$\partial p_{2}/\partial w_{2}$ -2.52 2.88 -0.56 1.87 0.13 1.84 0.11 1.85
+$\partial p_{0}/\partial x$ 3.75 3.85 0.64 1.59 -0.44 1.65 -0.40 1.68
+$\partial p_{0}/\partial w_{1}$ -3.23 3.32 -0.70 1.41 0.10 1.21 0.08 1.24
+$\partial p_{0}/\partial w_{2}$ 1.23 1.41 0.28 0.93 -0.07 0.92 -0.06 0.93
+$\tau=3/4$
+$\partial p_{1}/\partial x$ -13.35 13.38 -6.83 7.32 0.71 3.29 0.71 3.29
+$\partial p_{1}/\partial w_{1}$ 9.69 9.80 4.95 5.61 0.01 2.62 0.01 2.62
+$\partial p_{1}/\partial w_{2}$ 1.81 1.94 1.01 1.35 -0.01 0.98 -0.01 0.98
+$\partial p_{2}/\partial x$ 7.48 7.52 4.06 4.45 -0.37 2.35 -0.37 2.35
+$\partial p_{2}/\partial w_{1}$ -4.83 4.88 -2.47 2.81 -0.01 1.31 -0.01 1.31
+$\partial p_{2}/\partial w_{2}$ -3.51 3.76 -1.99 2.66 0.03 1.97 0.03 1.97
+$\partial p_{0}/\partial x$ 5.87 5.92 2.77 3.14 -0.34 1.80 -0.34 1.80
+$\partial p_{0}/\partial w_{1}$ -4.87 4.93 -2.48 2.81 -0.01 1.31 -0.01 1.31
+$\partial p_{0}/\partial w_{2}$ 1.70 1.82 0.98 1.31 -0.02 0.99 -0.02 0.99
+  • This table reports the simulated finite sample bias and RMSE of the MLE and the MERM estimators for the partial derivatives $\partial p_{j}(x,w,\theta_{0})/\partial x$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{1}$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{2}$ for $j\in\{1,2,0\}$ evaluated at the population mean.
+The true values of the marginal effects are $(\partial p_{1}/\partial x,\partial p_{2}/\partial x,\partial p_{0}/\partial x)=(0.222,-0.111,-0.111)$ and zeros for the rest.
+The results are based on 5,000 replications.
 4.2 Multiple Mismeasured Variables
@@
 $\blacksquare$
-4.2 Weakly Classical and Non-Classical Measurement Errors
-In some applications, assuming that the measurement error is classical, i.e., $\varepsilon_{i}$ is independent from $(X_{i}^{*},S_{i})$, may be too restrictive.
-In this section, we demonstrate how the MERM framework can be used to address this issue.
-To address general non-classical measurement errors it is convenient to proceed in two steps.
-First, in Section 4.2.1, we consider estimation of models with weakly classical measurement errors satisfying $\mathbb{E}\left[\varepsilon_{i}|X_{i}^{\ast},S_{i}\right]=0$.
-This restriction on $\varepsilon_{i}$ is much weaker than the classical measurement error assumption.
-For example, the variance of the measurement error $\mathbb{E}\left[\varepsilon_{i}^{2}|X_{i}^{\ast},S_{i}\right]$ (and other moments) could depend on $X_{i}^{\ast}$ and $S_{i}$.
-Second, in Section 4.2.2, we consider estimation with general non-classical measurement errors that may be correlated with the mismeasured variable.
-4.2.1 Weakly Classical Measurement Errors
-In this section, we relax Assumption 2.2 and assume that the measurement error is weakly classical.
-Specifically, assume that $\mathbb{E}\left[\varepsilon_{i}|X_{i}^{*},S_{i}\right]=0$ or, equivalently, $\mathbb{E}\left[X_{i}|X_{i}^{*},S_{i}\right]=X_{i}^{*}$.
-Suppose that $\mathbb{E}\left[\left|\varepsilon_{i}\right|^{k}|X_{i}^{*},S_{i}\right]=O(\tau_{n}^{k})$ for $k\in\{2,\ldots,K\}$.
-Conditional on $X_{i}^{*}$ and $S_{i}$, the moment function can be expanded as
-$\mathbb{E}[g(X_{i},S_{i},\theta)|X_{i}^{\ast},S_{i}]={g(X_{i}^{\ast},S_{i},\theta)}+\sum_{k=2}^{K}\frac{\mathbb{E}\left[\varepsilon_{i}^{k}|X_{i}^{*},S_{i}\right]}{k!}{g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)}+O(\tau_{n}^{K+1}).$ (21)
-Suppose $\mathbb{E}\left[\varepsilon_{i}^{k}|X_{i}^{*},S_{i}\right]=v_{k}(X_{i}^{*},S_{i},\omega_{0})$, $k\in\{2,\ldots,K\}$ where $\omega_{0}\in\mathbb{R}^{\dim(\omega)}$ are unknown parameters.
-Typically, functions $v_{k}$ are assumed to depend only on some of the components of $(X_{i}^{*},S_{i})$.
-For example, in the nonlinear regression (2), it may be natural for $v_{k}$ to depend on $X_{i}^{*}$ and/or $W_{i}$, but not on $Y_{i}$.
-Equation (21) then implies that
-$\mathbb{E}\left[g(X_{i}^{*},S_{i},\theta)\right]=\mathbb{E}\left[g(X_{i},S_{i},\theta)-\sum_{k=2}^{K}\frac{v_{k}(X_{i}^{*},S_{i},\omega_{0})}{k!}{g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)}\right]+O(\tau_{n}^{K+1}).$ (22)
-Equation (22) is an analog of equation (10), which motivates correcting the moment conditions by estimating and subtracting the bias terms $\frac{v_{k}(X_{i}^{*},S_{i},\omega_{0})}{k!}{g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)}$.
-The corrected moment function takes the form
-$\psi(X_{i},S_{i},\theta,\omega)=g(X_{i},S_{i},\theta)-\sum_{k=2}^{K}f_{k}(X_{i},S_{i},\theta,\omega),$ (23)
-where
-$f_{2}(x,s,\theta,\omega)=$ $\frac{v_{2}(x,s,\omega)}{2}g_{x}^{(2)}(x,s,\theta),\quad f_{3}(x,s,\theta,\omega)=\frac{v_{3}(x,s,\omega)}{6}g_{x}^{(3)}(x,s,\theta),$
-$f_{k}(x,s,\theta,\omega)=$ $\frac{v_{k}(x,s,\omega)}{k!}g_{x}^{(k)}(x,s,\theta)-\sum_{\ell=2}^{k-2}\frac{v_{k-\ell}(x,s,\omega)}{(k-\ell)!}\frac{\partial^{k-\ell}}{\partial x^{k-\ell}}f_{\ell}(x,s,\theta,\omega),\text{ for $k\geq 4$}.$
-Notice that for $K\leq 3$ the corrected moment functions are similar to the classical EIV case except for $v_{2}(x,s,\omega)$ and $v_{3}(x,s,\omega)$ taking the places of $\mathbb{E}\left[\varepsilon_{i}^{2}\right]$ and $\mathbb{E}\left[\varepsilon_{i}^{3}\right]$ respectively.
-For $K\geq 4$, we need to bias correct the bias correction terms such as $f_{2}(x,s,\theta,\omega)$ accounting for $v_{2}(X_{i},S_{i},\theta,\omega)$ being evaluated at $X_{i}$ instead of $X_{i}^{*}$.
-Also, when the conditional moments of $\varepsilon_{i}$ do not depend on $X_{i}^{*}$, the corrected moment function matches the classical EIV case with $v_{k}(s,\omega)$ replacing $\mathbb{E}\left[\varepsilon_{i}^{k}\right]$.
-Under the regularity conditions
-$\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\omega_{0})\right]=\mathbb{E}\left[g(X_{i}^{*},S_{i},\theta_{0})\right]+O(\tau_{n}^{K+1})=o(n^{-1/2}),$
-provided that $O(\tau_{n}^{K+1})=o(n^{-1/2})$.
-Thus, the model can be estimated using GMM estimator with the corrected moment function (23), where parameters $\theta_{0}$ and $\omega_{0}$ are estimated together, i.e., the estimator in equation (13) with $\beta\equiv(\theta^{\prime},\omega^{\prime})^{\prime}$.
-The following example provides a convenient parameterization of functions $v$.
-Example 4.1 (Exponential Specification, $K=4$).
-Suppose $\varepsilon_{i}=\exp(\omega_{01}X_{i}^{*})\zeta_{i}$, where $\zeta_{i}$ is mean zero and independent of $(X_{i}^{*},S_{i})$.
-In this case, $\mathbb{E}\left[\varepsilon_{i}^{k}|X_{i}^{*},S\right]=\mathbb{E}\left[\zeta_{i}^{k}\right]\exp(k\omega_{01}X_{i}^{*})$.
-Then, we can take $v_{k}(x,s,\omega)=\omega_{k}\exp(k\omega_{1}x)$, $k\in\{2,\ldots,4\}$, with $\omega_{0}=(\omega_{01},\mathbb{E}\left[\zeta_{i}^{2}\right],\mathbb{E}\left[\zeta_{i}^{3}\right],\mathbb{E}\left[\zeta_{i}^{4}\right])^{\prime}$.
-Functions $f_{k}$ in equation (23) are
-$f_{2}(x,s,\theta,\omega)$ $=\frac{\omega_{2}}{2}e^{2\omega_{1}x}g_{x}^{(2)}(x,s,\theta),\quad f_{3}(x,s,\theta,\omega)=\frac{\omega_{3}}{6}e^{3\omega_{1}x}g_{x}^{(3)}(x,s,\theta),$
-$f_{4}(x,s,\theta,\omega)$ $=\textstyle e^{4\omega_{1}x}\left(\frac{\omega_{4}-6\omega_{2}^{2}}{24}g_{x}^{(4)}(x,s,\theta)-\omega_{1}\omega_{2}^{2}g_{x}^{(3)}(x,s,\theta)-\omega_{1}^{2}\omega_{2}^{2}g_{x}^{(2)}(x,s,\theta)\right).\ \ \hfill\blacksquare$
-Identifiability of parameters $\omega_{0}$ depends on the specific model and moment conditions $g$.
-For the nonlinear regression model with weakly classical EIV Evdokimov and Zeleneev (2022) establish nonparametric identification using instrumental variables.
-4.2.2 Non-Classical Measurement Errors
-Since $X_{i}^{\ast}$ is not observed, in the absence of some a priori information about its distribution, some assumptions such as $\mathbb{E}\left[\left.X_{i}\right|X_{i}^{\ast},S_{i}\right]=X_{i}^{\ast}$ appear to be necessary, and are routinely imposed in the EIV literature.
-For example, see Hu and Schennach (2008), Hu and Sasaki (2015), Freyberger (2021), and Schennach (2021) for recent discussions of such assumptions.
-In the absence of such assumptions one cannot determine even the scale of the true unobserved covariate.
-To be specific, suppose the true covariate is denoted by $\mathcal{X}_{i}^{\ast}$ and is mismeasured according to the following model
-$X_{i}=\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast}+\mathcal{E}_{i},\qquad\mathbb{E}\left[\left.\mathcal{E}_{i}\right|\mathcal{X}_{i}^{\ast},S_{i}\right]=0,$ (24)
-so $\mathbb{E}\left[\left.X_{i}\right|\mathcal{X}_{i}^{\ast}\right]=\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast}$ for some unknown $\alpha_{0}$ and $\alpha_{1}>0$. Note that
-$\mathbb{E}\left[\left.X_{i}-\mathcal{X}_{i}^{\ast}\right|\mathcal{X}_{i}^{\ast}\right]\neq 0,$
-and the measurement error is non-classical. For concreteness, consider the nonlinear regression model:
-$\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast},W_{i}\right]=\rho\left(\theta_{01}+\theta_{0\mathcal{X}}\mathcal{X}_{i}^{\ast}+\theta_{0W}^{\prime}W_{i}\right)$ (25)
-for some known function $\rho$.
-Since $\mathcal{X}_{i}^{\ast}$ is unobservable, we can write an observationally equivalent model in terms of another unobservable $X_{i}^{\ast}$ defined as
-$X_{i}^{\ast}$ $\equiv\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast},$ (26)
-$\mathbb{E}\left[Y_{i}|X_{i}^{\ast},W_{i}\right]$ $=\rho\left(\tilde{\theta}_{01}+\tilde{\theta}_{0X}{\large X}_{i}^{\ast}+\theta_{0W}^{\prime}W_{i}\right),$ (27)
-where $\tilde{\theta}_{0X}\equiv\theta_{0\mathcal{X}}/\alpha_{1},\quad\tilde{\theta}_{01}\equiv\theta_{01}-\tilde{\theta}_{0X}\alpha_{0}$.
-Thus, parameters $\left(\theta_{01},\theta_{0\mathcal{X}}\right)$ and $\left(\alpha_{0},\alpha_{1}\right)$ are not separately identified without additional information about the location and scale of $\mathcal{X}_{i}^{\ast}$.
-Note that neither instrumental variables nor multiple measurements of the form (24) provide such information.
-Equation (24) and the definition (26) of $X_{i}^{\ast}$ imply that condition $\mathbb{E}\left[\left.X_{i}\right|X_{i}^{\ast},S_{i}\right]=X_{i}^{\ast}$ holds for the observationally equivalent model (27).
-This observationally equivalent model can be identified and estimated using the approach of Section 4.2.1.
-This then allows identification of many parameters of interest in the original model (25) because they do not depend on the (unidentified) scale of $\mathcal{X}_{i}^{\ast}$.
-First, parameters $\theta_{0W}$ are the same in both models.
-Moreover, we have $\left(\mathcal{X}_{i}^{\ast}-\mu_{\mathcal{X}^{\ast}}\right)/\sigma_{\mathcal{X}^{\ast}}=\left(X_{i}^{\ast}-\mu_{X^{\ast}}\right)/\sigma_{X^{\ast}}$, and hence the following average structural function
-$m(t,w)\equiv\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast}=\mu_{\mathcal{X}^{\ast}}+t\sigma_{\mathcal{X}^{\ast}},W_{i}=w\right]$
-is identified, since $m(t,w)=\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=\mu_{X^{\ast}}+t\sigma_{X^{\ast}},W_{i}=w\right]$ is identified.
-As a result, we can identify marginal effects such as $\frac{\partial}{\partial w}\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast}=\mu_{\mathcal{X}^{\ast}},W_{i}=w\right]=\frac{\partial}{\partial w}m(0,w)$, or the effect of increasing $\mathcal{X}_{i}^{\ast}$ from $\mu_{\mathcal{X}^{\ast}}$ to $\mu_{\mathcal{X}^{\ast}}+\lambda\sigma_{\mathcal{X}^{\ast}}$, i.e., $m(\lambda,w)-m(0,w)$.
-The corresponding marginal effects averaged with respect to the distribution of $\left(\mathcal{X}_{i}^{\ast},W_{i}\right)$ are also identified and can be estimated following Remark 3.
-Thus, in this model, the MERM estimator from Section 4.2.1 allows estimation of many economically meaningful marginal effects, even though the true measurement model is given by equation (24), and the measurement error $X_{i}-\mathcal{X}_{i}^{\ast}$ is non-classical.
-[^10] In the constructed observationally equivalent model (27) we have $\varepsilon_{i}=\mathcal{E}_{i}$, and the MERM approach is justified when $\sigma_{\mathcal{E}}^{2}/\sigma_{X}^{2}$ is moderate.
-The true measurement error $X_{i}-\mathcal{X}_{i}^{\ast}$ does not need to be small.
-At the same time, identification of the true values of $\theta_{01}$ and $\theta_{0\mathcal{X}}$ or some marginal effects such as $\frac{\partial}{\partial\varkappa}\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast}=\varkappa,W_{i}=w\right]$ requires identifying $\alpha_{0}$ and $\alpha_{1}$.
-Using the summary statistics for $\mathcal{X}^{\ast}$
-In some applications the researchers may have information about the marginal distribution of $\mathcal{X}^{\ast}$ such as its mean $\mu_{\mathcal{X}^{\ast}}$ and variance $\sigma_{\mathcal{X}^{\ast}}^{2}$, e.g., from an administrative dataset.
-[^11] It is important that $\mathcal{X}_{j}^{\ast}$ and $\left(X_{i},S_{i}\right)$ are drawn from the same population., [^12] Knowing $\sigma_{\mathcal{X}^{\ast}}^{2}$ does not identify even the unconditional variance $\sigma_{\varepsilon}^{2}$, since $\sigma_{X}^{2}=\alpha_{1}^{2}\sigma_{\mathcal{X}^{\ast}}^{2}+\sigma_{\varepsilon}^{2}$.
-Then, the MERM approach can be used to estimate $\alpha_{0}$ and $\alpha_{1}$ together with parameters $\theta_{0}$.
-For any original moment condition $\mathbb{E}\left[g\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)\right]=0$ consider the augmented moment condition $\mathbb{E}\left[g_{A}\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)\right]=0$, where
-$g_{A}\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)\equiv\left(g\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)^{\prime},\mathcal{X}_{i}^{\ast}-\mu_{\mathcal{X}^{\ast}},\left(\mathcal{X}_{i}^{\ast}-\mu_{\mathcal{X}^{\ast}}\right)^{2}-\sigma_{\mathcal{X}^{\ast}}^{2}\right)^{\prime}.$
-Define $X_{i}^{\ast}\equiv\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast}$, $\tilde{\theta}_{0}\equiv\left(\theta_{0}^{\prime},\alpha_{0},\alpha_{1}\right)^{\prime}$, and $\tilde{g}_{A}(X_{i}^{\ast},S_{i},\tilde{\theta}_{0})\equiv g_{A}\left(\left.\left(X_{i}^{\ast}-\alpha_{0}\right)\right/\alpha_{1},S_{i},\theta_{0}\right)$.
-Then one can take $\mathbb{E}[\tilde{g}_{A}(X_{i}^{\ast},S_{i},\tilde{\theta}_{0})]=0$ as the new “original” moment condition (1).
-Since the condition $\mathbb{E}\left[X_{i}|X_{i}^{\ast},S_{i}\right]=X_{i}^{\ast}$ holds, $\tilde{\theta}_{0}$ can be estimated using the MERM approach and the corrected moments (23) applied to $\tilde{g}_{A}$ in place of $g$.
-Evdokimov and Zeleneev (2022) show how an identification argument based on the assumption of weakly classical measurement errors can be used for nonparametric identification in the settings with non-classical measurement errors.
-The linearity of equation (24) simplifies the estimation procedure, but the underlying approach is valid nonparametrically.
-Using a parsimonious model, this section provides a practical way of dealing with non-classical measurement errors in empirical applications, avoiding any nonparametric estimation.
 Appendix A Regularity Conditions
@@
   4. (iv)
-[…]
+    $\hat{\Xi}\overset{p}{\rightarrow}\Xi$, where $\Xi$ is a symmetric positive definite matrix;
   5. (v)
@@
 $B_{n}=\begin{bmatrix}1&0&\ldots&0&0\\ \mathbb{E}[\varepsilon_{i}/\sigma_{n}]&1&\ldots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \frac{\mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K-3}]}{(K-3)!}&\frac{\mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K-4}]}{(K-4)!}&\ldots&1&0\\ \frac{\mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K-2}]}{(K-2)!}&\frac{\mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K-3}]}{(K-3)!}&\ldots&\mathbb{E}[(\varepsilon_{i}/\sigma_{n})]&1\\ \end{bmatrix},\quad c_{n}=\begin{bmatrix}\mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{2}]/2!\\ \mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{3}]/3!\\ \vdots\\ \mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K-1}]/(K-1)!\\ \mathbb{E}[(\varepsilon_{i}/\sigma_{n})^{K}]/K!\end{bmatrix}$.
-[…]
+Since $B_{n}$ is invertible, (B.6) has a unique solution $a_{n}=B_{n}^{-1}c_{n}$.
+Moreover, $a_{n}$ is bounded since both $B_{n}^{-1}$ and $c_{n}$ are bounded (Assumption 2.1).
+Hence, we conclude that (B.5) has a unique solution $\gamma_{0n}=\left(\sigma_{n}^{2}a_{2n},\dots,\sigma_{n}^{K}a_{Kn}\right)^{\prime}$.
+Since (B.5) is satisfied, using Assumption 2.2, we also conclude that
 $\mathbb{E}\left[\sum_{k=2}^{K}g_{x}^{(k)}(X_{i}^{*},S_{i},\theta_{0})\left(\dfrac{1}{k!}\varepsilon_{i}^{k}-\sum_{\ell=2}^{k}\dfrac{1}{(k-\ell)!}\varepsilon_{i}^{k-\ell}\gamma_{0\ell n}\right)\right]=0.$ (B.7)
@@
 $\leqslant a_{kn}\left(\sigma_{n}^{K+1}{\mathbb{E}\left[b_{1}(X_{i}^{*},S_{i},\theta_{0})\right]\mathbb{E}\left[\left|\varepsilon_{i}/\sigma_{n}\right|^{K+1-k}\right]}+\sigma_{n}^{M}{\mathbb{E}\left[b_{2}(X_{i}^{*},S_{i},\theta_{0})\right]\mathbb{E}\left[\left|\varepsilon_{i}/\sigma_{n}\right|^{M-k}\right]}\right).$
-[…]
+Since $a_{kn}$ is bounded, we conclude that (B.9) holds analogously to (B.8).
 Combining (B.3) with (B.4), and (B.7)-(B.9), we conclude that $\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right]=o(n^{-1/2})$.
@@
 Similarly, we let $a_{i}^{*}(\alpha)\equiv a(X_{i}^{*},S_{i},\alpha)$, $\overline{a}^{*}(\alpha)\equiv n^{-1}\sum_{i=1}^{n}a_{i}^{*}(\alpha)$, $a^{*}(\alpha)\equiv\mathbb{E}[a_{i}^{*}(\alpha)]$.
-For the true value of the parameter $\alpha_{0}$, we often write $a_{i}\equiv a(\alpha_{0})$, $\overline{a}\equiv\overline{a}(\alpha_{0})$, $a\equiv a(\alpha_{0})$, $a_{i}^{*}\equiv a(\alpha_{0})$, $\overline{a}^{*}\equiv\overline{a}^{*}(\alpha_{0})$, $a^{*}\equiv a^{*}(\alpha_{0})$.
+[…]
 C.1 Auxiliary lemmas
@@
 Hence, $\sup_{\theta\in\Theta}\left\|\overline{g}(\theta)-\overline{g}^{*}(\theta)\right\|=o_{p}(1)$, and we are left to show $\sup_{\theta\in\Theta}\left\|\overline{g}^{*}(\theta)-g^{*}(\theta)\right\|=o_{p}(1)$.
-This, in turn, follows from the standard ULLN (e.g., Lemma 2.4 in Newey and McFadden, 1994), which also ensures continuity of $g^{*}(\theta)$ on $\Theta$.
+This, in turn, follows from the standard ULLN (e.g., Lemma 2.4 in NeweyMcFadden1994HB), which also ensures continuity of $g^{*}(\theta)$ on $\Theta$.
 Hence, we conclude that the assertion of the lemma holds for $g$.
@@
 Lemma C.2.
-[…]
+Suppose that the hypotheses of Lemma C.1 are satisfied.
+Then, $g_{x}^{(k)}\rightarrow g_{x}^{(k)*}$ and $G_{x}^{(k)}\rightarrow G_{x}^{(k)*}$ for $k\in\{0,\dots,K\}$.
+Suppose also $\hat{\theta}\overset{p}{\rightarrow}\theta_{0}$.
+Then, $\overline{g}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}g_{x}^{(k)*}$ and $\overline{G}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}G_{x}^{(k)*}$ for $k\in\{0,\dots,K\}$.
 Proof of Lemma C.2.
@@
 where the convergence follows from Assumptions 2.1, A.1(ii) and A.3(ii). Hence, we conclude $g_{x}^{(k)}\rightarrow g_{x}^{(k)*}$.
-[…]
+Next, we show $\overline{g}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}g_{x}^{(k)*}$. By the triangle inequality,
 $\left\|\overline{g}_{x}^{(k)}(\hat{\theta})-g_{x}^{(k)*}\right\|\leqslant\sup_{\theta\in B_{\delta}(\theta_{0})}\left\|\overline{g}_{x}^{(k)}(\theta)-g_{x}^{(k)*}(\theta)\right\|+\left\|g_{x}^{(k)*}(\hat{\theta})-g_{x}^{(k)*}(\theta_{0})\right\|,$
-[…]
+where the inequality holds with probability approaching one since $\hat{\theta}\in B_{\delta}(\theta_{0})$ with probability approaching one.
+Note that, By Lemma C.1, $\sup_{\theta\in B_{\delta}(\theta_{0})}\left\|\overline{g}_{x}^{(k)}(\theta)-g_{x}^{(k)*}(\theta)\right\|=o_{p}(1)$ and $\left\|g_{x}^{(k)*}(\hat{\theta})-g_{x}^{(k)*}(\theta_{0})\right\|=o_{p}(1)$, where the second result follows from consistency of $\hat{\theta}$ and continuity of $g_{x}^{(k)*}(\theta)$.
+Hence, $\overline{g}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}g_{x}^{(k)*}$, which completes the proof of the results for $g_{x}^{(k)}$ for all $k\in\{0,\dots,K\}$.
 A nearly identical argument, can be invoked to establish the same results for $G_{x}^{(k)}$ for $k\in\{0,\dots,K\}$, with Assumptions A.3(i) and (iii) taking the places of Assumptions A.1(i) and A.3(ii), respectively.
 Lemma C.3.
-[…]
+Suppose that the hypotheses of Lemma C.1 are satisfied.
+Then, under additional Assumptions A.3(iv) and 2.3(ii), we have $\hat{\theta}\overset{p}{\rightarrow}\theta_{0}$, $\hat{\gamma}\overset{p}{\rightarrow}0$ and $\hat{\gamma}\overset{p}{\rightarrow}\gamma_{0n}$.
 Proof of Lemma C.3.
@@
 Moreover, Lemma C.1(i) also ensures that $\psi^{*}(\beta)$ is continuous on compact $\mathcal{B}$ and, consequently, is bounded.
-[…]
+Let $\hat{Q}(\beta)=\overline{\psi}(\beta)^{\prime}\hat{\Xi}\overline{\psi}(\beta)$ and $Q^{*}(\beta)=\psi^{*}(\beta)^{\prime}\Xi\psi^{*}(\beta)$.
+Notice that (C.1), boundedness of $\psi^{*}(\beta)$, and Assumption A.3(iv) together guarantee that $\sup_{\beta\in\mathcal{B}}\left|\hat{Q}(\beta)-Q^{*}(\beta)\right|=o_{p}(1)$.
+Next, recall that $\gamma_{0n}\rightarrow 0_{K-1}$ (Lemma 1).
+Since $\Gamma$ is compact and $\gamma_{0n}\in\Gamma$ (Assumption A.2(i)), $0_{K-1}\in\Gamma$.
+Consequently, Assumptions 2.3(ii) and A.3(iv) together guarantee that $Q^{*}(\beta)$ is uniquely minimized at $\theta=\theta_{0}$ and $\gamma=0_{K-1}$.
+Consequently, applying the standard consistency argument (e.g., Theorem 2.1 of NeweyMcFadden1994HB), we conclude that $\hat{\theta}\rightarrow\theta_{0}$ and $\hat{\gamma}\rightarrow 0_{K-1}$.
+Finally, since $\gamma_{0n}\rightarrow 0$ (Lemma 1), we also have $\hat{\gamma}\overset{p}{\rightarrow}\gamma_{0n}$.
 Lemma C.4.
 Suppose that $\{(X_{i}^{*},S_{i}^{\prime},\varepsilon_{i})\}_{i=1}^{n}$ are i.i.d.. Then, under Assumptions 2.1, 2.2, A.1, and A.3(ii) and (v), we have
-[…]
+$n^{1/2}\overline{\psi}(\beta_{0n})\overset{d}{\rightarrow}N(0,\Omega_{gg}^{*}),$
 where $\Omega_{gg}^{*}\equiv\mathbb{E}\left[g(X_{i},S_{i},\theta_{0})g(X_{i},S_{i},\theta_{0})^{\prime}\right]$.
@@
 $-\sum_{k=2}^{K}\dfrac{\gamma_{0kn}}{(K-k)!}n^{-1/2}\sum_{i=1}^{n}\left(g_{x}^{(K)}(\tilde{X}_{ki},S_{i},\theta_{0})-g_{x}^{(K)}(X_{i}^{*},S_{i},\theta_{0})\right)\varepsilon_{i}^{K-k}.$ (C.2)
-[…]
+First, note that, by the standard CLT, $n^{-1/2}\sum_{i=1}^{n}g(X_{i}^{*},S_{i},\theta_{0})\overset{d}{\rightarrow}N(0,\Omega_{gg}^{*})$.
+The rest of the proof is to show that the remaining terms are $o_{p}(1)$.
+By Assumptions 2.1, 2.2, A.3(ii), Chebyshev’s inequality guarantees
 $n^{-1/2}\sum_{i=1}^{n}g_{x}^{(1)}(X_{i}^{*},S_{i},\theta_{0n})\varepsilon_{i}=o_{p}(1)$
@@
 $n^{1/2}\overline{\psi}(\beta_{0n})=$ $n^{-1/2}\sum_{i=1}^{n}g(X_{i}^{*},S_{i},\theta_{0})+o_{p}(1)$
-[…]
+$\overset{d}{\rightarrow}N(0,\Omega_{gg}^{*}),$
 which completes the proof.
@@
 Equipped with Lemmas C.1-C.4, we are ready to prove Theorem 2.
-Proof of Theorem 2.
 Since (i) $\hat{\theta}$ and $\hat{\gamma}$ are consistent for $\theta_{0}$ and $\gamma_{0n}$, respectively (Lemma C.3) and (ii) both $\theta_{0}$ and $\gamma_{0n}\rightarrow 0$ (Assumption 2.1) are bounded away from the boundaries of $\Theta$ and $\Gamma$ respectively (Assumption A.2), the standard GMM FOC is satisfied with probability approaching one, i.e., we have (with probability approaching one)
 $\overline{\Psi}(\hat{\beta})^{\prime}\hat{\Xi}\overline{\psi}(\hat{\beta})=0.$
-Expanding $\overline{\psi}(\hat{\beta})$ around $\overline{\psi}(\beta_{0n})$ gives
+[…]
 $\overline{\Psi}(\hat{\beta})^{\prime}\hat{\Xi}\left(\overline{\psi}(\beta_{0n})+\overline{\Psi}(\tilde{\beta})(\hat{\beta}-\beta_{0})\right)=0,$ (C.3)
+where $\tilde{\beta}$ lies between $\beta_{0n}$ and $\hat{\beta}$ (and, consequently, $\tilde{\theta}\overset{p}{\rightarrow}\theta_{0}$ and $\tilde{\gamma}\overset{p}{\rightarrow}0$).
+Next, we argue that $\overline{\Psi}(\hat{\beta})=\Psi^{*}+o_{p}(1)$.
+Observe
 […]
-$\overline{\Psi}(\hat{\beta})=\left[\overline{G}(\hat{\theta})-\sum_{k=2}^{K}\hat{\gamma}_{k}\overline{G}_{x}^{(k)}(\hat{\theta}),-\overline{g}_{x}^{(2)}(\hat{\theta}),\dots,-\overline{g}_{x}^{(K)}(\hat{\theta})\right].$
-[…]
+Since $\hat{\theta}\overset{p}{\rightarrow}\theta_{0}$ (Lemma C.3), we can invoke the result of Lemma C.2 to argue that $\overline{g}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}g_{x}^{(k)*}$ and $\overline{G}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}G_{x}^{(k)*}$ for all $k\in\{0,\dots,K\}$.
+This, combined with $\hat{\gamma}\rightarrow 0$ (Lemma C.3), ensures that $\overline{\Psi}(\hat{\beta})=\Psi^{*}+o_{p}(1)$ and, analogously, $\overline{\Psi}(\tilde{\beta})=\Psi^{*}+o_{p}(1)$.
+Coupling these result with Assumption A.3(iv), we conclude that $\overline{\Psi}(\hat{\beta})^{\prime}\hat{\Xi}\overline{\Psi}(\tilde{\beta})\overset{p}{\rightarrow}\Psi^{*\prime}\Xi\Psi^{*}$, which is invertible by Assumption 2.3(i).
+Hence, (C.3) can be rearranged as (with probability approaching one)
 $n^{1/2}(\hat{\beta}-\beta_{0n})$ $=-\left(\overline{\Psi}(\hat{\beta})^{\prime}\hat{\Xi}\overline{\Psi}(\tilde{\beta})\right)^{-1}\overline{\Psi}(\hat{\beta})^{\prime}\hat{\Xi}n^{1/2}\overline{\psi}(\beta_{0n})$
 $=-\left(\Psi^{*\prime}\Xi\Psi^{*}\right)^{-1}\Psi^{*\prime}\Xi n^{1/2}\overline{\psi}(\beta_{0n})+o_{p}(1),$
-[…]
-[…]
+where, by Lemma C.4, $n^{1/2}\overline{\psi}(\beta_{0n})\overset{d}{\rightarrow}N(0,\Omega_{gg}^{*})$. Hence, we conclude
+$n^{1/2}(\hat{\beta}-\beta_{0n})\overset{d}{\rightarrow}N(0,\Sigma^{*}),$
 where
@@
 which completes the proof.
-Appendix D Details on the sufficient condition (17)
+Appendix D Proof of Lemma 3
+As in the previous proofs, we use notation $\sigma_{n}^{2}\equiv\mathbb{E}[\varepsilon_{i}^{2}]\equiv\sigma_{\varepsilon}^{2}$, $\hat{\sigma}^{2}_{n}\equiv\hat{\sigma}_{\varepsilon}^{2}$, and $\hat{\sigma}_{n}^{K}\equiv\hat{\sigma}_{\varepsilon}^{K}$.
+First, we show that (19) holds. By Theorem 2, we have $\hat{\sigma}_{n}^{2}=\sigma_{n}^{2}+O_{p}(n^{-1/2})$. Since $K$ is even, (19) can be obtained by expanding
+$\hat{\sigma}_{n}^{K}$ $=(\sigma_{n}^{2}+O_{p}(n^{-1/2}))^{K/2}=\sum_{\ell=0}^{K/2}\binom{n}{\ell}(\sigma_{n}^{2})^{K/2-\ell}(O_{p}(n^{-1/2}))^{\ell}$
+$=\sigma_{n}^{K}+\sum_{\ell=1}^{K/2}\binom{n}{\ell}(\sigma_{n}^{2})^{K/2-\ell}O_{p}(n^{-\ell/2})$
+$=\sigma_{n}^{K}+O_{p}(\max\{\sigma_{n}^{K-2}n^{-1/2},n^{-K/4}\}).$
+Next, we prove the first statement of the lemma.
+First, using Lemma C.2, we have $\overline{g}_{x}^{(K)}(\hat{\theta})=g_{x}^{(K)*}+o_{p}(1)$.
+Here, $\|g_{x}^{(K)*}\|$ is bounded away from zero and above, with the former implied by Assumption 2.3(i).
+Likewise, since the Jacobian corresponding to $\hat{\beta}_{L}$ (involved in the construction of $B$) is a submatrix of the Jacobian corresponding to $\hat{\beta}_{K}$, we conclude that $\|Bg_{x}^{(K)*}\|$ is also bounded from zero and above.
+Hence, using consistency of $\hat{B}$ and $\hat{\Sigma}$, we conclude that, with probability approaching one, for some $\underline{C},\overline{C}>0$, we have
+$\underline{C}\sqrt{n}\hat{\sigma}_{n}^{K}\leq\max_{1\leq\ell\leq\dim(\beta)}|\hat{\Sigma}_{\ell\ell}^{-1/2}\sqrt{n}\hat{\sigma}^{K}\hat{B}_{\ell\cdot}\overline{g}_{x}^{(K)}(\hat{\theta})|\leq\overline{C}\sqrt{n}\hat{\sigma}_{n}^{K}.$ (D.1)
+To prove the first part of the lemma it is sufficient to show that $\delta_{n}\sigma_{n}^{L+1}=o_{p}(n^{-1/2})$.
+Let $n_{j}$ index a subsequence of $\delta_{n_{j}}$ such that $\delta_{n_{j}}=1$.
+If this subsequence is finite, the statement is trivial.
+Otherwise, we need to show that $\sigma_{n_{j}}^{L+1}=o(n_{j}^{-1/2})$ along this subsequence.
+We prove this by contradiction.
+Suppose $\sigma_{n_{j}}^{L+1}=o(n_{j}^{-1/2})$ does not hold meaning that there exists another subsequence within $n_{j}$, denoted by $n_{m}$ such that for all sufficiently large $m$ we have $\sqrt{n_{m}}\sigma_{n_{m}}^{L+1}\geq C>0$ and $\delta_{n_{m}}=1$.
+Since we have $\sigma_{n_{m}}\geq Cn_{m}^{-\frac{1}{2(L+1)}}$ and $\sqrt{n_{m}}\sigma_{n_{m}}^{K}\geq Cn_{m}^{-(K-L-1)/2(L+1)}$, using (19), we conclude that, for some $C>0$, with probability approaching one, we have
+$\sqrt{n_{m}}\hat{\sigma}_{n_{m}}^{K}=\sqrt{n_{m}}\sigma_{n_{m}}^{K}+O_{p}\left(\sigma_{n_{m}}^{K-2}+n_{n_{m}}^{-\left(K-2\right)/4}\right)\geq Cn_{m}^{-(K-L-1)/2(L+1)}$ (D.2)
+because the remainder $O_{p}(\sigma_{n_{m}}^{K-2}+n_{n_{m}}^{-\left(K-2\right)/4})$ is dominated by $\sqrt{n_{m}}\sigma_{n_{m}}^{K}$ in this case.
+At the same time, $\delta_{n_{m}}=1$ together with (D.1) implies that $\sqrt{n_{m}}\hat{\sigma}_{n_{m}}^{K}\leq C\varkappa_{n_{m}}$ for some $C>0$.
+This contradicts (D.2) since $\varkappa_{n}n^{\left(K-L-1\right)/\left(2L+2\right)}\rightarrow 0$.
+The contradiction completes the proof of the first part.
+We now turn to the second part of the lemma. First, note that, using Jensen’s inequality, we obtain
+$\sqrt{n}\hat{\sigma}_{n}^{K}=\sqrt{n}(\sigma_{n}^{2}+O_{p}(n^{-1/2}))^{K/2}\leq C\sqrt{n}\sigma_{n}^{K}+O_{p}(n^{-(K-2)/4}).$
+In particular, for $\varkappa_{n}=n^{-\left(K-L-1\right)/\left(2L+2\right)}\left(\ln n\right)^{-a}$ and $\sigma_{n}=o(n^{-\frac{1}{2L+2}-\epsilon})$, the above implies that $\sqrt{n}\hat{\sigma}_{n}^{K}/\varkappa_{n}=o_{p}(1)$.
+This, together with (D.1), implies that $\delta_{n}=1$ with probability approaching one, which completes the proof.
+Appendix E Details on the sufficient condition (17)
 In this section, we demonstrate that the Jacobian $\Psi^{\ast}$ associated with the moment conditions (16) in Section 2.4 is guaranteed to have a full rank (for $K=2$) when the sufficient condition (17) holds.
@@
 where we used $\mathbb{E}\left[\varepsilon_{Q,i}|X_{i}^{\ast},S_{i},\varepsilon_{i}\right]=0$.
-Dividing moments $J+2,\ldots,2J+1$ by $\alpha_{1}$ and subtracting them from the moments $12,\ldots,J+$ 1, we obtain
+Dividing moments $J+2,\ldots,2J+1$ by $\alpha_{1}$ and subtracting them from the moments $2,\ldots,J+1$, we obtain
 $\mathop{\mathrm{R}k}\left(\Psi^{\ast}\right)$ $=$ $\mathop{\mathrm{R}k}\mathbb{E}\begin{pmatrix}u_{\theta,i}\left(X_{i}^{\ast},\theta\right)&-u_{x,i}^{(2)}\left(X_{i}^{\ast},\theta\right)\\ 0&-2\varrho_{x,i}^{(1)}\left(X_{i}^{\ast},\theta_{0}\right)\\ \varrho_{\theta,i}\left(X_{i}^{\ast},\theta_{0}\right)X_{i}^{\ast}&-\varrho_{x,i}^{(2)}\left(X_{i}^{\ast},\theta_{0}\right)X_{i}^{\ast}\end{pmatrix}$
@@
 $\mathbb{E}\left[\varrho_{x,i}^{(1)}\left(X_{i}^{\ast},\theta_{0}\right)\right]=\mathbb{E}\left[u_{x}^{(1)}\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\left(1,X_{i}^{\ast},\ldots,\left(X_{i}^{\ast}\right)^{J-1}\right)^{\prime}\right]\neq 0,$
-where the last equality follows from (17). $\blacksquare$
-Appendix E Evaluating the remaining bias of the MERM estimator
-In this section, we provide additional details on the derivation of the (higher-order) bias of the MERM estimator of $v^{\prime}\hat{\beta}$ and discuss potential implications for inference.
-The derivation of the expression for $\text{Bias}(v^{\prime}\hat{\beta})$ provided in equation (18) is based on the standard expansion of the GMM first order conditions around $\beta_{0}$.
-Following the lines of the proof of Lemma 1, we can consider a higher order expansion of $\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})$ to the order $o(\tau^{\overline{K}})$, which implies
-$\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right]\approx\gamma_{0\overline{K}}\mathbb{E}\left[g_{x}^{(\overline{K})}(X_{i},S_{i},\theta_{0})\right],$
-where $\overline{K}=K+2$ if $K$ is even and $\varepsilon_{i}$ is symmetric, and $\overline{K}=K+1$ otherwise. Hence,
-$\text{Bias}(v^{\prime}\hat{\beta})\approx-v^{\prime}(\Psi^{\prime}\Xi\Psi)^{-1}\Psi^{\prime}\Xi\;\mathbb{E}\left[g_{x}^{(\overline{K})}(X_{i},S_{i},\theta_{0})\right]\gamma_{0\overline{K}}.$ (E.1)
-Notice that, except for $\gamma_{0\overline{K}}$, all the remaining elements on the right hand side of (E.1) are either known or can be consistently estimated.
-The unknown $\gamma_{0\overline{K}}$ is given by the recursive relationship (12), and the only unknown determinant of $\gamma_{0\overline{K}}$ is $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ because the lower order moments of $\varepsilon_{i}$ can be recovered from $\hat{\gamma}$ with a required precision.
-Hence, the ranges of the possible values of $\gamma_{0\overline{K}}$ and, subsequently, of $\text{Bias}(v^{\prime}\hat{\beta})$ can be determined based on an a priori range of possible values of $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ (or, equivalently, of $\mathbb{E}[(\varepsilon_{i}/\sigma_{\varepsilon})^{\overline{K}}]$ combined with the estimated $\sigma_{\varepsilon}^{2}$) specified by the researcher.
-Example (Evaluating the bias with $K=2$).
-Suppose the researcher uses the MERM estimator with $K=2$, and believes that $\mathbb{E}\left[\varepsilon_{i}^{3}\right]=0$.
-Then, the bound on the bias depends on the $\mathbb{E}\left[\varepsilon_{i}^{4}\right]$.
-The variance $\sigma_{\varepsilon}^{2}$ is estimated by the MERM estimator, hence it is convenient to bound the kurtosis of $\varepsilon_{i}$ as $\mathbb{E}\left[\varepsilon_{i}^{4}\right]/\sigma_{\varepsilon}^{4}\in[1,\overline{\kappa}]$ for some $\overline{\kappa}$.
-For example, one could take $\overline{\kappa}=10$ as a conservative bound for the kurtosis.
-[^13] For Gaussian $\varepsilon_{i}$ the kurtosis $\kappa$ is $3$.
-Student’s $t(\nu)$ distribution has $\kappa<10$ for all $\nu\geq 5$ (the kurtosis and the $4$th moment do not exist for $\nu\leq 4$).
-Since $\hat{\gamma}$ estimates $\gamma_{0}=\gamma_{02}=\sigma_{\varepsilon}^{2}/2$, the range of possible values of $\gamma_{04}$ can be approximately bounded using
-$\gamma_{04}=\frac{\mathbb{E}\left[\varepsilon_{i}^{4}\right]-6\sigma_{\varepsilon}^{4}}{24}\in\left[-\frac{5}{24}\sigma_{\varepsilon}^{4},\;\frac{\overline{\kappa}-6}{24}\sigma_{\varepsilon}^{4}\right]=\left[-\frac{5}{6}\gamma_{02}^{2},\;\frac{\overline{\kappa}-6}{6}\gamma_{02}^{2}\right].{}$
-Hence, the range of possible values of $\text{Bias}(v^{\prime}\hat{\beta})$ can be approximated as
-$\text{Bias}(v^{\prime}\hat{\beta})\in\begin{cases}\left[-5\overline{b}\hat{\gamma}^{2}/6,\;\overline{b}\hat{\gamma}^{2}(\overline{\kappa}-6)/6\right],&\text{if $\overline{b}\geq 0$},\\ \left[\overline{b}\hat{\gamma}^{2}(\overline{\kappa}-6)/6,\;-5\overline{b}\hat{\gamma}^{2}/6\right],&\text{otherwise},\end{cases}$
-where
-$\overline{b}\equiv-v^{\prime}(\hat{\Psi}^{\prime}\hat{\Xi}\hat{\Psi})^{-1}\hat{\Psi}^{\prime}\hat{\Xi}\;\left(\frac{1}{n}\sum_{i=1}^{n}g_{x}^{(4)}(X_{i},S_{i},\hat{\theta})\right).$
-Notice that if $\overline{\kappa}\leq 6$, we have $\gamma_{04}\leq 0$, and the sign of $\text{Bias}(v^{\prime}\hat{\beta})$ becomes known.
-If (the absolute value of the worst case) $\text{Bias}(v^{\prime}\hat{\beta})$ is sufficiently small relative to its standard error $\sqrt{v^{\prime}\hat{\Sigma}v/n}$, the estimator $v^{\prime}\hat{\beta}$ is approximately unbiased and the inference on $v^{\prime}\beta_{0}$ can be based on equation (14).
-Otherwise, the researcher could consider using a higher $K$ and/or possibly employing inference tools that take into account the remaining bias (e.g., ArmstrongKolesar2021QE).
-Appendix F MERM derivation when $\sigma_{\varepsilon}$ is not small
+where the last equality follows from (17).
+Appendix F Additional Numerical Results
+In this section, we provide additional numerical results for the experiments considered in Sections 3.2 and 4.1.
+F.1 Additional numerical results for Section 3.2
+In this section, we provide additional simulation results for the same experiment as considered in Section 3.2 but for larger values of $\tau\in\{1,1.5,2\}$.
+The results are provided in Table 6 below, reporting the same statistics as in Table LABEL:tab:_MC_mult_logit_2 in the main text.
+Table 6: Simulation results for the multinomial logit model
+MLE $K=2$ $K=4$
+bias, $10^{-2}$ std, $10^{-2}$ rmse, $10^{-2}$ size bias, $10^{-2}$ std, $10^{-2}$ rmse, $10^{-2}$ size bias, $10^{-2}$ std, $10^{-2}$ rmse, $10^{-2}$ size
+$\tau=1$
+$\partial p_{1}/\partial x$ -16.14 0.70 16.15 100.00 -11.45 2.21 11.66 99.16 0.21 3.53 3.54 5.44
+$\partial p_{1}/\partial w_{1}$ 11.82 1.40 11.91 100.00 8.49 2.46 8.84 96.14 0.48 2.88 2.92 6.30
+$\partial p_{1}/\partial w_{2}$ 2.08 0.67 2.19 89.28 1.63 0.84 1.84 56.78 0.10 1.04 1.05 5.94
+$\partial p_{2}/\partial x$ 8.78 0.65 8.80 100.00 6.59 1.50 6.76 97.64 -0.04 2.53 2.53 6.42
+$\partial p_{2}/\partial w_{1}$ -5.87 0.71 5.91 100.00 -4.24 1.23 4.42 96.08 -0.24 1.44 1.46 6.26
+$\partial p_{2}/\partial w_{2}$ -4.03 1.27 4.22 89.50 -3.20 1.62 3.59 57.34 -0.18 2.09 2.10 6.24
+$\partial p_{0}/\partial x$ 7.36 0.61 7.39 100.00 4.86 1.33 5.04 94.72 -0.17 1.95 1.96 5.30
+$\partial p_{0}/\partial w_{1}$ -5.95 0.73 6.00 100.00 -4.24 1.24 4.42 96.14 -0.24 1.44 1.46 6.28
+$\partial p_{0}/\partial w_{2}$ 1.94 0.61 2.04 89.52 1.57 0.79 1.76 57.96 0.08 1.05 1.05 6.24
+$\tau=3/2$
+$\partial p_{1}/\partial x$ -19.01 0.50 19.01 100.00 -16.59 1.32 16.64 100.00 -2.29 3.67 4.33 16.56
+$\partial p_{1}/\partial w_{1}$ 14.07 1.34 14.13 100.00 12.83 1.90 12.97 99.98 2.70 3.40 4.34 24.14
+$\partial p_{1}/\partial w_{2}$ 2.31 0.64 2.40 95.70 2.20 0.75 2.32 87.72 0.56 1.18 1.31 12.18
+$\partial p_{2}/\partial x$ 9.97 0.47 9.98 100.00 9.05 0.95 9.10 100.00 1.58 2.70 3.13 16.10
+$\partial p_{2}/\partial w_{1}$ -6.96 0.69 6.99 100.00 -6.40 0.96 6.48 99.98 -1.35 1.71 2.18 24.08
+$\partial p_{2}/\partial w_{2}$ -4.44 1.21 4.60 95.88 -4.29 1.43 4.52 87.88 -1.09 2.35 2.59 12.68
+$\partial p_{0}/\partial x$ 9.04 0.46 9.05 100.00 7.54 0.94 7.60 100.00 0.71 2.05 2.18 9.46
+$\partial p_{0}/\partial w_{1}$ -7.11 0.71 7.15 100.00 -6.42 0.98 6.50 99.98 -1.35 1.70 2.17 23.90
+$\partial p_{0}/\partial w_{2}$ 2.13 0.58 2.21 95.90 2.09 0.69 2.20 88.08 0.53 1.17 1.29 13.00
+$\tau=2$
+$\partial p_{1}/\partial x$ -20.28 0.38 20.29 100.00 -18.86 0.88 18.88 100.00 -5.98 3.91 7.15 55.08
+$\partial p_{1}/\partial w_{1}$ 15.09 1.32 15.15 100.00 14.91 1.66 15.00 100.00 5.87 3.87 7.03 55.08
+$\partial p_{1}/\partial w_{2}$ 2.39 0.63 2.47 97.42 2.37 0.71 2.47 94.38 1.12 1.30 1.72 26.44
+$\partial p_{2}/\partial x$ 10.45 0.36 10.46 100.00 9.97 0.69 9.99 100.00 3.80 2.82 4.73 47.94
+$\partial p_{2}/\partial w_{1}$ -7.45 0.68 7.48 100.00 -7.43 0.85 7.48 100.00 -2.95 1.95 3.54 54.96
+$\partial p_{2}/\partial w_{2}$ -4.58 1.18 4.73 97.42 -4.61 1.36 4.81 94.54 -2.20 2.54 3.36 27.46
+$\partial p_{0}/\partial x$ 9.83 0.36 9.84 100.00 8.88 0.70 8.91 100.00 2.18 2.21 3.11 28.50
+$\partial p_{0}/\partial w_{1}$ -7.64 0.71 7.68 100.00 -7.48 0.87 7.53 100.00 -2.92 1.93 3.50 54.80
+$\partial p_{0}/\partial w_{2}$ 2.19 0.57 2.26 97.42 2.24 0.66 2.34 94.62 1.08 1.26 1.65 27.78
+  • This table reports the simulated finite sample bias, standard deviation, RMSE, and size of the MLE and the MERM estimators and the corresponding t-tests for the partial derivatives $\partial p_{j}(x,w,\theta_{0})/\partial x$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{1}$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{2}$ for $j\in\{1,2,0\}$ evaluated at the population mean.
+The true values of the marginal effects are $(\partial p_{1}/\partial x,\partial p_{2}/\partial x,\partial p_{0}/\partial x)=(0.222,-0.111,-0.111)$ and zeros for the rest.
+The results are based on 5,000 replications.
+F.2 Additional numerical results for Section 4.1
+In this section, we provide additional numerical results for the experiment considered in Section 4.1.
+In particular, Tables 7 and 8 below report the same statistics as Table 5 in the main text, but for smaller and larger sample sizes $n=1000$ and $n=4000$.
+Table 7: Choice of $K$ simulation results for the multinomial logit model, $n=1000$
+MLE $K=2$ $K=4$ data-driven $K$
+bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$
+$\tau=1/4$
+$\partial p_{1}/\partial x$ -3.22 3.74 1.34 3.96 2.10 4.56 1.37 4.05
+$\partial p_{1}/\partial w_{1}$ 2.29 3.27 -0.13 3.32 -0.57 3.39 -0.15 3.34
+$\partial p_{1}/\partial w_{2}$ 0.52 1.19 -0.01 1.22 -0.10 1.26 -0.02 1.23
+$\partial p_{2}/\partial x$ 1.93 2.55 -0.73 2.78 -1.18 3.11 -0.75 2.82
+$\partial p_{2}/\partial w_{1}$ -1.15 1.64 0.07 1.66 0.29 1.70 0.07 1.67
+$\partial p_{2}/\partial w_{2}$ -1.02 2.36 0.03 2.46 0.21 2.53 0.04 2.46
+$\partial p_{0}/\partial x$ 1.28 1.93 -0.61 2.16 -0.92 2.41 -0.62 2.20
+$\partial p_{0}/\partial w_{1}$ -1.14 1.63 0.07 1.66 0.29 1.70 0.07 1.67
+$\partial p_{0}/\partial w_{2}$ 0.51 1.17 -0.02 1.24 -0.11 1.27 -0.02 1.24
+$\tau=1/2$
+$\partial p_{1}/\partial x$ -8.97 9.10 -1.82 4.10 1.76 4.57 1.35 4.82
+$\partial p_{1}/\partial w_{1}$ 6.42 6.79 1.88 4.04 -0.35 3.52 -0.10 3.73
+$\partial p_{1}/\partial w_{2}$ 1.32 1.67 0.42 1.33 -0.05 1.32 0.00 1.34
+$\partial p_{2}/\partial x$ 5.21 5.39 1.16 2.91 -0.97 3.20 -0.73 3.30
+$\partial p_{2}/\partial w_{1}$ -3.21 3.39 -0.94 2.02 0.18 1.76 0.05 1.87
+$\partial p_{2}/\partial w_{2}$ -2.59 3.26 -0.83 2.65 0.11 2.65 0.01 2.68
+$\partial p_{0}/\partial x$ 3.76 3.95 0.66 2.18 -0.79 2.50 -0.62 2.56
+$\partial p_{0}/\partial w_{1}$ -3.21 3.40 -0.94 2.02 0.18 1.76 0.05 1.86
+$\partial p_{0}/\partial w_{2}$ 1.27 1.60 0.41 1.33 -0.06 1.33 -0.01 1.35
+$\tau=3/4$
+$\partial p_{1}/\partial x$ -13.36 13.41 -7.17 7.97 1.07 4.74 1.07 4.75
+$\partial p_{1}/\partial w_{1}$ 9.68 9.90 5.59 6.67 0.17 3.79 0.17 3.79
+$\partial p_{1}/\partial w_{2}$ 1.85 2.09 1.15 1.69 0.07 1.41 0.07 1.41
+$\partial p_{2}/\partial x$ 7.48 7.57 4.26 4.91 -0.54 3.36 -0.54 3.36
+$\partial p_{2}/\partial w_{1}$ -4.82 4.94 -2.80 3.34 -0.09 1.90 -0.09 1.90
+$\partial p_{2}/\partial w_{2}$ -3.59 4.05 -2.27 3.33 -0.13 2.82 -0.13 2.82
+$\partial p_{0}/\partial x$ 5.87 5.96 2.91 3.53 -0.53 2.67 -0.53 2.68
+$\partial p_{0}/\partial w_{1}$ -4.86 4.97 -2.79 3.34 -0.08 1.90 -0.08 1.90
+$\partial p_{0}/\partial w_{2}$ 1.74 1.96 1.12 1.64 0.06 1.41 0.06 1.42
+  • This table reports the simulated finite sample bias and RMSE of the MLE and the MERM estimators for the partial derivatives $\partial p_{j}(x,w,\theta_{0})/\partial x$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{1}$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{2}$ for $j\in\{1,2,0\}$ evaluated at the population mean.
+The true values of the marginal effects are $(\partial p_{1}/\partial x,\partial p_{2}/\partial x,\partial p_{0}/\partial x)=(0.222,-0.111,-0.111)$ and zeros for the rest.
+The results are based on 5,000 replications.
+Table 8: Choice of $K$ simulation results for the multinomial logit model, $n=4000$
+MLE $K=2$ $K=4$ data-driven $K$
+bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$
+$\tau=1/4$
+$\partial p_{1}/\partial x$ -3.28 3.41 0.39 1.90 0.56 1.93 0.39 1.90
+$\partial p_{1}/\partial w_{1}$ 2.32 2.57 -0.08 1.59 -0.15 1.57 -0.08 1.59
+$\partial p_{1}/\partial w_{2}$ 0.48 0.72 -0.04 0.62 -0.05 0.61 -0.04 0.62
+$\partial p_{2}/\partial x$ 1.97 2.13 -0.23 1.35 -0.33 1.37 -0.23 1.35
+$\partial p_{2}/\partial w_{1}$ -1.16 1.28 0.04 0.80 0.08 0.78 0.04 0.79
+$\partial p_{2}/\partial w_{2}$ -0.96 1.43 0.07 1.23 0.11 1.23 0.07 1.23
+$\partial p_{0}/\partial x$ 1.31 1.49 -0.16 1.01 -0.23 1.03 -0.16 1.01
+$\partial p_{0}/\partial w_{1}$ -1.16 1.29 0.04 0.80 0.08 0.78 0.04 0.80
+$\partial p_{0}/\partial w_{2}$ 0.47 0.71 -0.04 0.62 -0.05 0.62 -0.04 0.62
+$\tau=1/2$
+$\partial p_{1}/\partial x$ -9.00 9.03 -1.52 2.41 0.49 2.03 0.48 2.04
+$\partial p_{1}/\partial w_{1}$ 6.43 6.52 1.06 1.99 -0.11 1.65 -0.10 1.65
+$\partial p_{1}/\partial w_{2}$ 1.28 1.38 0.21 0.68 -0.04 0.65 -0.04 0.65
+$\partial p_{2}/\partial x$ 5.22 5.27 0.90 1.64 -0.30 1.46 -0.29 1.47
+$\partial p_{2}/\partial w_{1}$ -3.21 3.25 -0.53 1.00 0.05 0.82 0.05 0.83
+$\partial p_{2}/\partial w_{2}$ -2.51 2.70 -0.42 1.35 0.09 1.30 0.09 1.30
+$\partial p_{0}/\partial x$ 3.78 3.83 0.62 1.19 -0.20 1.10 -0.19 1.10
+$\partial p_{0}/\partial w_{1}$ -3.22 3.26 -0.53 1.00 0.05 0.82 0.05 0.83
+$\partial p_{0}/\partial w_{2}$ 1.23 1.32 0.21 0.67 -0.05 0.65 -0.05 0.65
+$\tau=3/4$
+$\partial p_{1}/\partial x$ -13.37 13.38 -6.49 6.79 0.41 2.25 0.41 2.25
+$\partial p_{1}/\partial w_{1}$ 9.68 9.73 4.44 4.84 -0.02 1.80 -0.02 1.80
+$\partial p_{1}/\partial w_{2}$ 1.80 1.87 0.91 1.12 -0.03 0.69 -0.03 0.69
+$\partial p_{2}/\partial x$ 7.48 7.50 3.83 4.08 -0.24 1.63 -0.24 1.63
+$\partial p_{2}/\partial w_{1}$ -4.82 4.85 -2.22 2.42 0.01 0.90 0.01 0.90
+$\partial p_{2}/\partial w_{2}$ -3.50 3.63 -1.80 2.22 0.06 1.39 0.06 1.39
+$\partial p_{0}/\partial x$ 5.89 5.91 2.66 2.87 -0.16 1.22 -0.16 1.22
+$\partial p_{0}/\partial w_{1}$ -4.86 4.89 -2.22 2.42 0.01 0.90 0.01 0.90
+$\partial p_{0}/\partial w_{2}$ 1.70 1.76 0.89 1.09 -0.03 0.69 -0.03 0.69
+  • This table reports the simulated finite sample bias and RMSE of the MLE and the MERM estimators for the partial derivatives $\partial p_{j}(x,w,\theta_{0})/\partial x$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{1}$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{2}$ for $j\in\{1,2,0\}$ evaluated at the population mean.
+The true values of the marginal effects are $(\partial p_{1}/\partial x,\partial p_{2}/\partial x,\partial p_{0}/\partial x)=(0.222,-0.111,-0.111)$ and zeros for the rest.
+The results are based on 5,000 replications.
+Appendix G MERM derivation when $\sigma_{\varepsilon}$ is not small
 Note that $\tau$ can be small without $\sigma_{\varepsilon}$ being small in absolute magnitude.
@@
 Remember the model of Section 2:
-$\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta_{0})]=0,\quad X_{i}=X_{i}^{\ast}+\varepsilon_{i},\quad\mathbb{E}[\varepsilon_{i}]=0.$ (F.1)
+$\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta_{0})]=0,\quad X_{i}=X_{i}^{\ast}+\varepsilon_{i},\quad\mathbb{E}[\varepsilon_{i}]=0.$ (G.1)
 Let $\xi_{i}$ denote a random variable with $\mathbb{E}\left[\xi_{i}\right]=0$ and $\mathbb{E}\left[\xi_{i}^{2}\right]=1$, $\mathbb{E}\left[\left|\xi_{i}\right|^{L+1}\right]$ is bounded, and $\varepsilon_{i}\equiv\sigma_{\varepsilon}\xi_{i}$.
@@
 $\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}},\quad\widetilde{X}_{i}\equiv X_{i}/\sigma_{X^{\ast}},\quad\widetilde{X}_{i}^{\ast}\equiv X_{i}^{\ast}/\sigma_{X^{\ast}},\quad\widetilde{g}(\widetilde{x},s,\theta)\equiv g(\sigma_{X^{\ast}}\widetilde{x},s,\theta).$
-Then, we can rewrite equation (F.1) as
+Then, we can rewrite equation (G.1) as
 $\mathbb{E}[\widetilde{g}(\widetilde{X}_{i}^{\ast},S_{i},\theta_{0})]=0,\quad\widetilde{X}_{i}=\widetilde{X}_{i}^{\ast}+\tau\xi_{i},\quad\mathbb{E}[\xi_{i}]=0.$
@@
 Then, the corrected moment condition has the form
-$\widetilde{\psi}(\widetilde{X}_{i},S_{i},\theta,\widetilde{\gamma})=\widetilde{g}(\widetilde{X}_{i},S_{i},\theta)-\sum_{k=2}^{K}\widetilde{\gamma}_{k}\widetilde{g}_{x}^{(k)}(\widetilde{X}_{i},S_{i},\theta),$ (F.2)
+$\widetilde{\psi}(\widetilde{X}_{i},S_{i},\theta,\widetilde{\gamma})=\widetilde{g}(\widetilde{X}_{i},S_{i},\theta)-\sum_{k=2}^{K}\widetilde{\gamma}_{k}\widetilde{g}_{x}^{(k)}(\widetilde{X}_{i},S_{i},\theta),$ (G.2)
 where true parameter values $\widetilde{\gamma}_{0}$ are $\widetilde{\gamma}_{02}=\tau^{2}\mathbb{E}\left[\xi_{i}^{2}\right]/2=\tau^{2}/2$, $\widetilde{\gamma}_{03}=\tau^{3}\mathbb{E}\left[\xi_{i}^{3}\right]/6$, and $\widetilde{\gamma}_{0k}=\frac{\tau^{k}\mathbb{E}\left[\xi_{i}^{k}\right]}{k!}-\sum_{\ell=2}^{k-2}\frac{\tau^{k-\ell}\mathbb{E}\left[\xi_{i}^{k-\ell}\right]}{(k-\ell)!}\widetilde{\gamma}_{0\ell}$ for $k\geq 4$.
@@
 $=$ $\sigma_{X^{\ast}}^{k}\left(\frac{\tau^{k}\mathbb{E}\left[\xi_{i}^{k}\right]}{k!}-\sum_{\ell=2}^{k-2}\frac{\tau^{k-\ell}\mathbb{E}\left[\xi_{i}^{k-\ell}\right]}{(k-\ell)!}\widetilde{\gamma}_{0\ell}\right)=\sigma_{X^{\ast}}^{k}\widetilde{\gamma}_{0k}\text{.}$
-Finally, let us now show that moment condition $\widetilde{\psi}$ in equation (F.2) is numerically identical to $\psi$ in equation (11) with $\gamma_{k}=\sigma_{X^{\ast}}^{k}\widetilde{\gamma}_{k}$.
+Finally, let us now show that moment condition $\widetilde{\psi}$ in equation (G.2) is numerically identical to $\psi$ in equation (11) with $\gamma_{k}=\sigma_{X^{\ast}}^{k}\widetilde{\gamma}_{k}$.
 Note that for $\widetilde{x}=x/\sigma_{X^{\ast}}$ we have $\widetilde{g}_{\widetilde{x}}^{(k)}(\widetilde{x},s,\theta)\equiv\nabla_{\widetilde{x}}^{k}g(\sigma_{X^{\ast}}\widetilde{x},s,\theta)=\sigma_{X^{\ast}}^{k}g_{a}^{(k)}(a,s,\theta)|_{a=\sigma_{X^{\ast}}\widetilde{x}}=\sigma_{X^{\ast}}^{k}g_{x}^{(k)}(x,s,\theta)$, and hence
-$\widetilde{\psi}(\widetilde{X}_{i},S_{i},\theta,\widetilde{\gamma})$ $=$ $g(\sigma_{X^{\ast}}\widetilde{X}_{i},S_{i},\theta)-\sum_{k=2}^{K}\left(\widetilde{\gamma}_{k}\sigma_{X^{\ast}}^{k}\right)g_{x}^{(k)}(\sigma_{X^{\ast}}\widetilde{X}_{i},S_{i},\theta)$
+[…]
 $=$ $g(X_{i},S_{i},\theta)-\sum_{k=2}^{K}\left(\widetilde{\gamma}_{k}\sigma_{X^{\ast}}^{k}\right)g_{x}^{(k)}(X_{i},S_{i},\theta)$
 $=$ $\psi(X_{i},S_{i},\theta,\gamma)\text{.}$
-Appendix G Some Implementation Details
+Appendix H Some Implementation Details
 Numerical Optimization
@@
 As for the standard GMM estimator, the optimal weighting matrix can be estimated by
-$\hat{\Xi}_{\text{eff}}\equiv\hat{\Omega}_{\psi\psi}^{-1}(\tilde{\theta},\tilde{\gamma}),$
+[…]
 where $\tilde{\theta}$ and $\tilde{\gamma}$ are some preliminary estimators of $\theta_{0}$ and $\gamma_{0}$, and $\hat{\Omega}_{\psi\psi}(\theta,\gamma)\equiv n^{-1}\sum_{i=1}^{n}\psi_{i}(\theta,\gamma)\psi_{i}(\theta,\gamma)^{\prime}.$ One example of such a preliminary estimator would be the 1-step (GMM-)MERM estimator using $\hat{\Xi}_{\text{GMM}1}\equiv\hat{\Omega}_{\psi\psi}^{-1}(\hat{\theta}_{\text{Naive}},0)$ as the first-step GMM weighting matrix, where $\hat{\theta}_{\text{Naive}}$ is a naive estimator of $\theta_{0}$ that ignores EIV.
@@
 $\hat{\Sigma}\equiv(\hat{\Psi}^{\prime}\hat{\Xi}\hat{\Psi})^{-1}\hat{\Psi}^{\prime}\hat{\Xi}\hat{\Omega}_{\psi\psi}\hat{\Xi}\hat{\Psi}(\hat{\Psi}^{\prime}\hat{\Xi}\hat{\Psi})^{-1},$
-where, $\hat{\Xi}$ is the chosen weighting matrix, and $\hat{\Psi}\equiv\overline{\Psi}(\hat{\theta},\hat{\gamma})=n^{-1}\sum_{i=1}^{n}\Psi_{i}(\hat{\theta},\hat{\gamma})$ and $\hat{\Omega}_{\psi\psi}=\hat{\Omega}_{\psi\psi}(\hat{\theta},\hat{\gamma})$ are estimators of $\Psi$ and $\Omega_{\psi\psi}$.
-Appendix H Implementation Details of the Empirical Illustration
+[…]
+Appendix I Implementation Details of the Empirical Illustration
 In this section, we provide additional details on the implementation of the numerical experiment in Section 3.3.
@@
 This dataset has been extensively used in transportation research.
-For a detailed description of the dataset see, for example, Koppelman and Wen (2000), Wen and Koppelman (2001), and Hansen (2022).
-As in Koppelman and Wen (2000), we use only the subset of travelers who chose train, air, or car (and had all of those alternatives available for them), which leaves $n=2769$ observations.
+For a detailed description of the dataset see, for example, koppelman2000paired, wen2001generalized, and Hansen2021-Textbook.
+As in koppelman2000paired, we use only the subset of travelers who chose train, air, or car (and had all of those alternatives available for them), which leaves $n=2769$ observations.
 Monte-Carlo design
@@
 $\tau=1/4$
-$\theta_{1}$0.0021 0.0035 0.0041 8.70 0.0001 0.0042 0.0042 5.48 0.0005 0.0057 0.0058 7.38
+$\theta_{1}$ -0.0021 0.0035 0.0041 8.70 0.0001 0.0042 0.0042 5.48 0.0005 0.0057 0.0058 7.38
 $\theta_{2}$ 0.0047 0.0932 0.0933 5.10 0.0028 0.0957 0.0957 5.36 0.0022 0.0960 0.0960 5.40
 $\theta_{3}$ 0.1152 0.4452 0.4599 6.00 -0.0048 0.4821 0.4821 5.94 -0.0251 0.5336 0.5342 6.68
-$\theta_{4}$0.0004 0.0031 0.0031 4.52 -0.0001 0.0034 0.0034 5.32 -0.0001 0.0036 0.0036 6.86
-$\theta_{5}$0.0023 0.0894 0.0895 5.18 -0.0088 0.0918 0.0922 5.54 -0.0113 0.0922 0.0929 5.96
+$\theta_{4}$ -0.0004 0.0031 0.0031 4.52 -0.0001 0.0034 0.0034 5.32 -0.0001 0.0036 0.0036 6.86
+$\theta_{5}$ -0.0023 0.0894 0.0895 5.18 -0.0088 0.0918 0.0922 5.54 -0.0113 0.0922 0.0929 5.96
 $\theta_{6}$ 0.0232 0.1821 0.1836 4.64 0.0250 0.1982 0.1998 5.72 0.0329 0.2089 0.2115 6.74
-$\theta_{7}$0.0001 0.0035 0.0035 5.58 -0.0002 0.0036 0.0036 6.24 -0.0003 0.0036 0.0037 6.04
-$\theta_{8}$0.0001 0.0007 0.0007 4.82 -0.0001 0.0007 0.0007 5.48 -0.0002 0.0007 0.0007 5.58
+$\theta_{7}$ -0.0001 0.0035 0.0035 5.58 -0.0002 0.0036 0.0036 6.24 -0.0003 0.0036 0.0037 6.04
+$\theta_{8}$ -0.0001 0.0007 0.0007 4.82 -0.0001 0.0007 0.0007 5.48 -0.0002 0.0007 0.0007 5.58
 $\tau=1/2$
-$\theta_{1}$0.0073 0.0032 0.0080 60.08 -0.0016 0.0043 0.0046 6.86 0.0005 0.0061 0.0061 6.60
+$\theta_{1}$ -0.0073 0.0032 0.0080 60.08 -0.0016 0.0043 0.0046 6.86 0.0005 0.0061 0.0061 6.60
 $\theta_{2}$ 0.0109 0.0930 0.0936 5.18 0.0050 0.0959 0.0960 5.54 0.0026 0.0964 0.0965 5.36
 $\theta_{3}$ 0.4080 0.4452 0.6039 17.12 0.0936 0.4874 0.4963 6.58 -0.0263 0.5475 0.5481 6.38
-$\theta_{4}$0.0012 0.0029 0.0031 6.46 -0.0003 0.0035 0.0035 5.22 -0.0002 0.0038 0.0038 6.52
-$\theta_{5}$0.0006 0.0894 0.0894 5.16 -0.0083 0.0919 0.0923 5.52 -0.0110 0.0924 0.0930 5.92
+$\theta_{4}$ -0.0012 0.0029 0.0031 6.46 -0.0003 0.0035 0.0035 5.22 -0.0002 0.0038 0.0038 6.52
+$\theta_{5}$ -0.0006 0.0894 0.0894 5.16 -0.0083 0.0919 0.0923 5.52 -0.0110 0.0924 0.0930 5.92
 $\theta_{6}$ 0.0655 0.1752 0.1870 6.22 0.0348 0.2035 0.2064 5.86 0.0326 0.2158 0.2183 6.42
-$\theta_{7}$0.0003 0.0035 0.0036 5.64 -0.0003 0.0036 0.0037 6.34 -0.0003 0.0037 0.0037 6.06
-$\theta_{8}$0.0001 0.0007 0.0007 5.06 -0.0001 0.0007 0.0007 5.54 -0.0002 0.0007 0.0007 5.48
+$\theta_{7}$ -0.0003 0.0035 0.0036 5.64 -0.0003 0.0036 0.0037 6.34 -0.0003 0.0037 0.0037 6.06
+$\theta_{8}$ -0.0001 0.0007 0.0007 5.06 -0.0001 0.0007 0.0007 5.54 -0.0002 0.0007 0.0007 5.48
 $\tau=3/4$
-$\theta_{1}$0.0132 0.0029 0.0135 99.34 -0.0056 0.0043 0.0071 25.12 0.0003 0.0065 0.0065 6.06
+$\theta_{1}$ -0.0132 0.0029 0.0135 99.34 -0.0056 0.0043 0.0071 25.12 0.0003 0.0065 0.0065 6.06
 $\theta_{2}$ 0.0180 0.0923 0.0940 5.36 0.0102 0.0961 0.0966 5.76 0.0033 0.0973 0.0973 5.44
 $\theta_{3}$ 0.7336 0.4496 0.8604 41.66 0.3203 0.4859 0.5820 12.00 -0.0130 0.5666 0.5667 6.20
-$\theta_{4}$0.0024 0.0026 0.0035 14.00 -0.0009 0.0035 0.0036 5.94 -0.0002 0.0041 0.0041 5.76
+$\theta_{4}$ -0.0024 0.0026 0.0035 14.00 -0.0009 0.0035 0.0036 5.94 -0.0002 0.0041 0.0041 5.76
 $\theta_{5}$ 0.0021 0.0890 0.0891 5.08 -0.0071 0.0921 0.0924 5.68 -0.0109 0.0926 0.0932 5.82
 $\theta_{6}$ 0.1204 0.1654 0.2046 9.76 0.0648 0.2048 0.2148 6.56 0.0334 0.2294 0.2318 5.98
-$\theta_{7}$0.0004 0.0036 0.0036 6.00 -0.0004 0.0036 0.0037 6.34 -0.0003 0.0037 0.0037 6.06
-$\theta_{8}$0.0001 0.0007 0.0007 5.54 -0.0002 0.0007 0.0008 6.00 -0.0002 0.0007 0.0008 5.42
+$\theta_{7}$ -0.0004 0.0036 0.0036 6.00 -0.0004 0.0036 0.0037 6.34 -0.0003 0.0037 0.0037 6.06
+$\theta_{8}$ -0.0001 0.0007 0.0007 5.54 -0.0002 0.0007 0.0008 6.00 -0.0002 0.0007 0.0008 5.42
   • This table reports the simulated finite sample bias, standard deviation, RMSE, and size of the MLE and the MERM estimators and the corresponding t-tests for the components of $\theta_{0}$.